### v1

# Some insights into depth estimators for location and scatter in the multivariate setting

Abstract
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to other statistical models. These procedures aim to formalize the idea of identifying deeply embedded fits to a model that are less influenced by contamination. Since contamination introduces bias in estimators, it is well known in the location model that the median minimizes the worst-case performance, in terms of maximum bias, among all equivariant estimators. In the multivariate case, Tukey’s median was a groundbreaking concept for location estimation, and its counterpart for scatter matrices has recently attracted considerable interest. The breakdown point and the maximum asymptotic bias are key concepts used to summarize an estimator’s behavior under contamination. For the location and scale model, we consider two closely related depth formulations, whose deepest estimators display significantly different behavior in terms of breakdown point. In the multivariate setting, we analyze recently introduced concentration inequalities that provide a unified framework for studying both the statistical convergence rate and robustness of Tukey’s median and depth-based scatter matrices. We observe that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. Since the maximum bias for depth-based scatter matrices had not previously been derived, we explicitly calculate both the breakdown point and the maximum bias curve for the deepest scatter matrices.
Keywords: Multivariate scatter depth, Breakdown Point, Minimax Bias Function
1 Introduction
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to more general statistical models. Originally introduced by Tukey (1975), depth was defined as the minimum proportion of data points lying on either side of a point, and this idea was later extended to the bivariate case. Donoho and Gasko (1992) formalized the general notion of the depth of a point $\mathbf{z}$ in a $p$-dimensional space with respect to a probability measure $P$.
The seminal paper by Huber (1964) highlighted the median as an estimator with optimal worst-case behavior under contamination (i.e., its maximum bias), making it the most robust among the class of equivariant location estimators. A key feature of the median is that it is flanked on each side by half of the data mass, helping to shield it from the effects of outliers. This robustness—being well-surrounded by data—is echoed in Tukey’s concept of the multivariate median. In this case, a point in Euclidean space is considered a median if it maximizes the minimum mass contained in any closed half-space whose boundary includes the point.
This pursuit of a point that is ”deeply inserted” into the data, to enhance robustness in univariate and multivariate location models, has been extended to other statistical frameworks. In the regression setting, Rousseeuw and Hubert (1999) defined regression depth, which measures how deeply a linear fit is embedded within the data. This is evaluated by the smallest amount of data mass in the two opposing wedges formed by the fit plane and vertical planes orthogonal to the explanatory subspace. Mizera and Müller (2004) further explored depth concepts in the context of location-scale models.
Chen and Tyler (2002) investigated various properties of Tukey’s median, including its influence function and maximum contamination bias. Chen et al. (2018) extended the idea of depth to covariance matrix estimation by introducing the concept of matrix depth. Their estimator achieves the optimal rate under Huber’s $\varepsilon$-contamination model for estimating covariance/scatter matrices with various structural assumptions, such as bandedness and sparsity. Paindaveine and Van Bever (2018) also developed halfspace depth concepts for scatter, concentration, and shape matrices. While their concept of scatter depth is similar to that in Chen et al. (2018) and Zhang (2002), rather than focusing on the deepest scatter matrix, they studied the properties of the depth function and its associated depth regions.
Nagy et al. (2019) discussed the relationships between halfspace depth and problems in affine and convex geometry, offering an extensive overview of various depth notions.
Section 2 presents two proposals for joint location and scale depth estimation in the univariate location-scale model, comparing their differing breakdown behaviors despite conceptual similarity. Section 3 discusses the importance of maximum asymptotic bias in understanding how estimator behavior evolves as sample size $n$ becomes large. By examining the concentration inequalities in Chen et al. (2018), we show that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. In particular, Section 4 derives the maximum bias function for the deepest scatter matrix estimator introduced by Chen et al. (2018) in the case of known location, showing that this estimator shares Tukey’s median’s asymptotic breakdown point of $1/3$. All proofs are deferred to the Appendix.
2 The location-scale model, depth and breakdown point
Let us take the multivariate location and scatter model,.
$X\mathbf{=\mu}_{0}\mathbf{+V}_{0}\mathbf{u}$ (1)
with $\mathbf{\mu}_{0}\in\mathbb{R}^{p},$ $\mathbf{V}_{0}\in\mathbb{R}^{p\times p}$ an inversible matrix, and $\mathbf{u}$ an inobservable random vector with centrosymmetric distribution $P_{\mathbf{0}}$ around $\mathbf{0}.$ $\mathbf{\mu}_{0}$ and $\mathbf{V}_{0}\mathbf{V}_{0}^{t}$ (except for a constant) are identifiable. Tukey’s depth of a vector $\mathbf{\theta}\in\mathbb{R}^{p}$ is given by
$D_{T}\left(\mathbf{\theta},P\right)=\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min P\left(\mathbf{u}^{t}X\leq\mathbf{u}^{t}\mathbf{\theta}\right).$
with $\mathcal{S}^{p-1}{\Large=}\left\{\mathbf{z\in}\mathbb{R}^{p}:\left\|\mathbf{z}\right\|=1\right\}.$ Carrizosa (1996) and Adrover et al. (2002) independently came up with a residual smallness concept which comprises the notion of depth given in the univariate and multivariate model as well as the regression model. More precisely,
$\mathcal{D}_{T}\left(\mathbf{\theta},P\right){\Large=}\inf_{\left\|\mathbf{\lambda}\right\|=1,\mathbf{\gamma\in\mathbb{R}}^{p}}P\left(\left|\mathbf{\lambda}^{t}(\mathbf{x-\theta)}\right|\leq\left|\mathbf{\lambda}^{t}(\mathbf{x-\gamma)}\right|\right)$ (2)
and Tukey’s median is taken to be
$\mathbf{\hat{\theta}}_{T}\left(P\right)=\arg\sup_{\mathbf{\theta}}\mathcal{D}_{T}\left(\mathbf{\theta},P\right).$
The idea behind that definition is that the depth of a fit $\mathbf{\theta}$ is determined by the bad performance displayed by the residuals $\left|\mathbf{\lambda}^{t}(\mathbf{x-\theta)}\right|$ compared to the best competitor $\mathbf{\gamma,}$ whose residuals $\left|\mathbf{\lambda}^{t}(\mathbf{x-\gamma)}\right|$ have the minimum probability of being worse than those of $\left|\mathbf{\lambda}^{t}(\mathbf{x-\theta)}\right|$, therefore the $\mathbf{\theta}$ with the best worst performance is singled out. Observe that, if $P=P_{\mathbf{0}}\left(\mathbf{V}_{0}^{-1}\left(\cdot-\mathbf{\mu}_{0}\right)\right),$ $\mathbf{\hat{\theta}}_{T}\left(P\right)=\mathbf{\mu}_{0}.$
If $p=1,$ the deepest location estimator is the median. In the univariate case the model (1) is the usual location-scale model $Y=\mu_{0}+\sigma_{0}U$, $\sigma_{0}>0,$ with $U\sim P_{0}$ centrosymmetric. The residual smallness concept in (2) is easily adapted for a joint estimation of location and scale as it follows: The depth of $\left(\mu,\sigma\right)\in\mathbb{R}\times(0,\infty)$ is taken to be,
$D_{LS}^{1}\left(\mu,\sigma,P\right)$ $=$ $\min\left\{\begin{array}[]{c}\inf_{\lambda\in\mathbb{R}}P\left(\left[\left|Y-\mu\right|\leq\left|Y-\lambda\right|\right]\right),\\ \inf_{\gamma>0}P\left(\left[\left|\left|\frac{Y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{Y-\mu}{\gamma}\right|-1\right|\right]\right)\end{array}\right\}$
$\left(\hat{\mu}_{1},\hat{\sigma}_{1}\right)$ $=$ $\arg\max_{\mu,\sigma}D_{LS}^{1}\left(\mu,\sigma,P\right).$
With this definition, we get very well known functionals for location and scale, as it is stated in the following lemma.
[…]
If $\Theta$ denotes the set in which the parameters are assumed to lie in the central model (for instance, $\Theta=\mathbb{R}\times\left(0,\infty\right)$ in the location-scale model), the breakdown point of an estimator is a value that quantifies the level of contamination required to cause the estimator to move outside any compact subset of $\Theta$. A usual way to model the presence of spurious data in a sample is the $\varepsilon-$contamination neighborhood, which is defined to be,
$\mathcal{P}_{\varepsilon}\left(F_{0}\right)=\left\{\left(1-\varepsilon\right)F_{0}\left(\frac{\cdot-\mu_{0}}{\sigma_{0}}\right)+\varepsilon G\left(\cdot\right),\text{ }G\text{ any distribution on }\mathbb{R}\right\}.$
and the asymptotic breakdown point of $\left(\varepsilon_{L}^{\ast},\varepsilon_{S}^{\ast}\right)$ of a functional $\left(\hat{\mu},\hat{\sigma}\right)$:$\mathcal{P}_{\varepsilon}\left(F_{0}\right)\rightarrow\Theta$ is defined to be,
$\varepsilon_{L}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}|\hat{\mu}\left(\cdot\right)|=\infty\right\},\text{ }$
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}\max\left\{\hat{\sigma}\left(\cdot\right),\hat{\sigma}^{-1}\left(\cdot\right)\right\}=\infty\right\}.$
$\hat{\mu}_{1}$ as well as $\hat{\sigma}_{1}$ are estimators whose breakdown point is $0.5,$ the best breakdown point attainable for location and scale equivariant estimators. Let us suppose now that we bind together the estimation of both parameters in one expression, by considering the following second proposal for location-scale depth,
$D_{LS}^{2}\left(\mu,\sigma,P\right)$
$=$ $\inf_{\gamma>0,\lambda\in\mathbb{R}}P\left(\left[\left|Y-\mu\right|\leq\left|Y-\lambda\right|\right]\cap\left[\left|\left|\frac{Y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{Y-\mu}{\gamma}\right|-1\right|\right]\right),$
and the deepest estimator as
$\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\arg\max_{\mu\in\mathbb{R},\sigma>0}D_{LS}^{2}\left(\mu,\sigma,P\right).$
It is easy to see that
$D_{LS}^{2}\left(\mu,\sigma,P\right)=\min\left\{\begin{array}[]{cc}P\left(\mu-\sigma\leq y\leq\mu\right),&P\left(\mu\leq y\leq\mu+\sigma\right),\\ P\left(\left[y\leq\mu-\sigma\right]\right),&P\left(\left[y\geq\mu+\sigma\right]\right).\end{array}\right\}.$
If $P=F_{0}\left(\sigma_{0}^{-1}\left(\cdot-\mu_{0}\right)\right),\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\left(\mu_{0},\sigma_{0}\right).$ If $P$ is a symmetric distribution, $\hat{\mu}^{2}=med\left(P\right),$ $\hat{\sigma}^{2}=med\left(P\right)-q_{1}\left(P\right),$ $q_{1}\left(P\right)$ the first quartile of $P.$ However, we can see in the Appendix that $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ has breakdown point much lower than that of the $D_{LS}^{1}.$ We will assume that $F_{0}=\Phi$, the standard normal distribution. Let $\varepsilon_{0}$ be the solution to the equation
$\delta=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right).$
with $h\left(x,y\right)=\Phi\left(x\right)-\Phi\left(\frac{y+x}{2}\right)$ and $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$
Theorem 1
The breakdown point $\varepsilon^{\ast}$ of $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ is $\varepsilon^{\ast}=\varepsilon_{0}$ with $41/5<\varepsilon_{0}<1/$ 4 .
The second proposal for depth estimators in the location-scale model yields estimators with much lower breakdown point compared to that of the first proposal. This phenomenum can be considered in a broader scenario showing a behavior also described in some other simultaneous procedures in the robust literature. For instance, a similar behavior is described in the simultaneous M-estimation for location and scale (Huber and Ronchetti, 2009, p.141) and in the context of simultaneous M-estimation for regression and scale in Maronna and Yohai (1991) which seems to prevent from doing simultaneous estimation in order to improve its robustness.
3 Maximum bias and concentration inequalities
The asymptotic breakdown point is an aspect of a more powerful tool to measure the performance of an estimator. The so-called maximum asymptotic bias function depicts more accurately the behavior of the estimator since it captures the global behavior by quantifying how much the estimator deviates in the whole contamination neighborhood under diferent levels of contaminations from the parameters at the central model. In spite of the deep understanding that the maxbias function provides, it is usually neglected because of the technicalities its derivation requires. .
$\mathcal{P}_{\varepsilon}\left(P_{0}\right)=\left\{\left(1-\varepsilon\right)P_{\mathbf{0}}\left(\mathbf{V}_{0}^{-1}\left(\cdot-\mathbf{\mu}_{0}\right)\right)+\varepsilon G\left(\cdot\right),\text{ }G\text{ any distribution on }\mathbb{R}^{p}\right\}$
[…]
$\mathbf{\hat{\mu}}\left(\mathcal{L}\left(A\mathbf{x+b}\right)\right)$ $=$ $A\mathbf{\hat{\mu}}\left(\mathcal{L}\left(\mathbf{x}\right)\right)\mathbf{+b}$
$\hat{\Gamma}\left(\mathcal{L}\left(A\mathbf{x+b}\right)\right)$ $=$ $A\hat{\Gamma}\left(\mathcal{L}\left(\mathbf{x}\right)\right)A^{t}$
We say that the functionals $\mathbf{\hat{\mu}}$ and $\hat{\Gamma}$ are Fisher consistent if
$\mathbf{\hat{\mu}}\left(P_{0}^{E}\right)=\mathbf{\mu}_{0}\text{ and }\hat{\Gamma}\left(P_{0}^{E}\right)=c\left(\mathbf{V}_{0}\mathbf{V}_{0}^{t}\right),\text{ }c>0\text{ for any }\mathbf{\mu}_{0},\mathbf{V}_{0}\in\mathbb{R}^{p\times p}\text{ invertible. }$.
The effect of the distorsion caused by having $P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$, it is customary measured in the following invariant manner,
$b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right)$ $=$ $\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right)^{t}\left(\hat{\Gamma}\left(P_{0}^{E}\right)\right)^{-1}\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right)$
$b_{S}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}\left(P_{0}\right)\mathbf{u}}$
$=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\lambda_{(1)}\left(\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1}\right),$
$b_{I}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\inf_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}\left(P_{0}\right)\mathbf{u}}$
$=$ $\lambda_{(p)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\lambda_{(p)}\left(\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1}\right)$
$=$ $\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}^{-1}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}^{-1}\left(P_{0}\right)\mathbf{u}}=\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{1/2}\hat{\Gamma}^{-1}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{1/2}\right).$
Then we can define the asymptotic biases for the largest and smallest eigenvalues as
$B_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right)$
$B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{S}\left(\hat{\Gamma},\varepsilon,P\right),\text{ }B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}^{E}\right)=\sup_{P\in\mathcal{P}_{\varepsilon}}b_{I}\left(\hat{\Gamma},\varepsilon,P\right).$
$B\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\max\left\{B_{S}\left(\hat{\Gamma},\varepsilon\right),B_{I}\left(\hat{\Gamma},\varepsilon\right)\right\}$
We say that the asymptotic explosion and implosion breakdown points are given by
$\varepsilon_{L}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{L}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\}$
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{S}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\},\text{ }\varepsilon_{I}^{\ast}=\inf\left\{\varepsilon>0:B_{I}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\},$
$\varepsilon^{\ast}$ $=$ $\min\left(\varepsilon_{S}^{\ast},\varepsilon_{I}^{\ast}\right).$
If we consider equivariant and Fisher consistent functionals $\mathbf{\hat{\mu}}$ and $\hat{\Gamma}$ for location and scatter, it is easily proved that
$B_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}^{E}\right)$ $=$ $cB_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}\right)$
$B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}\right)$
$B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}^{E}\right)$ $=$ $B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}\right)$
which entails that the maximum bias can be computed using $\mathbf{\mu}_{0}=\mathbf{0}$ and $\Sigma_{0}=I.$ The maximum bias for Tukey’s median turns out to be $B\left(\hat{\theta}_{T},\varepsilon,\Phi\right)=\Phi^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$ (Chen and Tyler (2002)). Interest in scatter depth has been revitalized over the past decade, particularly following the seminal papers by Chen et al. (2018) and Paindaveine and Van Bever (2018), which explored different aspects of the concept of depth for multivariate scatter. If $X\sim P,$ $X\in\mathbb{R}^{p},\Gamma\in\mathcal{E}$, the depth of $\Gamma$ is taken to be
$D\left(\Gamma,P\right)=\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{\begin{array}[]{c}P\left(\left|\mathbf{u}^{t}\left(X-\mathbf{v}_{0}\left(P\right)\right)\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P\left(\left|\mathbf{u}^{t}\left(X-\mathbf{v}_{0}\left(P\right)\right)\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\},$
with $\mathbf{v}_{0}$ a preliminary affine equivariant location functional to yield an affine equivariant multivariate scatter functional If location is known we can assume without loss of generality that $\mathbf{v}_{0}=\mathbf{0.}$ For the known location case, the depth and the deepest estimator are defined to be,
$D\left(\Gamma,P\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}.$
$\hat{\Gamma}$ $=$ $\arg\underset{\Gamma\succeq\mathbf{0}}{\max}D\left(\Gamma,P\right).\text{ \ \ }D_{M}\left(P\right)=D\left(\hat{\Gamma},P\right),$
($\succeq$ defines the usual partial order in the set $\mathcal{E}$ of symmetric positive matrices).
Chen et al.. (2018) broke new ground by introducing a unified way to study the statistical convergence rate and robustness jointly. Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ They derived that, for $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\boldsymbol{\hat{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha$ (4)
(Theorem 2.1 of Chen et al. (2018)). The constant $C$ in the concentration inequality (4) is actually affected by the asymptotic maximum bias of Tukey’s median.
However, the bound (4) can be derived in a more illuminating manner by explicitly incorporating the maximum bias, as the maximum bias governs the behavior of the estimator when the sample size is sufficiently large. Without enlarging the bound in (4), we obtain a more informative inequality,
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\hat{\boldsymbol{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq\tilde{C}\left(\max\left\{\frac{p}{n},B^{2}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,\Phi\right)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
(see Lemma 9 in the Appendix).
The reasoning behind expecting the asymptotic maximum bias function to appear in the concentration inequality is as follows: As the sample size tends to infinity, the estimator is expected to converge to the functional value $\hat{\boldsymbol{\theta}}_{T}\left(P\right).$The quantity $\left\|\hat{\boldsymbol{\theta}}_{T}\left(P\right)-\boldsymbol{\theta}\right\|$ remains within a range bounded above by the maximum bias corresponding to the given level of contamination, since the distributions vary over the entire $\varepsilon$-contamination neighborhood.
On the other hand, Chen et al. (2018) also derived a concentration inequality for the deepest estimator for the dispersion matrix. They showed that, with probability at least $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$ (5)
With a similar reasoning to that of Lemma 9 we can obtain that (see Lemma 10 below)
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
with $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right].$ By these means, the concentration inequality is also able to uncover the maximum bias of the deepest as it will be shown in the next Section.
4 Maximum bias for the deepest scatter matrix
From now onwards we will assume that
A1. $P_{0}$ is a multivariate normal around $\mathbf{0}$ and covariance matrix $I.$
The bounds found using the concentration inequalities of Chen et al. (2018) suggest to study the asymptotic maximum bias of the deepest scatter matrix to verify the accuracy of the bounds. We will summaryze the steps to get the maximum bias function for the deepest scatter matrix. First of all we compute the depth of any matrix under the normal model and the deepest estimator under this distribution, which is given by $\Phi^{-1}\left(3/4\right)I.$ We next verify that in case of having a sequence of contaminations in the $\varepsilon-$contamination neighborhood which yields the deepest estimator to have either the largest eigenvalue going to infinity or the smallest one going to $0,$ then the level of contamination $\varepsilon\geq 1/3,$ which entails that $\varepsilon^{\ast}\geq 1/3.$ Next, we calculate the depth of any matrix in $\mathcal{E}$ under point mass contaminations located along the direction given by a vector $\mathbf{e\in\mathbb{R}}^{p}$. Moreover we consider the set of the matrices whose eigenvector associated with the largest eigenvalue coincides with the vector $\mathbf{e}$, which yields a more specific formula rather the general one for any matrix given in the previous step. Then, in this class, we calculate a deepest scatter matrix (since uniqueness cannot be suspected at all) in which the largest and smallest eigenvalues  attainable for point mass contaminations coincide with those of the bounds found in the concentration inequalities. On the one hand we show that in case of having a sequence of matrices such that the maximum eigenvalue tends to infinity its depth could only converge to a values less or equal than $\min\left(\varepsilon,1-\varepsilon\right).$ On the other hand, we prove that the deepest estimator in the case of using point mass contaminations is less or equal to $\left(1-\varepsilon\right)/2$. Therefore we can conclude that if we have a level of contamination $\varepsilon$ such that the deepest estimator have smallest and largest eigenvalues bounded above and below for any contamination, the level of contamination must be less or equal to $1/3$ and we get that the breakdown $\varepsilon^{\ast}=1/3,$ which coincide with that of Tukey’s median. The reminiscence of Tukey’s median is even emphasized since we can use a similar reasoning to that of Chen and Tyler (2002) to get a bound for the asymptotic maximum bias curve. And the bound found let us confirm that it agrees with that of the bounds obtained in the concentration inequalities given in the former Section, that is
Theorem 2
$B\left(\hat{\Gamma},\varepsilon,P_{0}\right)=\max\left\{\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\sqrt{\beta}}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}.$
All the proofs are deferred to the Appendix.
5 Appendix
Let $g\left(\delta\right)=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ with $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$Let $\varepsilon_{0}$ be the fixed point of $g$; i.e.,
$\varepsilon_{0}=g(\varepsilon_{0}).$ (6)
Let $P_{n}=(1-\varepsilon)\Phi\left(\cdot\right)+\varepsilon H_{n}\left(\cdot\right)$ be a sequence of contaminated distributions in the $\varepsilon-$contamination neighborhood $\mathcal{P}_{\varepsilon}\left(\Phi\right)$, $D_{n}$ the depth induced by these contaminated distributions and $\left\{\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right\}_{n}$ the deepest estimators associated to those distributions. Take the sets determined by $\left(\mu,\sigma\right)\in\mathbb{R}\times(0,\infty),$ $A^{(1)}\left(\mu,\sigma\right)=(-\infty,\mu-\sigma],$ $A^{(2)}\left(\mu,\sigma\right)=[\mu-\sigma,\mu],$ $A^{(3)}\left(\mu,\sigma\right)=[\mu,\mu+\sigma]$ and $A^{(4)}\left(\mu,\sigma\right)=[\mu+\sigma,\infty)$. Let $P_{\varepsilon}=(1-\varepsilon)\Phi$. Note that
$\forall n\geq 1:\,P_{n}(A^{(j)}(\mu,\sigma))\geq P_{\varepsilon}(A^{(j)}),\;j=1,2,3,4.$
Lemma 3
The following statements hold.
  • (i)
    Given $y_{0}<0$ fixed, the maximum of the function $h_{0}(x)=h\left(x,y_{0}\right),\;x>0$ occurs at $M\left(y_{0}\right)=\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$.
  • (ii)
    For $\epsilon\in(0,1/3)$ the following equivalence holds: $\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$ if and only if $\varepsilon<\varepsilon_{0}$.
Proof. (i) Fixed $y_{0},$ we want to maximize $h_{0}\left(x\right)=h\left(x,y_{0}\right)=\Phi\left(x\right)-\Phi\left(\frac{y_{0}+x}{2}\right).$ If $\varphi$ stands for the derivative of $\Phi,$ the derivative of $h_{0}$turns out to be $h_{0}^{\prime}\left(x\right)=\varphi\left(x\right)-\frac{1}{2}\varphi\left(\frac{y_{0}+x}{2}\right)\geq 0$ if and only if $2exp\left(-\frac{1}{2}x^{2}\right)\geq exp\left(-\frac{1}{8}\left(\left(x+y_{0}\right)^{2}\right)\right)$ if and only if $\ln 2-\frac{3x^{2}}{8}+\frac{xy_{0}}{4}+\frac{y_{0}^{2}}{8}\geq 0$ if and only if $\frac{1}{3}\left(y_{0}-2\sqrt{y_{0}^{2}+6\ln 2}\right)$leq $x$leq $\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$. Then $M\left(y_{0}\right)=\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$ yields the maximum value given by
$h_{0}\left(M\left(y_{0}\right)\right)=\Phi\left(\frac{y_{0}}{3}+\frac{2}{3}\sqrt{y_{0}^{2}+6\ln 2}\right)-\frac{1}{2}\Phi\left(\frac{2}{3}y_{0}+\frac{1}{3}\sqrt{y_{0}^{2}+6\ln 2}\right).$
(ii) If $\varepsilon<\varepsilon_{0}$, using (6), we have that
$(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)>(1-\varepsilon_{0})h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=g(\varepsilon_{0})=\varepsilon_{0}>\varepsilon.$
Assume now that $\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$ holds. Using again (6), we have that
$\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}(1-\varepsilon_{0}))h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}\varepsilon_{0}.$
and the converse follows.
Lemma 4
The function $g\left(\delta\right)=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ is strictly decreasing for $\delta\in\left(0,1/3\right).$ Therefore, $\delta<g\left(\delta\right)$ if and only if $\delta<\varepsilon_{0}.$
Proof. Note that $g(\delta)=(1-\delta)f(y_{\delta})=f\left(\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)\right)$with
$f\left(y\right)$ $=$ $h\left(\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2},y\right)$
$=$ $\Phi\left(\frac{1}{3}y+\frac{2}{3}\sqrt{y^{2}+6\ln 2}\right)-\Phi\left(\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}\right).$
Let us see that the function $f$ is decreasing in $y\in(-\infty,0).$ Note that
$\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2}=\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}+\frac{1}{3}\left(\sqrt{y^{2}+6\ln 2}-y\right)$
and $\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2}>0$ for all $y.$ The derivative of this function turns out to be
$f^{\prime}\left(y\right)$ $=$ $\varphi\left(\frac{1}{3}y+\frac{2}{3}\sqrt{y^{2}+6\ln 2}\right)\left(\frac{2}{3}+\frac{1}{3}\frac{y}{\sqrt{y^{2}+6\ln 2}}+\frac{1}{3}\left(\frac{y}{\sqrt{y^{2}+6\ln 2}}-1\right)\right)$
$-\varphi\left(\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}\right)\left(\frac{2}{3}+\frac{1}{3}\frac{y}{\sqrt{y^{2}+6\ln 2}}\right).$
Set $a=\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2},$ $b=\frac{1}{3}\left(\sqrt{y^{2}+6\ln 2}-y\right)>0$ and
$f^{\prime}\left(y\right)=\frac{1}{\sqrt{y^{2}+6\ln 2}}\left(\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)\right).$
Observe that $\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}>0$ if and only if $y>-\sqrt{\ln\left(4\right)}.$ Therefore, if $y\leq-\sqrt{\ln\left(4\right)},$ then $a\leq 0,$ $\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)<0$ and
$f^{\prime}\left(y\right)=\frac{1}{\sqrt{y^{2}+6\ln 2}}\left[\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)\right]<0$
If $y>-\sqrt{\ln\left(4\right),}$then $a>0,$ $a<a+b,$ $\varphi\left(a+b\right)<\varphi\left(a\right)$ and $f^{\prime}\left(y\right)<0,$ which entails the result. Since the function $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ is increasing in $\delta$ then the composition $f\left(y_{\delta}\right)=h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ is decreasing in $\delta$. Consequently, $g\left(y\right)$ is a product of two positive decreasing functions and it is decresing in $\delta$. Therefore, the claim of the Lemma follows.
Lemma 5
The following statements are valid.
  • (i)
    If $\varepsilon<1/5$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
    $\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\geq\varepsilon\text{.}$
  • (ii)
    If there exists $0<\alpha\leq 1$ verifying that $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))\geq\alpha$ and $\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}<\varepsilon<\varepsilon_{0},$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
    $\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon\alpha\text{.}$
[…]
  • (iii)
    If $P_{\varepsilon}\left(A^{(j)}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)\underset{n\rightarrow\infty}{\rightarrow}0$ for some $j\in\left\{1,2,3,4\right\}$ then $\underset{n\rightarrow\infty}{\overline{\lim}}D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)$leq $\varepsilon$ and $\varepsilon>1/5.$
  • (iv)
    Given a sequence $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ such that $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\mu_{n},\sigma_{n}\right)>0$ then $\underset{n\rightarrow\infty}{\underline{\lim}}P_{n}\left(A^{(j)}\left(\mu_{n},\sigma_{n}\right)\right)>0$ for any $j\in\left\{1,2,3,4\right\}.$
Proof. (i) It suffices to take $x_{U}=-x_{L}$ such that $\Phi\left(x_{L}\right)=1/4.$ Indeed, considering $\mu=(x_{L}+x_{U})/2$ and $\sigma=(x_{L}-x_{U})/2$, the probabilities that define the notion of depth satisfy: $P_{n}(A^{(1)}\left(\mu,\sigma)\right)\geq P_{\varepsilon}(A^{(1)}(\mu,\sigma))=(1-\epsilon)\Phi(x_{L})=(1-\epsilon)\frac{1}{4}$, $P_{n}(A^{(2)}(\mu,\sigma)))\geq P_{\varepsilon}(A^{(2)}(\mu,\sigma))=(1-\epsilon)[\Phi(0)-\Phi(x_{L})]=(1-\epsilon)\frac{1}{4}$, and similarly for the other terms. Therefore $D\left(0,-x_{L}\right)\geq\left(1-\varepsilon\right)1/4>\varepsilon$ and the statement follows.
(ii) Given $y_{0}=y_{\varepsilon_{0}},$ and using Lemma 3 we can write $\ (1-\epsilon)h_{0}(x)$ $=$ $P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+x}{2},\frac{x-y_{\varepsilon_{0}}}{2}\right)\right)$.The maximum of $h_{0}$ occurs at
$M\left(y_{\varepsilon_{0}}\right)=\frac{1}{3}\left(y_{\varepsilon_{0}}+2\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right)$
and the maximum value is given by
$h_{0}\left(M\left(y_{\varepsilon_{0}}\right)\right)=\Phi\left(\frac{y_{\varepsilon_{0}}}{3}+\frac{2}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right)-\frac{1}{2}\Phi\left(\frac{2}{3}y_{\varepsilon_{0}}+\frac{1}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right).$
If $\varepsilon<\varepsilon_{0}$, by (ii) of Lemma 3
$P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)=(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)>\varepsilon$
Moreover
$P_{\varepsilon}\left(A^{(1)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)=\varepsilon_{0}>\varepsilon,$
and then, $P_{n}\left(A^{(j)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)>\varepsilon$, $j=1,3$. Observe that $\frac{2}{3}y_{\delta}+\frac{1}{3}\sqrt{y_{\delta}^{2}+6\ln 2}>0$ if and only if $y_{\delta}\in\left[-\ln\left(4\right),\ln\left(4\right)\right]$ if and only if $\delta\geq\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}.$ Therefore $\frac{2}{3}y_{\varepsilon_{0}}+\frac{1}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}>0,$and we can say that
$P_{\varepsilon}\left(A^{(2)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$ $>$
$P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$ $>$ $\varepsilon,$
and then $P_{n}\left(A^{(2)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)>\epsilon$.
The remaining set $A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)$ has probability greater than $\varepsilon\alpha.$ Set
$\eta=\frac{(1-\varepsilon)}{2}\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right),$
then we get, for $n$ large enough that
$(1-\varepsilon)\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$
$+\varepsilon H_{n}\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$
$\geq$ $(1-\varepsilon)\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)+\varepsilon\alpha-\eta=\varepsilon\alpha+\eta,$
and (ii) follows.
(iii) Take $-x_{L}$ obtained in item (i) and $j_{0}$ such that $\lim_{n}P_{\varepsilon}\left(A^{(j_{0})}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)=0.$ Then,
$\frac{(1-\varepsilon)}{4}\leq D\left(0,-x_{L}\right)\leq D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\leq P_{\varepsilon}\left(A^{(j_{0})}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)+\varepsilon$
and the result follows.
(iv) It follows straightforwardly.
Lemma 6
If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ then there exists $-\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon$
if and only if $\varepsilon<\varepsilon_{0}.$
Proof. If $\varepsilon<\varepsilon_{0}$ then Lemma (5) (i) and (ii) says the conclusion. If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ and there exists $-\infty<x_{L}<x_{U}<\infty$ such that $\varepsilon$ $<$  $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)$ then
$P_{\varepsilon}\left(A^{(j)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)>\varepsilon\text{ forall }j\in\left\{1,2,3\right\}.$
If $P_{\varepsilon}\left(A^{(1)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)=(1-\epsilon)\Phi(x_{L})>\varepsilon$ then $x_{L}>\Phi^{-1}(\epsilon/(1-\epsilon))=y_{\epsilon}$. Moreover, by (i) of Lemma 1
$\varepsilon<P_{\varepsilon}\left(A^{(3)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)$ $=$ $(1-\epsilon)\left(\Phi(x_{U})-(\Phi(\frac{x_{L}+x_{U}}{2})\right)$
$=$ $(1-\epsilon)h(x_{L},x_{U})$
$\leq$ $(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$
$=$ $g(\epsilon).$
By Lemma 4 we get that $\varepsilon<\varepsilon_{0}.$
Lemma 7
Given the contaminated distributions $P_{n}=(1-\varepsilon)\Phi\left(\cdot\right)+\varepsilon\delta_{a_{n}}\left(\cdot\right)$ with point mass contaminations, $\delta_{a_{n}}$ verifying that $\lim_{n\rightarrow\infty}a_{n}=\infty,$ then there exists $\left(M_{0},s_{0}\right)\in\left(0,\infty\right)\times\left(0,\infty\right)$ such that the deepest estimators $\left\{\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right\}_{n\in\mathbb{N}}\in\left(-M_{0},M_{0}\right)\times\left(s_{0},M_{0}\right)$ if and only $\varepsilon<\varepsilon_{0}.$
Proof: Take $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ any sequence such that
$\lim_{n\rightarrow\infty}\mu_{n}+\sigma_{n}=\infty\text{ and}\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\mu_{n},\sigma_{n}\right)>0.$ (7)
Then,
(i) If $\lim_{n\rightarrow\infty}\mu_{n}-\sigma_{n}=-\infty$ then $\lim_{n\rightarrow\infty}P_{n}\left(A^{(1)}(\mu_{n},\sigma_{n})\right)=0$ and Lemma 5 (iv) yields a contradiction.
(ii) If $\mu_{n}+\sigma_{n}>a_{n}$ for $a_{n}$ large enough, then $\lim_{a\rightarrow\infty}P_{n}(A^{(1)}(\mu_{n},\sigma_{n}))=0$ and Lemma 5 (iv) yields a contradiction. Consequently,
$-\infty<\inf_{n}\left(\mu_{n}-\sigma_{n}\right)\text{ and }\mu_{n}+\sigma_{n}\leq a_{n}\text{ for }n\text{ sufficiently large.}$.
Set $b=\lim_{n\rightarrow\infty}\mu_{n}-\sigma_{n}.$ This allows to say that $\lim_{n\rightarrow\infty}\mu_{n}=\infty$ since $\lim_{n\rightarrow\infty}\mu_{n}+\sigma_{n}-\left(\mu_{n}-\sigma_{n}\right)=\infty.$
(iii) If $\mu_{n}+\sigma_{n}<a_{n}$ for $n$ sufficiently large then $\lim_{a\rightarrow\infty}P_{n}\left(A^{(3)}(\mu_{n},\sigma_{n})\right)=0,$ and Lemma 5 (iv) yields a contradiction. Then $\mu_{n}+\sigma_{n}=a_{n}$ for $n$ sufficiently large.
(iv) Summing up, we can assume without loss of generality that
$\lim_{n\rightarrow\infty}P_{n}\left(A^{(1)}(\mu_{n},\sigma_{n})\right)$ $=$ $\left(1-\varepsilon\right)\Phi\left(b\right),\text{ \ \ }$
$\lim_{a\rightarrow\infty}P_{n}\left(A^{(2)}(\mu_{n},\sigma_{n})\right)$ $=$ $\left(1-\varepsilon\right)\left(1-\Phi\left(b\right)\right),\text{ \ \ }$
$\lim_{n\rightarrow\infty}P_{n}\left(A^{(3)}(\mu_{n},\sigma_{n})\right)$ $=$ $\varepsilon,\text{ \ \ }\lim_{a\rightarrow\infty}P_{n}\left(A^{(4)}(\mu_{n},\sigma_{n})\right)=\varepsilon.$
By Lemma 6 there exists $-\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon$
if and only if $\varepsilon<\varepsilon_{0}.$ Therefore, for $n$ large enough,
$D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>D_{n}\left(\mu_{n},\sigma_{n}\right)$ (8)
if and only if $\varepsilon<\varepsilon_{0}.$ Consequently, in case of having of point mass contaminations, $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ cannot be the deepest estimators if and only if $\varepsilon<\varepsilon_{0}.$
Theorem 8
The breakdown point of the depth estimator is $\varepsilon^{\ast}=\varepsilon_{0}.$
Proof. Given $\varepsilon>0,$
(i)  Assume that the deepest estimator remains bounded in the $\varepsilon-$contamination neighborhood, that is there exists $M>0$ and $r>0$ such that
$\left(\mu\left(P\right),\sigma\left(P\right)\right)\in\left[-M,M\right]\times\left[r,M\right]$
for all $P\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$. An special case is given by the point mass contaminations. Thus Lemma 7  let us infer that $\varepsilon<\varepsilon_{0}$ and the breakdown point verifies that $\varepsilon^{\ast}\leq\varepsilon_{0}.$
(ii) Assume that there is a sequence of distributions $P_{n}\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$ such that the associated deepest estimators $\left\{\left(\hat{\mu}_{n}\hat{\sigma}_{n}\right)\right\}$ verifies that $\lim_{n\rightarrow\infty}\left(\hat{\mu}_{n}+\hat{\sigma}_{n}\right)=\infty.$ Set $A_{n}^{(j)}=A^{(j)}(\hat{\mu}_{n},\hat{\sigma}_{n}).$ Then $\lim_{n\rightarrow\infty}P_{\varepsilon}\left(A_{n}^{(j)}\right)=0$ for $j=4$ and Lemma 5 (iii) says that $\varepsilon>1/5.$ Without loss of generality we can assume that there is a -$\infty\leq b$ such that
$\lim_{n\rightarrow\infty}P_{n}\left(A_{n}^{(1)})\right)$ $=$ $\left(1-\varepsilon\right)\Phi\left(b\right)+\varepsilon\alpha_{1},\text{ }$
$\lim_{a\rightarrow\infty}P_{n}\left(A_{n}^{(2)}\right)$ $=$ $\left(1-\varepsilon\right)\left(1-\Phi\left(b\right)\right)+\varepsilon\alpha_{2},\text{ \ \ }$
$\lim_{n\rightarrow\infty}P_{n}\left(A_{n}^{(3)}\right)$ $=$ $\varepsilon\alpha_{3},\text{ \ \ }\lim_{a\rightarrow\infty}P_{n}\left(A_{n}^{(4)}\right)=\varepsilon\alpha_{4}.$
Then
$\lim_{n\rightarrow\infty}D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\leq\varepsilon\alpha,\text{ with }\alpha=\min\left(\alpha_{3},\alpha_{4}\right).$ (9)
Suppose by contradiction that $\varepsilon<\varepsilon_{0}.$ By Lemma 5 (ii), there exists $-\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon\alpha$ (10)
Then, by (9) and (10), for $n$ large enough we have
$D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)$
which is a contradiction. Therefore $\varepsilon\geq\varepsilon_{0}$ and $\varepsilon^{\ast}\geq\varepsilon_{0}$ and the result follows.
Lemma 9
Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$ it holds that
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\boldsymbol{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq\tilde{C}\left(\max\left\{\frac{p}{n},B_{T}^{2}\left(\varepsilon,I\right)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
Proof. Chen et al. (2018) in the proof of Theorem 2.1 showed that, for $\mathbf{\theta}=\mathbf{0}$
$\Phi\left(\left\|\hat{\theta}\right\|\right)$$\leq$ $\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$
$=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$
The upper bound becomes useless for $\varepsilon\geq 1/3$ since it is greater than 1. Therefore, we need to consider $\varepsilon<1/3-c.$ Put $b\left(p,n\right)=40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$ and $a\left(\varepsilon\right)=\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}=\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}.$ Thus, (5) is equivalent to
$\left\|\hat{\theta}\right\|$$\leq$ $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)-\Phi^{-1}\left(\frac{1}{2}\right)+\Phi^{-1}\left(\frac{1}{2}\right)$
$=$ $\frac{1}{\phi\left(\Phi^{-1}\left(\gamma_{C}\right)\right)}\left(\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)\leq C\left(\sqrt{\frac{p}{n}}\vee\varepsilon+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),\text{ }$
with $\text{with }\gamma_{C}$$\in$ $\left(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right).$
Therefore, set and we get
$\left\|\hat{\theta}\right\|$$\leq$ $\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)-\Phi^{-1}\left(a\left(\varepsilon\right)\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
$=$ $\frac{1}{\phi\left(\Phi^{-1}\left(\eta_{B}\right)\right)}b\left(p,n\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right),\text{ \ }\eta_{B}\in\left(a\left(\varepsilon\right),a\left(\varepsilon\right)+b\left(p,n\right)\right)$
$\leq$ $\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)\right)}b\left(p,n\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
If $\varepsilon<1/3-c,$ $c$ a positive constant, $a\left(\varepsilon\right)$ is increasing on $[0,1/3-c)$, let $a(m_{c})$ be the maximum value. Take $d$ such that $01-a\left(m_{c}\right)>d>$ 0 and $\left(p,n\right)$in $A_{c,d}$ $=$ $\left\{\left(p,n\right):b\left(p,n\right)<1-a\left(m_{c}\right)-d\right\}.$Then, there exists a constant $C_{c,d}=\sup_{\left(p,n\right)\in A_{c,d}}$ $\left[\phi\left(\Phi^{-1}\left(a\left(m_{c}\right)+b\left(p,n\right)\right)\right)\right]^{-1}$ such that
$\left\|\hat{\theta}\right\|\leq\tilde{C}_{c,d}\left(\sqrt{\frac{p}{n}}\vee B_{L}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,I\right)+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),$
and the lemma follows.
Lemma 10
Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\leq 1/3-c$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
Proof. In the process of obtaining (5), In the process of obtaining such an inequality, Chen et al. (2018, p. 1955) showed that, with probability greater than $1-2\delta,$
$\sup_{\mathbf{u\in}\emph{S}^{p-1}}\left|\Phi\left(\sqrt{\beta}\right)-\Phi\left(\frac{\mathbf{u}^{t}\hat{\Gamma}\mathbf{u}}{\mathbf{u}^{t}\Sigma\mathbf{u}}\right)\right|\leq\frac{\varepsilon}{2(1-\varepsilon)}+\sqrt{\frac{2400e\pi}{1-e^{-1}}}\sqrt{\frac{3+2p}{n}}+\frac{7}{4}\sqrt{\frac{\log\left(1/\delta\right)}{n}}.$ (13)
holds. Put $a_{n,\delta}=40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{3+2p}{n}}+\frac{7}{2}\sqrt{\frac{\log(1/\delta)}{n}}$and $\sqrt{\beta}\sqrt{\frac{\mathbf{u}^{T}\left(\hat{\Gamma}/\beta\right)u}{\mathbf{u}^{T}\Sigma\mathbf{u}}}=\sqrt{\beta}x_{\mathbf{u}}$, with $a_{n,\delta}\geq 0$. Then,
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|$ (14)
$=$ $\max\left(\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}),\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)\right)$ (15)
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$, $\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right).$If $1\leq\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ then $\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}).$ If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq 1\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ we have to analyze both cases. If the maximum on the righthand side of (14) occurs in $\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta})$ then $\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\geq 1,$ and we have,
$\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)\leq\Phi(\sqrt{\beta})+\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}.$
By calling $a\left(\varepsilon\right)=\frac{3}{4}+\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}$ an increasing function in $\varepsilon$, we have
$\sup_{u\in S^{p-1}}x_{u}$$\leq$ $\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)+1$
$+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]$
Then there exists $\eta\in\left[a\left(\varepsilon\right),a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right]$ such that
$\sup_{u\in S^{p-1}}x_{u}-1\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(\eta\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right],\text{ }$
00 0 $\leq$ $\sup_{u\in S^{p-1}}x_{u}-1\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]=g_{S}\left(\varepsilon,n,\delta,p\right)$
Then,
$B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right],\text{ }$
If the maximum in (14) occurs at $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right),$ then $\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$ and we call $b\left(\varepsilon\right)=\frac{3}{4}-\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)},$ which is decreasing in $\varepsilon.$ Thus, we get that
00 0 $\leq$ $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)\leq\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}$
$\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)$$\geq$ $\Phi(\sqrt{\beta})-\frac{\varepsilon}{2\left(1-\varepsilon\right)}-\frac{1}{2}a_{n,\delta}$
$\inf_{u\in S^{p-1}}x_{u}$$\geq$ $\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1+1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)$
$+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)-\frac{1}{2}a_{n,\delta}\right)$
Then, there exists $\eta\in\left(b\left(\varepsilon\right)-\frac{1}{2}a_{n,\delta},b\left(\varepsilon\right)\right)$ such that
$\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}-1$$\geq$ $-\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(\eta\right)\right)}a_{n,\delta}+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1,\text{ }$
$\geq$ $-\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(b\left(\varepsilon\right)\right)\right)}a_{n,\delta}+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1.$
Then,
$0\leq 1-\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(b\left(\varepsilon\right)\right)\right)}a_{n,\delta}+1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)=g_{I}\left(\varepsilon,n,\delta,p\right).$
Call
$B_{I}\left(\varepsilon\right)=1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right).$
It is easy to show that $B_{I}\left(\varepsilon\right)\leq B_{E}\left(\varepsilon\right).$ Observe that there is an $\eta\in\left[\beta\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}},\beta\right]\cup\left[\beta,\beta\sup_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\right]$ such that
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta x_{\mathbf{u}}}\right)\right|$ $=$ $\sup_{\mathbf{u}\in S^{p-1}}\left|\left(\beta-\beta x_{\mathbf{u}}\right)\frac{\varphi\left(\eta\right)}{\sqrt{\eta}}\right|$
$\geq$ $\frac{\varphi\left(1+g_{S}\left(\varepsilon,n,p,\delta\right)\right)}{\sqrt{1+g_{S}\left(\varepsilon,n,p,\delta\right)}}\sup_{\mathbf{u}\in S^{p-1}}\left|\left(\beta-\beta x_{\mathbf{u}}\right)\right|.$
Then, we can conclude that
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(F_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
5.1 Maximum bias of the deepest scatter estimator
We know that $\hat{\Gamma}\left(P_{0}\right)$ is Fisher consistent up to a constant, that is $\hat{\Gamma}\left(P_{0}\right)=\sqrt{\beta}\Sigma.$ Without loss of generality to study the breakdown point and maximum bias, we can assume that $\Sigma=I.$ If $F$ is a centrosymmetric distribution in $\mathbb{R}^{p}$ and $X\sim F\,$with then $\mathbf{w}^{t}X\sim\mathbf{z}^{t}X$ for all $\mathbf{w,z\in\mathcal{S}}^{p-1}.$ Let us take a symmetric positive matrix $\Gamma,$with eigenvalues $l_{1}\geq l_{2}\geq\dots$ $\geq l_{p}>0$ and eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{p}$ respectively
$\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}.$
Observe that
$P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)=P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)$
$P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(\mathbf{u}^{t}X\leq-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+P\left(\mathbf{u}^{t}X\geq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$
Let $g$ be the function,
$\dot{g}$ ::: $\emph{S}^{p-1}\rightarrow\left[0,1\right],$
$g\left(\mathbf{u}\right)$ $=$ $P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right).$
Thus, we have the following result.
Lemma 11
$\ D\left(\Gamma,P_{0}\right)=\min_{\mathbf{u\in}\emph{S}^{p-1}}\left(g\left(\mathbf{u}\right),1-g\left(\mathbf{u}\right)\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$
Proof. Given $g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$ the Lagrangian is
$h\left(\mathbf{u,}\lambda\right)=g\left(\mathbf{u}\right)+\lambda\left(\mathbf{u}^{t}\mathbf{u-}1\right).$
The derivatives are given by
$\mathbf{0}\mathbf{=}\frac{\partial h\left(\mathbf{u,}\lambda\right)}{\partial\mathbf{u}}=\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\left[\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}\right]$ $-\varphi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\left[-\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}\right]+2\lambda\mathbf{u}$
[…]
Therefore, we get that
$\mathbf{0}\mathbf{=}\frac{\partial h\left(\mathbf{u,\lambda}\right)}{\partial\mathbf{u}}=\left[2\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right]\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}+2\lambda\mathbf{u}$
Call $b\left(\mathbf{u}\right)=-\left[2\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right]/\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}<0,$ which yields $b\left(\mathbf{u}\right)\Gamma\mathbf{u}\mathbf{=}2\lambda\mathbf{u}$ with $\lambda=\frac{1}{2}b\left(\mathbf{u}\right)\mathbf{u}^{t}\Gamma\mathbf{u}$ and consequently, $\left(\Gamma\mathbf{-}\left[\mathbf{u}^{t}\Gamma\mathbf{u}\right]I\right)\mathbf{u}\mathbf{=}\mathbf{0.}$Thus, this entails that the critical points of the Lagrangian are the eigenvectors of $\Gamma.$Therefore, the function,
$g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)$
has the following Lagrangian critical points we have
$g\left(\mathbf{v}_{j}\right)$ $=$ $\Phi\left(l_{j}^{1/2}\right)-\Phi\left(-l_{j}^{1/2}\right)=2\Phi\left(l_{j}^{1/2}\right)-1,\text{ }j=1,...,p$
$g\left(\mathbf{v}_{p}\right)$$\leq$ $g\left(\mathbf{v}\right)\leq g\left(\mathbf{v}_{1}\right)\text{ for all }\mathbf{v\in}\emph{S}^{p-1}$
$1-g\left(\mathbf{v}_{j}\right)$ $=$ $1-\left[\Phi\left(l_{j}^{1/2}\right)-\Phi\left(-l_{j}^{1/2}\right)\right]=2\left[1-\Phi\left(l_{j}^{1/2}\right)\right],\text{ }j=1,...,p$
$1-g\left(\mathbf{v}_{p}\right)$$\geq$ $g\left(\mathbf{v}\right)\geq 1-g\left(\mathbf{v}_{1}\right)\text{ for all }\mathbf{v\in}\emph{S}^{p-1}$
and the minimum is given by $m=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right).$
Corollary 12
The deepest estimator is a multiple of the identity and its depth is $1/2$
Proof. Since $g\left(\mathbf{v}_{p}\right)\leq g\left(\mathbf{v}_{1}\right)$ and are increasing functions of $\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}$, then $g\left(\mathbf{v}_{p}\right)=1-g\left(\mathbf{v}_{1}\right)=0.5$ which says. $l_{p}^{1/2}=l_{1}^{1/2}=\Phi^{-1}\left(3/4\right).$ Then we $\Gamma=\left[\Phi^{-1}\left(3/4\right)\right]^{2}I_{p}$ ($I_{p}$ the identity matrix $p\times p).$
Corollary 13
Let $\left\{P_{\varepsilon,n}\right\}_{n=1}^{\infty}$ be such that $P_{\varepsilon,n}=\left(1-\varepsilon\right)P_{0}+\varepsilon P_{n}$ such that the associated depth estimators $\left\{\Gamma_{n}\right\}_{n=1}^{\infty}$, $\Gamma_{n}=\sum_{j=1}^{p}l_{j}^{(n)}\mathbf{v}_{j}^{(n)}\left[\mathbf{v}_{j}^{(n)}\right]$ have either their largest eigenvalue $l_{1}^{(n)}\rightarrow\infty$ or their smallest eigenvalue $l_{p}^{(n)}\rightarrow 0.$ Then $\varepsilon>1/3.$
Proof. Put
$a_{\varepsilon,n}\left(\mathbf{w}\right)$ $=$ $\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)+$
$\varepsilon P_{n}\left(\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)$
$b_{\varepsilon,n}\left(\mathbf{w}\right)$ $=$ $\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)+$
$\varepsilon P_{n}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=1}^{p}l_{j}^{{}^{(n)}-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}\right)$
Let $\mathbf{w}_{n\text{ }}$ be such that $D\left(\Gamma_{n},P_{n}\right)=\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right).$
i) If $l_{1}^{(n)}\rightarrow\infty,$ since $\min\left(a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right),b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\right)\geq\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right)$ then, either $b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ or $a_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq b_{\varepsilon,n}\left(\mathbf{v}_{1}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ which yields that
$\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)$
$+\varepsilon P_{n}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)$
$\geq$ $D\left(\Gamma_{n},P_{n}\right)$
Observe that $P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\geq l_{1}^{(n)}\right)\rightarrow 0$ and we get
$\varepsilon\geq\lim_{n\rightarrow\infty}D\left(\Gamma_{n},P_{n}\right)\geq\lim_{n\rightarrow\infty}D\left(I,P_{n}\right)\geq\left(1-\varepsilon\right)D\left(I,P_{0}\right)=\left(1-\varepsilon\right)\frac{1}{2}$
which says that $\varepsilon\geq 1/3.$
(ii) If $l_{p}^{(n)}\rightarrow 0,$since $\min\left(a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right),b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\right)\geq\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right),$ then either $b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)$ $\geq a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ or $a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ which yields that
$\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)$
$+\varepsilon P_{n}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)$
$\geq$ $D\left(\Gamma_{n},P_{n}\right)$
Thus, since $P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)\rightarrow 0$, we get
$\varepsilon\geq\lim_{n\rightarrow\infty}D\left(\Gamma_{n},P_{n}\right)\geq\lim_{n\rightarrow\infty}D\left(I,P_{n}\right)\geq\left(1-\varepsilon\right)D\left(I,P_{0}\right)=\left(1-\varepsilon\right)\frac{1}{2}$
which says that $\varepsilon\geq 1/3.$
Given a unit vector $\mathbf{e}$ and $r>0,$ take the point mass contaminations $\delta_{r}=\delta_{r\mathbf{e}}$. If we have $P_{\varepsilon,r}=\left(1-\varepsilon\right)P_{0}+\varepsilon\delta_{r}$ and $\delta$ stands for the Kronecker delta (it is $1$ if the inequality holds, 00 otherwise), we consider
$m_{r}\left(\mathbf{u}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+\varepsilon\delta\left(r^{2}\left(\mathbf{u}^{t}\mathbf{e}\right)^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ \left(1-\varepsilon\right)\left\{P\left(\mathbf{u}^{t}X\leq-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+P\left(\mathbf{u}^{t}X\geq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right\}+\varepsilon\delta\left(r^{2}\left(\mathbf{u}^{t}\mathbf{e}\right)^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}.$
Notation 14
Put
$h^{\mathbf{e}}\left(\mathbf{v}\right)$ $=$ $\mathbf{v}^{t}\Gamma\mathbf{v/}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}$
$B_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)<r^{2}\right\}$
$F_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)=r^{2}\right\}$
$A_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)>r^{2}\right\}$
and
$gb_{r}^{\mathbf{e}}=\inf_{B_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }Gb_{r}^{\mathbf{e}}=\sup_{B_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }Ga_{r}^{\mathbf{e}}=\sup_{A_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }ga_{r}^{\mathbf{e}}=\inf_{A_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right).$
If either $B_{r}^{\mathbf{e}}=\emptyset$ or $A_{r}^{\mathbf{e}}=\emptyset$ put $gb_{r}^{\mathbf{e}}=\infty$ and $Gb_{r}^{\mathbf{e}}=-\infty$ or $ga_{r}^{\mathbf{e}}=\infty$ and $Ga_{r}^{\mathbf{e}}=-\infty$
The following lemma gives us the depth of any symmetric positive matrix $\Gamma$ under point mass contaminations.
Lemma 15
It holds that
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\\ \left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\end{array}\right\}.$ (16)
Proof. if $h^{\mathbf{e}}\left(\mathbf{v}\right)<r^{2},$ then $\left(1-\varepsilon\right)g\left(\mathbf{v}\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\geq r^{2}\right)\geq\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}}$ and $\left(1-\varepsilon\right)(1-g\left(\mathbf{v}\right))+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\leq r^{2}\right)\geq\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon.$ if $h^{\mathbf{e}}\left(\mathbf{v}\right)>r^{2},$ then $\left(1-\varepsilon\right)g\left(\mathbf{v}\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\geq r^{2}\right)\geq\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon$ and $\left(1-\varepsilon\right)(1-g\left(\mathbf{v}\right))+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\leq r^{2}\right)\geq\left(1-\varepsilon\right)\left(1-Ga_{r}^{\mathbf{e}}\right)$ and the lemma follows.
Corollary 16
Take $\Gamma$ a matrix whose eigenvector $\mathbf{v}_{1}$coincides with $\mathbf{e.}$ If $r>l_{1}^{1/2}$ consider $\mathbf{v}_{0,r}\mathbf{\in}\mathcal{S}^{p-1}$ such that $\mathbf{v}_{0,r}^{t}\left(\Gamma-r^{2}\mathbf{v}_{1}\mathbf{v}_{1}^{t}\right)\mathbf{v}_{0,r}=0.$Then
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\left(1-\varepsilon\right)(1-g(\mathbf{v}_{1})),\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\left(1-\varepsilon\right)(1-g(\mathbf{v}_{1}))+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right..$
Proof. Observe that $q_{1}=\frac{l_{1}^{1/2}}{\mathbf{v}_{1}^{t}\mathbf{e}}=l_{1}^{1/2}$ and $q_{j}=\infty,$ $j=2...,p.$
If $r<l_{1}^{1/2}$ then
$B_{r}^{\mathbf{e}}=\emptyset,Gb_{r}^{\mathbf{e}}=-\infty\text{ and }gb_{r}^{\mathbf{e}}=\infty.$
$\left\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{p}\right\}\in A_{r}^{\mathbf{e}}\text{ then }Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right)\text{and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
Thus,
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\right\}.$
If $r>l_{1}^{1/2}$ then
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}\notin B_{r}^{e},\mathbf{v}_{1}\in B_{r}^{e},\text{ }Gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right).$
Take now $\mathbf{v}_{0,r}\mathbf{\in}\mathcal{S}^{p-1}$ such that $\mathbf{v}_{0,r}^{t}\left(\Gamma-r^{2}\mathbf{v}_{1}\mathbf{v}_{1}^{t}\right)\mathbf{v}_{0,r}=0.$ Since $g\left(\mathbf{v}_{2}\right)\leq g\left(\mathbf{v}_{0,r}\right)<g\left(\mathbf{v}_{1}\right),$
$gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{0,r}\right).$
Observe that $\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}\in A_{r}^{\mathbf{e}}$ and $\mathbf{v}_{1}\notin A_{r}^{\mathbf{e}}.$ Then
$Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{0,r}\right)\text{ and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
Since
$D\left(\Gamma,P_{\varepsilon,r}\right)$ $=$
$\min\left\{\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\right\},$
then we have that
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\\ \left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\end{array}\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right..$ (17)
The following theorem restricted to matrices whose eigenvector associated with the largest eigenvalue shares the same direction than the contamination gives us which should be the deepest estimator in this class.
Lemma 17
Given $\Gamma=\sum l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$, $l_{1}\geq...\geq l_{p},$ $\mathbf{v}_{i}^{t}\mathbf{v}_{j}=\delta_{ij}$ and $\mathbf{v}_{1}=\mathbf{e.}$ Then the deepest matrices in this class are given by
$\Gamma=\left\{\begin{array}[]{cc}\sqrt{\beta}I,&\text{if }r\leq l_{1}^{-1/2},\\ \sum l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t},&\text{if }r>l_{1}^{-1/2}.\end{array}\right.$
with $\beta=\left[\Phi^{-1}\left(\frac{3-4\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},$ $l_{1}=\left[\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}=\left[\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}\leq l_{p-1}\leq\dots\leq l_{2}<\Phi^{-1}\left(\frac{3}{4}\right)$ and $D\left(\Gamma,P_{\varepsilon,r}\right)=\left(1-\varepsilon\right)/2.$
Proof. If $r<l_{1}^{1/2}$ then the deepest matrix should verify
$\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon$ $=$ $\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)$
$g\left(\mathbf{v}_{p}\right)+g\left(\mathbf{v}_{1}\right)$ $=$ $\frac{1-2\varepsilon}{1-\varepsilon}$
To increase the depth, $g\left(\mathbf{v}_{p}\right)$ should increase and $g\left(\mathbf{v}_{1}\right)$ decrease, and then they should reach $1/2$ This implies that
$g\left(\mathbf{v}_{p}\right)$ $=$ $g\left(\mathbf{v}_{1}\right)=\frac{1-2\varepsilon}{2(1-\varepsilon)}$
$l_{p}^{1/2}$ $=$ $l_{1}^{1/2}=\Phi^{-1}\left(\frac{3-4\varepsilon}{4(1-\varepsilon)}\right)$
and the deepest matrix is a multiple of the identity matrix.
Let us take $l_{1}^{1/2}<r.$ Thus we get
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\\ \left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}.$
with
$g(\mathbf{v}_{p})\leq...\leq g(\mathbf{v}_{2})<g\left(\mathbf{v}_{0,r}\right)<g\left(\mathbf{v}_{1}\right).$
Given $\Gamma$ we have for any $l_{p}\leq$ $s\leq l_{1},$ $l_{p}/r^{2}\leq$ $s/r^{2}\leq l_{1}/r^{2}<1$ and the equation $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=s=r|\mathbf{v}_{0,r}^{t}\mathbf{v}_{1}|$ has a solution $\mathbf{v}_{0,r}.$ Geometrically, $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=s=r|\mathbf{v}_{0,r}^{t}\mathbf{v}_{1}|$ means that $\mathbf{v}_{0,r}$ belongs to the intersection of an ellipsoid and a cone around $\mathbf{v}_{1}.$ $\mathbf{v}_{0,r}$ is a function of $r,$ and $\mathbf{v}_{0,r}\underset{r\rightarrow\infty}{\rightarrow}\mathbf{v}_{1}.$The best estimator should be selected so that $\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right)=(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)$ which entails that $g\left(\mathbf{v}_{0,r}\right)=1/2$ or $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=\mathbf{\Phi}^{-1}\left(3/4\right).$   Then we take matrices $\Gamma$ such that $\Phi^{-1}\left(3/4\right)\leq l_{1}^{1/2}$ (in special $s=\Phi^{-1}\left(3/4\right)=0.6745).$ All the four quantities involved in $D\left(\Gamma,P_{\varepsilon,r}\right)$ must be equal in order to get the best estimator In this case we get that
$\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon$ $=$ $\frac{\left(1-\varepsilon\right)}{2}=\left(1-\varepsilon\right)g_{p}+\varepsilon$
$\left(1-g_{1}\right)+\frac{\varepsilon}{1-\varepsilon}$ $=$ 12 ; $\frac{1}{2}\text{ \ };\text{ \ }g_{p}+\frac{\varepsilon}{1-\varepsilon}=\frac{1}{2}$
$g_{1}$ $=$ $\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}\text{ \ \ ; \ \ }g_{p}=\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}$
$g_{1}$ $=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\text{ \ \ ; \ \ }g_{p}=\frac{1-2\varepsilon}{2\left(1-\varepsilon\right)}.$
which implies that
$g_{1}$ $=$ $2\Phi\left(l_{1}^{1/2}\right)-1=\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}$
$\Phi\left(l_{1}^{1/2}\right)$ $=$ $\frac{3}{4}+\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-3\varepsilon+2\varepsilon}{4\left(1-\varepsilon\right)}=\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}.$
$g_{p}$ $=$ $2\Phi\left(l_{p}^{1/2}\right)-1=\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}$
$\Phi\left(l_{p}^{1/2}\right)$ $=$ $\frac{3}{4}-\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-3\varepsilon-2\varepsilon}{4\left(1-\varepsilon\right)}=\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}.$
Remark 18
$\Phi\left(l_{1}^{1/2}\right)$ and $\Phi\left(l_{p}^{1/2}\right)$ are the quantities predicted by the bounds in Chen et al (2018), $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)-1\right]$ and $B_{I}\left(\varepsilon\right)=1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right).$ Both functions goes to the the maximum and minimum value when $\varepsilon=1/3.$
Lemma 19
Let us consider the case of having a family of matrices $\Gamma_{r}$ with the largest eigenvalue $l_{1}^{(r)}\rightarrow$infty, $r^{2}l_{1}^{{}^{(r)}-1/2}\rightarrow 1/2,$ and the other eigenvalues going to 0. Then, $\lim_{r\rightarrow\infty}D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left(\varepsilon,1-\varepsilon\right).$
Proof. We have to analyze,
$\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\leq 1\right),\\ \left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\geq 1\right)\end{array}\right\}$
where $\mathbf{w}_{r}$ is a vector which yields the minimum.
(i) If $\ \mathbf{w}_{r}$ $\in L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)$ we get for $r$ sufficiently large,
$\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$
(ii) If $\mathbf{w}_{r}$ $\in\left[L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)\right]^{c},$ we get for $r$ sufficiently large,
$\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$
Then, for $r$ large enough, since $\varepsilon\leq 1/2,$
$D\left(\Gamma_{r},P_{\varepsilon,r}\right)=\varepsilon.$ (18)
and we conclude the statement of the lemma.
Lemma 20
The critical points of $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)$ (respectively $1-\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right))$ subject to $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\geq 0$ (idem $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\leq 0)$ are either $\mathbf{v}_{p},$ $\mathbf{v}_{1}.$ or occurs at $F_{r}^{\mathbf{e}}.$
Proof. Take the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\lambda\left(\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right),$ Karush, Kuhn, Tucker condition says that if the minimum local occurs in $\mathbf{\tilde{v},}$ then either $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0$ or we have to consider the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right).$ In the last case, it was shown that the critical points are the eigenvectors of $\Gamma.$ Suppose that $\mathbf{v}_{j_{0}\text{,}}$ $j_{0}<p,$is the eigenvector corresponding to the minimum $l_{j_{0}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$, by the continuity of the functions we could find another $\mathbf{\bar{v}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>\mathbf{\bar{v}}_{\text{ }}^{t}\Gamma\mathbf{\bar{v}}>r^{2}\left(\mathbf{\bar{v}}^{t}\mathbf{e}\right)^{2}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$  which is a contradiction. Therefore the critical points occur at $\mathbf{\tilde{v}=v}_{p}$ $\left(\mathbf{v}_{1}\right)$ or $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0.$
The following lemma will give us the maximum bias for the deepest estimator under point mass contaminations.
Lemma 21
If $\hat{\Gamma}$ stands for a deepest estimator, then .$D\left(\hat{\Gamma},P_{\varepsilon,r}\right)\leq\left(1-\varepsilon\right)/2$.
Proof. Given $\Gamma=\sum l_{v}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$.The depth of a matrix is given by, if $r>l_{1}^{1/2}.$
$\min\left\{\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\right\}\text{.}$
Call $g_{m}=\min_{\mathbf{v\in}F_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right)$ and $g_{M}=\arg\max_{\mathbf{v\in}F_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right)$. The different configurations that a matrix $\Gamma$ might have are given in the following table depending on where $\mathbf{v}_{1}$ and $\mathbf{v}_{p}$ belong to,
$\begin{array}[]{cccccccccc}A_{r}^{\mathbf{e}}&\mathbf{v}_{1},\mathbf{v}_{p}&&&\mathbf{v}_{1}&\mathbf{v}_{p}&\mathbf{v}_{1}&\mathbf{v}_{p}&&\\ B_{r}^{\mathbf{e}}&&\mathbf{v}_{1},\mathbf{v}_{p}&&\mathbf{v}_{p}&\mathbf{v}_{1}&&&\mathbf{v}_{p}&\mathbf{v}_{1}\\ F_{r}^{\mathbf{e}}&&&\mathbf{v}_{1},\mathbf{v}_{p}&&&\mathbf{v}_{p}&\mathbf{v}_{1}&\mathbf{v}_{1}&\mathbf{v}_{p}\end{array}$
(i) $\mathbf{v}_{1},\mathbf{v}_{p}\in A_{r}^{\mathbf{e}}$: the depth of $\Gamma$ should be
$\min\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{m},\left(1-\varepsilon)\left(1-g_{1}\right)\right)$
(ii) $\mathbf{v}_{1},\mathbf{v}_{p}\in B_{r}^{\mathbf{e}}:$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{m}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(iii) $\mathbf{v}_{1},\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}:$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(iv) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in B_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{m}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(v) $\mathbf{v}_{p}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{1}\in B_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{m},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(vi) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(vii) $\mathbf{v}_{1}\in B_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(viii) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(ix) $\mathbf{v}_{1},\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
In any case, we have that $D\left(\Gamma,P_{\varepsilon,r}\right)\leq\min\left\{\left(1-\varepsilon\right)g_{s},\left(1-\varepsilon\right)\left(1-g_{t}\right)\right\},s\in\left\{p,m\right\},$ $t\in\left\{1,M\right\},$ $g_{s}\leq 1-g_{t}$ and the statement follows.
Theorem 22
The asymptotic breakdown point of the deepest estimator is $1/3$
Proof. (i) Lemma above says that if the estimators moves to the boundary of the domain of the eigenvalues, then $\varepsilon\geq 1/3$  and $\varepsilon^{\ast}\geq 1/3.$
(ii) If the depth estimators $\left\{\hat{\Gamma}_{r}\right\}$ remain bounded in the $\varepsilon-$neighborhood, we have that for point mass contaminatios we have that $D\left(\Gamma_{r},P_{\varepsilon,r}\right)\leq D\left(\hat{\Gamma}_{r},P_{\varepsilon,r}\right)$ for $r$ large enough. By the preceding lemmas, we have that $\varepsilon\leq\left(1-\varepsilon\right)/2,$ and $\varepsilon\leq 1/3.$Then, if the depth estimator remains bounded in the $\varepsilon-$contamination neighborhood then the level of contamination $\varepsilon$ is less than $1/3$ Equivalently, if $\varepsilon>1/3$ then the depth estimator becomes unbounded in the $\varepsilon-$contamination neighborhood. Therefore $\varepsilon^{\ast}\leq 1/3.$
Next, we follow closely to Theorem 4.1 and 4.2 of Chen and Tyler (2002). Let us take the contaminated distribution $P_{\varepsilon,Q}=\left(1-\varepsilon\right)P_{0}+\varepsilon Q$ and $A\subset\left\{\Gamma\succeq\mathbf{0}\right\}$, definir
$\left\|A\right\|=\sup_{\Gamma\in A}\left\{\frac{\left\|\Gamma\right\|_{op}}{\left\|\beta I\right\|_{op}},\frac{\left\|\beta I\right\|_{op}}{\left\|\Gamma^{-1}\right\|_{op}}\right\}$
$L\left(\eta,P_{0}\right)$ $=$ $\left\{\Gamma\succeq 0:D\left(\Gamma,P_{0}\right)\geq D_{M}\left(P_{0}\right)-\eta\right\}$
$\Lambda\left(\varepsilon,P_{0}\right)$ $=$ $\inf_{Q}D_{M}\left(P_{\varepsilon,Q}\right)$
$\delta\left(\varepsilon,P_{0}\right)$ $=$ $\frac{\Lambda\left(\varepsilon,P_{0}\right)-\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)}{1-\varepsilon}$
$M\left(P\right)$ $=$ $\left\{\Gamma\succeq 0:D\left(\Gamma,P\right)=D_{M}\left(P\right)\right\}=\bigcap_{0<\varepsilon<D_{M}\left(P\right)}L\left(\varepsilon,P\right)$
Lemma 23
Let $\varepsilon<1/3$ and $P_{\varepsilon,Q}=\left(1-\varepsilon\right)P_{0}+\varepsilon Q.$ It holds that
  1. 1.
    $\Lambda\left(\varepsilon,P_{0}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$ and $D_{M}\left(P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)+\varepsilon.$
  2. 2.
    Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ If $\Gamma\notin L\left(\alpha.P\right)$ then $D\left(\Gamma,P_{\varepsilon,Q}\right)<\Lambda\left(\varepsilon,P_{0}\right)$ and $\Gamma$ cannot be a deepest estimator.
  3. 3.
    $B\left(\hat{\Gamma},\varepsilon,P_{0}\right)\leq\left\|L\left(\frac{\varepsilon}{1-\varepsilon},P_{0}\right)\right\|=\max\left\{\frac{1}{\beta}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\beta}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}$
Proof. Since $P_{0}=\Phi$ then $D_{M}\left(\Phi\right)=1/2$ Observe that $\delta\geq 0$ since given $Q$ any distribution on $\mathbb{R}^{p},$
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$$\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\text{ \ },\text{ \ \ }$
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$$\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)$
$D\left(\Gamma,P_{\varepsilon,Q}\right)$$\geq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)$
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$D_{M}\left(P_{\varepsilon,Q}\right)$$\geq$ $D\left(\beta I,P_{\varepsilon,Q}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$
$\Lambda\left(\varepsilon,P_{0}\right)$ $=$ $\inf_{Q}D_{M}\left(P_{\varepsilon,Q}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$
Since
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$$\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$$\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
Then,
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\leq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}+\varepsilon$
and
$D\left(\Gamma,P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon$
and (i) follows.
(ii) Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ Then $20\leq\alpha<1/$ 2 .  If $\Gamma\notin L\left(\alpha.P_{0}\right)$ we have that $D\left(\Gamma,P_{0}\right)<D_{M}\left(P_{0}\right)-\alpha.$ Therefore,
$D\left(\Gamma,P_{\varepsilon,Q}\right)$$\leq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon<\left(1-\varepsilon\right)\left[D_{M}\left(P_{0}\right)-\frac{\varepsilon}{1-\varepsilon}+\delta\left(\varepsilon,P_{0}\right)\right]+\varepsilon$
$\leq$ $\left(1-\varepsilon\right)\left[D_{M}\left(P_{0}\right)-\frac{\varepsilon}{1-\varepsilon}+\frac{\Lambda\left(\varepsilon,P_{0}\right)-\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)}{1-\varepsilon}\right]+\varepsilon=\Lambda\left(\varepsilon,P_{0}\right)$
and (ii) follows.
(iii) $L\left(\alpha,P_{0}\right)=L\left(\varepsilon/\left(1-\varepsilon\right)-\delta,P_{0}\right)\subset.L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ Then, $\Gamma\in\left(L\left(\alpha,P_{0}\right)\right)^{c}$ implies, by  (ii) that $\Gamma\in\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ or $\left(L\left(\alpha,P_{0}\right)\right)^{c}\subset\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ which says that $M\left(P_{\varepsilon,Q}\right)\subset L\left(\alpha,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ for all distribution $Q.$
Since $D\left(\Gamma,P_{0}\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$ and $D_{M}\left(P_{0}\right)=1/2,$
$L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ $=$ $\left\{\Gamma\succeq 0:\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$=$ $\left\{\Gamma\succeq 0:1-g\left(\mathbf{v}_{1}\right)\geq g\left(\mathbf{v}_{p}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$\cup\left\{\Gamma\succeq 0:g\left(\mathbf{v}_{p}\right)\geq 1-g\left(\mathbf{v}_{1}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$=$ $\left\{\Gamma\succeq 0:\frac{1}{2}+\varepsilon/\left(1-\varepsilon\right)\geq g\left(\mathbf{v}_{1}\right)\geq g\left(\mathbf{v}_{p}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}.$
Then, $\left\|\Gamma\right\|_{op}\leq\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)$ and $\left\|\Gamma^{-1}\right\|_{op}\leq 1/\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)$ if $\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P\right)$ and
$\left\|L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)\right\|$ $=$ $\sup_{\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)}\max\left\{\frac{\left\|\Gamma\right\|_{op}}{\beta},\frac{\beta}{\left\|\Gamma^{-1}\right\|_{op}}\right\}$
$=$ $\max\left\{\frac{1}{\beta}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\beta}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}$
The bias functions turn out to be
$b_{S}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\beta^{-1}\lambda_{(1)}\left(\hat{\Gamma}\left(P\right)\right)=\beta^{-1}\left\|\hat{\Gamma}\left(P\right)\right\|_{op}$
$b_{I}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{1/2}\hat{\Gamma}^{-1}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{1/2}\right)=\beta\lambda_{(1)}\left(\hat{\Gamma}^{-1}\left(P\right)\right)=\beta\left\|\hat{\Gamma}^{-1}\left(P\right)\right\|_{op}.$
$B\left(\hat{\Gamma},\varepsilon,P_{0}\right)=\max\left\{\beta^{-1}\sup_{P\in\mathcal{P}_{\varepsilon}}\left\|\hat{\Gamma}\left(P\right)\right\|_{op},\text{ }\beta\sup_{P\in\mathcal{P}_{\varepsilon}}\left\|\hat{\Gamma}^{-1}\left(P\right)\right\|_{op}\right\}$
Since $M\left(P_{\varepsilon,Q}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ then $B\left(\hat{\Gamma},\varepsilon,P_{0}\right)\leq\left\|L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)\right\|.$
Corollary 24
$B\left(\hat{\Gamma},\varepsilon,P_{0}\right)=\left\|L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)\right\|.$
Proof. Since the point mass contaminations yield estimators whose bias reach the upper bound then