### v2

Abstract
The concept of statistical depth extends the notions of the median and quantiles to other statistical models. These procedures aim to formalize the idea of identifying deeply embedded fits to a model that are less influenced by contamination. In the multivariate case, Tukey’s median was a groundbreaking concept for multivariate location estimation, and its counterpart for scatter matrices has recently attracted considerable interest. The breakdown point and the maximum asymptotic bias are key concepts used to summarize an estimator’s behavior under contamination. In the multivariate and regression setting we analyze recently introduced concentration inequalities that provide a unified framework for studying both the statistical convergence rate and robustness of Tukey’s median, depth-based scatter matrices and depth-based multivariate regression estimators. We observe that slight variations in these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. We explicitly obtain the maximum bias curve and breakdown point of the deepest scatter matrices. For the location and scale model, we consider two closely related depth formulations, whose deepest estimators display significantly different behavior in terms of breakdown point. A numerical study is performed to compare the finite sample bias performance of several robust estimators in the multivariate setting.
Keywords: Multivariate scatter depth, Breakdown Point, Minimax Bias Function
1 Introduction
The concept of statistical depth has received considerable attention as a way to extend the notions of the median and quantiles to more general statistical models. Originally introduced by Tukey (1975), depth was defined as the minimum proportion of data points lying on either side of a point, and this idea was later extended to the bivariate case. Donoho and Gasko (1992) formalized the general notion of the halfspace depth of a point $\mathbf{z}$ in a $p$-dimensional space with respect to a probability measure $P$ in which the search is for a location or center of the probability measure $P$.
The seminal paper by Huber (1964) highlighted the median as an estimator with optimal worst-case behavior under contamination (i.e., its maximum bias), making it the most robust among the class of equivariant location estimators. A key feature of the median is that it is flanked on each side by half of the data mass, helping to shield it from the effects of outliers. This robustness—being well-surrounded by data—is echoed in Tukey’s concept of the multivariate median. In this case, a point in Euclidean space is considered a median if it maximizes the minimum mass contained in any closed half-space whose boundary includes the point. In this sense, the search for a point or location center that fits deeply within the data cloud is the idea that pervades some other statistical models.
This pursuit of a point that is ”deeply inserted” into the data, to enhance robustness in univariate and multivariate location models, has been extended to other statistical frameworks. In the regression setting, Rousseeuw and Hubert (1999) defined regression depth, which measures how deeply a linear fit is embedded within the data. This is evaluated by the smallest amount of data mass in the two opposing wedges formed by the fit plane and vertical planes orthogonal to the explanatory subspace. Mizera and Müller (2004) further explored depth concepts in the context of location-scale models. In fact, the univariate regression depth treated by Rousseeuw and Hubert (1999) is a special case of the multivariate regression depth defined by Mizera (2002). A very similar but earlier definition was proposed in Bern and Eppstein (2000).
Chen, Gao and Ren (2018a) extended the idea of depth to covariance matrix estimation by introducing the concept of matrix depth. Their estimator achieves the optimal rate under Huber’s $\varepsilon$-contamination model for estimating covariance/scatter matrices with various structural assumptions, such as bandedness and sparsity. Paindaveine and Van Bever (2018) also developed halfspace depth concepts for scatter, concentration, and shape matrices. While their concept of scatter depth is similar to that of Chen, Gao and Ren (2018a) and Zhang (2002), rather than focusing on the deepest scatter matrix, they studied the properties of the depth function and its associated depth regions. Gao (2020) considered estimators that are maximizers of multivariate regression depth functions, by studying its minimax rates in the settings of $\varepsilon$-contamination models for various regression problems including nonparametric regression, sparse linear regression, reduced rank regression, etc. Nagy, Schütt and Werner (2019) discussed the relationships between halfspace depth and problems in affine and convex geometry, offering an extensive overview of various depth notions.
The concept of asymptotic maximum bias for estimators in a neighborhood has been treated in the robust literature since the pioneering paper by Huber (1964). The maximum asymptotic bias function depicts more accurately the behavior of the estimator since it captures the global behavior by quantifying how much the estimator deviates in the whole contamination neighborhood under different levels of contamination from the parameters at the central model. Among several contributions in the context of maximum bias we can mention: He and Simpson (1993) obtained a lower bound on the contamination bias of an estimator that holds for a wide class of parametric families and a second lower bound which applies to locally linear estimates yielding that such estimates cannot be bias minimax among all Fisher-consistent estimates in higher dimensions; Martin and Zamar (1989) derived asymptotically minimax bias robust M-estimates of scale for positive random variables in $\varepsilon$-contamination neighborhoods; Martin and Zamar (1993a) dealt with minimax location M-estimators under an efficiency constraint; Martin and Zamar (1993b) established expressions for the maximum asymptotic bias of M-estimates of scale over the $\varepsilon$-contamination neighborhood and they construct asymptotically minimax bias robust estimates of scale; Ghement, Ruiz and Zamar (2008) considered robust scale estimators in the nonparametric regression setting and they derived their maximum bias curves; Martin, Yohai and Zamar (1989) found minimax bias estimators for two different classes of regression estimates: M-estimates with general scale and GM-estimates in the context of known intercept and elliptical regressors; Maronna and Yohai (1993) delved into a projection estimator for regression whose maximum bias is twice the lower bound attainable for maximum bias in regression; Berrendero and Zamar (2001) calculated the maximum bias for several classes of estimators in a general setup of unknown intercept and nonelliptical regressors; Svarc, Yohai and Zamar (2002) found a family of M-estimates of regression such that they minimize the asymptotic variance at the central model subject to a bound on the maximum bias over contamination neighborhoods; Yohai and Maronna (1990) calculated the maximum bias for the minimum volume ellipsoid and Tyler scatter estimator under point mass contaminations; Maronna, Stahel and Yohai (1992) came up with a projection estimator for multivariate scatter with remarkable bias performance; Maronna and Yohai (1995) studied the bias behavior of the Stahel-Donoho estimator for scatter among a bundle of multivariate dispersion estimators, Adrover (1998) computed the asymptotic maximum bias for M-estimators for multivariate dispersion showing that Tyler scatter M-estimator minimizes the maximum bias in the $\varepsilon$-neighborhood; Chen and Tyler (2002) investigated various properties of Tukey’s median, including its influence function and maximum contamination bias; Adrover, Maronna and Yohai (2002) derived the maximum asymptotic bias for the regression depth estimator and compared its performance with that of Maronna and Yohai (1993); Zuo, Cui and Young (2004) computed the maximum bias and influence functions for projection based estimators for multivariate location; Zamar (1992) calculated the maximum bias of M-estimators for orthogonal regression over the $\varepsilon$-contamination neighborhood and obtained the M-estimator minimizing the maximum bias in such a class.
Section 2 reviews the concept of halfspace depth in various settings: multivariate location and scatter, regression and multivariate regression models. Section 3 discusses the importance of maximum asymptotic bias in understanding how the estimator behavior evolves as sample size $n$ becomes sufficiently large. By examining the concentration inequalities in Chen, Gao and Ren (2018a), we show that slight variations in the derivation of these inequalities allow us to visualize the maximum bias behavior of the deepest estimators. In particular, Section 4 derives the maximum bias function for the deepest scatter matrix estimator introduced by Chen, Gao and Ren (2018a) in the case of known location, showing that this estimator shares Tukey’s median’s asymptotic breakdown point of $1/3$. Section 5 deals with two proposals for joint location and scale depth estimation in the univariate location-scale model, comparing their differing breakdown behaviors despite conceptual similarity. In Section 6 a simulation analysis is performed including several robust proposals to compare their performance in the presence of contamination, when the location estimator is included. All proofs are deferred to the Appendix.
2 Depth in several statistical models: a review
Let us consider the multivariate location and scatter model,
$\mathbf{X=\mu}_{0}\mathbf{+}V_{0}\mathbf{u}$ (1)
with $\mathbf{\mu}_{0}\in\mathbb{R}^{p},$ $V_{0}\in\mathbb{R}^{p\times p}$ an invertible matrix, and $\mathbf{u\sim}P_{0}$ an unobservable random vector. $\mathbf{\mu}_{0}$ and $V_{0}V_{0}^{t}$ (except for a constant) are assumed to be identifiable (for instance this holds if $P_{0}$ is a centrosymmetric distribution around $\mathbf{0)}$. Tukey’s depth of a vector $\mathbf{\theta}\in\mathbb{R}^{p}$ is given by
$D_{T}\left(\mathbf{\theta},P\right)=\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}P\left(\mathbf{u}^{t}\mathbf{X}\leq\mathbf{u}^{t}\mathbf{\theta}\right),$ (2)
with $\mathcal{S}^{p-1}{\ =}\left\{\mathbf{z\in}\mathbb{R}^{p}:\left\|\mathbf{z}\right\|=1\right\}.$ Then, $\mathbf{\theta}\left(P\right)=\arg\max_{\mathbf{\theta}\in\mathbb{R}^{p}}D_{T}\left(\mathbf{\theta},P\right)$ is taken to be the deepest estimator.
In the setting of multivariate linear regression, consider the model $Y=B^{t}\mathbf{X}+\sigma Z,$ where $B\in\mathbb{R}^{p\times m}$ and $\sigma>0$ are unknown, $Y\in\mathbb{R}^{m}$ is the random response vector and $\mathbf{X}\in\mathbb{R}^{p}$ is the vector of covariables with $Z\in\mathbb{R}^{m}$ an unobservable random vector independent of $\mathbf{X}$. We assume that $B$ and $\sigma$ are identifiable (for instance, if $Z\sim F_{0}$ is centrosymmetric around $\mathbf{0).}$ The multivariate regression depth of $B\in\mathbb{R}^{p\times m}$ is defined as
$D_{MR}\left(B,P\right)=\inf_{U\in\mathbb{R}^{p\times m}-\left\{0\right\}}P\left(\left\langle U^{t}\mathbf{X},Y-B^{t}\mathbf{X}\right\rangle\geq 0\right),$ (3)
with $\left\langle\cdot,\cdot\right\rangle$ the Euclidean scalar product in $\mathbb{R}^{m}$. The deepest estimator is defined to be
$\hat{B}\left(P\right)=\arg\max_{B\in\mathbb{R}^{p\times m}}D_{MR}\left(B,P\right).$
The definition of multivariate regression depth was considered by Bern and Eppstein (2000) and Mizera (2002). When $m=1$, this is reduced to the univariate regression depth found in Rousseeuw and Hubert (1999), that is,
$D_{R}\left(\mathbf{\beta},P\right)=\inf_{\mathbf{u}\in\mathbb{R}^{p}-\left\{0\right\}}P\left(\left(\mathbf{u}^{t}\mathbf{X}\right)\left(y-\mathbf{\beta}^{t}\mathbf{X}\right)\geq 0\right).$ (4)
Interest in multivariate scatter depth has been revitalized over the past decade, particularly following the seminal papers by Chen, Gao and Ren (2018a) and Paindaveine and Van Bever (2018), which explored different aspects of the concept of depth for multivariate scatter. If we take
$\mathcal{E}=\left\{A\in\mathbb{R}^{p\times p}:A=A^{t}\text{ and }\mathbf{x}^{t}A\mathbf{x>}0\text{ for all }\mathbf{x\neq 0}\right\},$ (5)
$\mathbf{X}\sim P,$ $\mathbf{X}\in\mathbb{R}^{p},$ the depth of $\Gamma\in\mathcal{E}$ is taken to be
$D_{S}\left(\Gamma,P\right)=\inf_{\mathbf{u}\in\mathcal{S}^{p-1}}\min\left\{\begin{array}[]{l}P\left(\left|\mathbf{u}^{t}\left(\mathbf{X}-\mathbf{v}_{0}(P)\right)\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\[6.0pt] P\left(\left|\mathbf{u}^{t}\left(\mathbf{X}-\mathbf{v}_{0}(P)\right)\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$
with $\mathbf{v}_{0}$ a preliminary affine equivariant location functional used to yield an affine equivariant multivariate scatter functional. If location is known, we can assume without loss of generality that $\mathbf{v}_{0}=\mathbf{0.}$ For the known location case, the depth and the deepest estimator are defined as,
$D_{S}\left(\Gamma,P\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\},$
$\hat{\Gamma}\left(P\right)$ $=$ $\arg\underset{\Gamma\in\mathcal{E}}{\max}D_{S}\left(\Gamma,P\right),\text{ \ \ }D_{M}\left(P\right)=D_{S}\left(\hat{\Gamma},P\right).$
Given a random sample either in the multivariate model $\left\{\mathbf{X}_{i}\right\}_{i=1}^{n}$ or in the multivariate regression model $\left\{\left(\mathbf{Y}_{i},\mathbf{X}_{i}\right)_{i=1}^{n}\right\}$ and the corresponding empirical distribution function $P_{n}$ based on the sample, then we set the depth estimators as $\mathbf{\hat{\theta}}_{n}=\arg\max_{\mathbf{\theta\in}\mathbb{R}^{p}}D_{T}\left(\mathbf{\theta},P_{n}\right)\mathbf{,}$ $\hat{\Sigma}_{n}=\arg\underset{\Gamma\succeq\mathbf{0}}{\max}D_{S}\left(\Gamma,P_{n}\right)$ and $\hat{B}_{n}=\arg\underset{B\in\mathbb{R}^{p\times m}}{\max}D_{MR}\left(B,P_{n}\right).$
Our framework assumes that we have observations coming from Huber’s $\varepsilon$-contamination neighborhood in which we have a majority of observations coming from a parametric model and a minority coming from an unknown distribution. Depth estimators are supposed to be much less affected by the presence of spurious observations as it can be assessed by using robustness measures. If $\Theta$ denotes the set in which the parameters are assumed to lie in the central model, the breakdown point of an estimator is a value that quantifies the level $\varepsilon$ of contamination required to cause the estimator to move outside any compact subset of $\Theta$. The asymptotic breakdown point is an aspect of a more powerful notion to measure the performance of an estimator, the asymptotic maximum bias, whose concept is treated in the next section. In spite of the deep understanding that the maxbias function provides, it is usually neglected because of the technicalities that its derivation requires.
3 Maximum bias and concentration inequalities
3.1 Maximum bias in the multivariate location and scatter model
Given an elliptical distribution $\mathbf{X}\sim P_{0}^{E}\left(\cdot\right)=P_{\mathbf{0}}\left(V_{0}^{-1}\left(\cdot-\mathbf{\mu}_{0}\right)\right)$ at (1), with $P_{0}$ a centrosymmetric distribution around $\mathbf{0}$ the $\varepsilon$-contamination neighborhood for the multivariate model is given by
$\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)=\left\{\left(1-\varepsilon\right)P_{0}^{E}\left(\cdot\right)+\varepsilon G\left(\cdot\right),\text{ }G\text{ any distribution on }\mathbb{R}^{p}\right\},$
with $\varepsilon\in[0,1).$ Set $\Sigma_{0}=V_{0}V_{0}^{t}$ and let $\mathcal{F}$ be a subset of distributions such that $\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)\subset\mathcal{F}$ and it also contains the empirical distribution functions. Observe that, if $\mathcal{L}\left(\mathbf{X}\right)\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$ then $\mathcal{L}\left(\mathbf{X}\right)=\mathcal{L}\left(\Sigma_{0}^{-1/2}(\mathbf{\tilde{X}}-\mathbf{\mu}_{0})\right)$ with $\mathcal{L}\left(\mathbf{\tilde{X}}\right)\in\mathcal{P}_{\varepsilon}\left(P_{0}\right).$ We say that the functionals $\mathbf{\hat{\mu}}:\mathcal{F}\rightarrow\mathbb{R}^{p}$ and $\hat{\Gamma}$$\mathcal{F}\rightarrow\mathcal{E}$ ($\mathcal{E}$ as in (5)) are affine and translation equivariant if, for any invertible matrix $A\in\mathbb{R}^{p\times p},$ it holds that
$\mathbf{\hat{\mu}}\left(\mathcal{L}\left(A\mathbf{X+b}\right)\right)$ $=$ $A\mathbf{\hat{\mu}}\left(\mathcal{L}\left(\mathbf{X}\right)\right)\mathbf{+b}$
$\hat{\Gamma}\left(\mathcal{L}\left(A\mathbf{X+b}\right)\right)$ $=$ $A\hat{\Gamma}\left(\mathcal{L}\left(\mathbf{X}\right)\right)A^{t}.$
We say that the functionals $\mathbf{\hat{\mu}}$ and $\hat{\Gamma}$ are Fisher consistent if and only if $\mathbf{\hat{\mu}}\left(P_{0}^{E}\right)=\mathbf{\mu}_{0}$ and $\hat{\Gamma}\left(P_{0}^{E}\right)=c\left(V_{0}V_{0}^{t}\right),$ $c>0$ for any $\mathbf{\mu}_{0},V_{0}\in\mathbb{R}^{p\times p}$ invertible. The effect of the distortion caused by having $P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$ is customarily measured in the following invariant manner,
$b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right)=\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right)^{t}\left(\hat{\Gamma}\left(P_{0}^{E}\right)\right)^{-1}\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right),$
$b_{E}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}\left(P_{0}\right)\mathbf{u}}$
$=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)$
$=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1}\right),$
$b_{I}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\inf_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}\left(P_{0}\right)\mathbf{u}}$
$=$ $\lambda_{(p)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\lambda_{(p)}\left(\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1}\right)$
$=$ $\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}^{-1}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}^{-1}\left(P_{0}\right)\mathbf{u}}=\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{1/2}\hat{\Gamma}^{-1}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{1/2}\right).$
The second and third equalities in (3.1) and (3.1) follow from standard arguments in multivariate analysis; see for instance Section A7, p. 523 of Seber (1984). (3.1) refers to the “explosion” behavior of the functional over the neighborhood by comparing the two quadratic forms based on the functional under the true probability $P$ and $P_{0},$ respectively. Similarly, (3.1) tries to display the ”implosion” behavior of the functional over the neighborhood. Then we can define the asymptotic maximum biases for the location and scatter functionals as
$B_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right),$
$B_{E}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{E}\left(\hat{\Gamma},\varepsilon,P\right),\text{ }B_{I}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)=\sup_{P\in\mathcal{P}_{\varepsilon}}b_{I}\left(\hat{\Gamma},\varepsilon,P\right),$
$B\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\max\left\{B_{E}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right),B_{I}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)\right\}.$
We say that the asymptotic explosion and implosion breakdown points are given by
$\varepsilon_{L}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{L}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)=\infty\right\}$
$\varepsilon_{E}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{E}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)=\infty\right\},\text{ }\varepsilon_{I}^{\ast}=\inf\left\{\varepsilon>0:B_{I}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)=\infty\right\},$
$\varepsilon^{\ast}$ $=$ $\min\left(\varepsilon_{E}^{\ast},\varepsilon_{I}^{\ast}\right).$
If we consider equivariant and Fisher consistent functionals $\mathbf{\hat{\mu}}$ and $\hat{\Gamma}$ for location and scatter, it is easily proved that
$B_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}^{E}\right)$ $=$ $cB_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}\right)$
$B_{E}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $B_{E}\left(\hat{\Gamma},\varepsilon,P_{0}\right)$
$B_{I}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $B_{I}\left(\hat{\Gamma},\varepsilon,P_{0}\right),$
which entails that the maximum bias can be computed using $\mathbf{\mu}_{0}=\mathbf{0}$ and $\Sigma_{0}=I.$
3.2 Maximum bias in the multivariate regression model
We will assume in the multivariate regression model that the intercept is known and the covariates have an elliptical distribution $G_{0}$ around $\mathbf{0\in\mathbb{R}}^{p}$ with finite second moments; see Gao (2020). Given the joint cumulative distribution function
$H_{B,\sigma}\left(\mathbf{y,x}\right)=E_{\mathbf{X}}\left[F_{0}\left(\frac{\mathbf{y-}B_{0}^{t}\mathbf{w}}{\sigma}\right)I_{\pi_{\mathbf{x}}}\left(\mathbf{w}\right)\right],$ (8)
with $\pi_{\mathbf{x}}=\prod_{j=1}^{p}(-\infty,x_{j}]$, $\mathbf{x}=\left(x_{1},\dots,x_{p}\right)$, the $\varepsilon$-contamination neighborhood for the multivariate regression model is given by
$\mathcal{P}_{\varepsilon}\left(H_{B,\sigma}\right)=\left\{\left(1-\varepsilon\right)H_{B,\sigma}\left(\mathbf{y,x}\right)+\varepsilon G\left(\mathbf{y,x}\right),\text{ }G\text{ any distribution on }\mathbb{R}^{m}\times\mathbb{R}^{p}\right\},$
with $\varepsilon\in[0,1),$ Let $\mathcal{F}$ be a subset of distributions such that $\mathcal{P}_{\varepsilon}\left(H_{B,\sigma}\right)\subset\mathcal{F}$   for all $B\in\mathbb{R}^{p\times m},\sigma>0$ and $\varepsilon\leq\varepsilon^{\prime}$ and it also contains the empirical distribution functions. Observe that, if $\mathcal{L}\left(\mathbf{Y,X}\right)\in\mathcal{P}_{\varepsilon}\left(H_{B,\sigma}\right)$ then $\mathcal{L}\left(\mathbf{Y,X}\right)=\mathcal{L}\left(\mathbf{\tilde{Y}+}B^{t}\mathbf{X,X})\right)$ with $\mathcal{L}\left(\mathbf{\tilde{Y},X}\right)\in\mathcal{P}_{\varepsilon}\left(H_{0,1}\right).$ We say that the functional $T:\mathcal{F}\rightarrow\mathbb{R}^{p\times m}$ is affine, regression and scale equivariant if, for any invertible matrix $A\in\mathbb{R}^{p\times p},$ $C\in\mathbb{R}^{p\times m}$ and $s\in\mathbb{R}^{+},$ it holds that
$T\left(\mathcal{L}\left(\mathbf{Y,}A\mathbf{X}\right)\right)$ $=$ $\left(A^{t}\right)^{-1}T\left(\mathcal{L}\left(\mathbf{Y,X}\right)\right)$
$T\left(\mathcal{L}\left(\mathbf{Y+}C^{t}\mathbf{X,X}\right)\right)$ $=$ $T\left(\mathcal{L}\left(\mathbf{Y,X}\right)\right)+C$
$T\left(\mathcal{L}\left(s\mathbf{Y,X}\right)\right)$ $=$ $sT\left(\mathcal{L}\left(\mathbf{Y,X}\right)\right)$
respectively. Therefore, we may define the bias of a functional $T:\mathcal{P}_{\varepsilon}\rightarrow\mathbb{R}^{p\times m}$ at a distribution $H$ as
$b_{MR}\left(T,\varepsilon,H\right)=\left[\frac{tr\left(\left(T\left(H\right)-T\left(H_{B,\sigma}\right)\right)^{t}S\left(G\right)\left(T\left(H\right)-T\left(H_{B,\sigma}\right)\right)\right)}{\sigma^{2}}\right]^{1/2},$
where $H_{B,\sigma}$ is the distribution under the central model, $S:\mathcal{P}_{\varepsilon}\rightarrow\mathbb{R}^{p\times p}$ is an affine equivariant estimator for the dispersion matrix of $\mathbf{X}$. This definition in the multivariate regression model coincides with that of the univariate regression model; see Adrover, Maronna and Yohai (2002). Then we can define the asymptotic maximum biases for the functionals as
$B_{MR}\left(\hat{B},\varepsilon,H_{B,\sigma}\right)=\sup_{H\in\mathcal{P}_{\varepsilon}}b_{MR}\left(\hat{B},\varepsilon,H\right)$
If we consider equivariant and Fisher consistent functionals $\hat{B}$, it is easily proved that
$B_{MR}\left(\hat{B},\varepsilon,H_{B,\sigma}\right)=B_{MR}\left(\hat{B},\varepsilon,H_{0,1}\right)$
which entails that the maximum bias can be computed using $\mathbf{\sigma}=1$ and $B=0.$ If $m=1,$ the deepest estimator $\boldsymbol{\hat{\beta}}^{t}\in\mathbb{R}^{p}$ and $B_{MR}\left(\boldsymbol{\hat{\beta}},\varepsilon,H_{0,1}\right)=\sup_{H\in\mathcal{P}_{\varepsilon}}\left\|\boldsymbol{\hat{\beta}}\left(H\right)\right\|$, the definition of maximum bias coincides with the definition given in univariate regression; see, for instance, Martin, Yohai and Zamar (1989), p.1610.
3.3 Concentration inequalities for depth estimators and its relationship with maximum bias
Chen, Gao and Ren (2018a) broke new ground by introducing a unified way to study the statistical convergence rate and robustness jointly. Given $\delta\in(0,1/2),$ put $\alpha=1-2\delta$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$ be the $\varepsilon$-contamination neighborhood with $P_{0}^{E}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ Theorem 2.1 of Chen, Gao and Ren (2018a) derived that, for $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)}P\left(\left\|\boldsymbol{\hat{\theta}}_{n}-\boldsymbol{\theta}\right\|^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$ (9)
The constant $C$ in the concentration inequality (9) is actually affected by the asymptotic maximum bias of Tukey’s median. Chen and Tyler (2002) derived the asymptotic maximum bias for Tukey’s median $\mathbf{\hat{\theta}}_{T}$ for $p\geq 2,$ which turns out to be $B_{L}\left(\hat{\boldsymbol{\theta}}_{T},\varepsilon,\Phi\right)=\Phi^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$. The heuristics behind expecting the asymptotic maximum bias function to appear in the concentration inequality is as follows: As the sample size tends to infinity, the estimator is expected to converge to the functional value $\hat{\boldsymbol{\theta}}_{T}\left(P\right).$ The quantity $\left\|\hat{\boldsymbol{\theta}}_{T}\left(P\right)-\boldsymbol{\theta}\right\|$ remains within a range bounded above by the maximum bias corresponding to the given level of contamination, since the distributions vary over the entire $\varepsilon$-contamination neighborhood. For $p=1$ the rationale is completely similar, although the bound $B_{L}^{2}\left(\boldsymbol{\hat{\theta}}_{T}\mathbf{,}\varepsilon,\Phi\right)$ is too large since the maximum bias for the univariate median $\mathbf{\hat{\theta}}_{M}$ is $B_{L}^{2}\left(\mathbf{\hat{\theta}}_{M}\mathbf{,}\varepsilon\mathbf{,}\Phi\right)=\Phi^{-1}\left(1/(2(1-\varepsilon))\right)$. However, we have not been able to modify the proof to get a tighter upper bound in the concentration inequality by including $B_{L}^{2}\left(\mathbf{\hat{\theta}}_{M}\mathbf{,}\varepsilon\mathbf{,}\Phi\right)$ rather than $B_{L}^{2}\left(\hat{\boldsymbol{\theta}}_{T}\mathbf{,}\varepsilon,\Phi\right)$. Anyway, for $p\geq 2,$ the bound (9) can be derived in a more illuminating manner by explicitly incorporating the maximum bias, as the maximum bias governs the behavior of the estimator when the sample size is sufficiently large, which turns out to be a more informative inequality without enlarging significantly the bound in (9).
Lemma 3.1
Given $\delta\in(0,1/2)$ put $\alpha=1-2\delta$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$ be the $\varepsilon$-contamination neighborhood with $P_{0}^{E}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$ it holds that
$\inf_{\boldsymbol{\theta},\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)}P\left(\left\|\hat{\boldsymbol{\theta}}_{T}-\boldsymbol{\theta}\right\|^{2}\leq\tilde{C}\left(\max\left\{\frac{p}{n},B_{L}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,\Phi\right)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$ (10)
On the other hand, Theorem 3.1 of Chen, Gao and Ren (2018a) also derived a concentration inequality for the deepest estimator for the dispersion matrix. They showed that, with probability at least $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C\left(\max\left\{\frac{p}{n},\varepsilon^{2}\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$ (11)
With a similar reasoning to that of Lemma 3.1 we can obtain the following result.
Lemma 3.2
Given $\delta\in(0,1/2)$ put $\alpha=1-2\delta$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$ be the $\varepsilon$-contamination neighborhood with $P_{0}^{E}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ For $\varepsilon\leq 1/3-c$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha,$ (12)
with $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right].$
Remark 3.3
By these means, the concentration inequality is also able to uncover the likely maximum bias of the deepest one for $p\geq 2$ as it will be shown in the next Section.
Similarly, in the multivariate regressión model, under the assumptions
(a) $\mathbf{X\sim}N_{p}\left(\mathbf{0},\Sigma\right)$
(b) $\left.\mathbf{Y}\right|\mathbf{X\sim}N_{m}\left(B^{t}\mathbf{X,}\sigma^{2}I_{m}\right),$
the depth estimator $\hat{B}$ for multivariate regression verifies that there exists a universal constant $C>0$ such that $tr\left(\left(\hat{B}_{n}-B\right)^{t}\Sigma\left(\hat{B}_{n}-B\right)\right)\leq C\sigma^{2}\left(\frac{pm}{n}\vee\varepsilon^{2}\right),$ with high probability and uniformly over the $\varepsilon$-contamination neighborhood and all $B\in\mathbb{R}^{p\times m},$ (Theorem 4.1 of Gao (2020)).
We can reformulate this result similarly to those of (10) and (12). Take the functions $h:[0,\infty)\times\left[0,1\right]^{p}\rightarrow\left[0.5,1\right]$ and $g:[0,\infty)\rightarrow\left[0.5,1\right]$ defined as $h\left(t,\theta_{1},\dots,\theta_{p}\right)=\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ and $g\left(t\right)=h\left(t,1,0,...,0\right).$ Now, note that $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right):\left[0,1/3\right)\rightarrow[0,\infty)$ and put $b_{MR}^{2}\left(\varepsilon\right)=g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right).$ Then, the following result follows.
Lemma 3.4
Given $\delta\in(0,1/2),$ put $\alpha=1-2\delta$. Suppose that assumptions (a) and (b) hold. Let $\mathcal{P}_{\varepsilon}\left(H_{B,\sigma}\right)$ be the $\varepsilon$-contamination neighborhood with $H_{B,\sigma}$ as in (3.4). For $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that, with probability greater than $1-2\delta,$ it holds that
$\inf_{\begin{subarray}{c}\boldsymbol{B\in\mathbb{R}}^{p\times m},\\ P\in\mathcal{P}_{\varepsilon}\left(H_{B,\sigma}\right)\end{subarray}}P\left[tr\left(\left(\hat{B}_{n}-B\right)^{t}\Sigma\left(\hat{B}_{n}-B\right)\right)\leq C\sigma^{2}\left(\frac{pm}{n}\vee b_{MR}^{2}\left(\varepsilon\right)\right)+\frac{\log\left(1/\delta\right)}{n}\right]\geq\alpha.$
Adrover, Maronna and Yohai (2002) derived the maximum bias of depth estimators in case of the univariate regression setting with known intercept and covariables with spherical distribution. They proved that given a bivariate normal vector $\left(U,V\right)$ with zero mean, $Var\left(U\right)=Var\left(V\right)=1$ and $\rho=Corr\left(U,V\right),$ the function $h\left(\rho\right)=P\left(sg\left(U\right)=sg\left(V\right)\right)$ determines the maximum bias $b$, since it solves the equation
$\frac{b}{\sqrt{1+b^{2}}}=h^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right).$
The following lemma shows that such a $b$ solving the equation should be $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right).$
Lemma 3.5
Suppose that assumptions (a) and (b) hold. Then $B_{MR}\left(\boldsymbol{\hat{\beta}},\varepsilon,H_{0,1}\right)=g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right).$
Remark 3.6
The concentration inequality given by Gao (2020) contains the information regarding to the maximum bias.
4 Maximum bias for the deepest scatter matrix
From now onwards we will assume that
A1. $P_{0}$ is a multivariate normal around $\mathbf{0}$ and covariance matrix $I.$
The bounds found using the concentration inequalities of Chen, Gao and Ren (2018a) suggest to study the asymptotic maximum bias of the deepest scatter matrix to verify the accuracy of the bounds. We will summaryze the steps to get the maximum bias function for the deepest scatter matrix. First of all we compute the depth of any matrix under the normal model and the deepest estimator under this distribution, which is given by $\Phi^{-1}\left(3/4\right)I.$ We next verify that in case of having a sequence of contaminations in the $\varepsilon$-contamination neighborhood which yields the deepest estimator to have either the largest eigenvalue going to infinity or the smallest one going to $0,$ then the level of contamination $\varepsilon\geq 1/3,$ which entails that $\varepsilon^{\ast}\geq 1/3.$ Next, we calculate the depth of any matrix in $\mathcal{E}$ defined in (5) under point mass contaminations located along the direction given by a vector $\mathbf{e\in\mathbb{R}}^{p}$. Moreover we consider the set of the matrices whose eigenvector associated with the largest eigenvalue coincides with the vector $\mathbf{e}$, which yields a more specific formula rather than the general one for any matrix given in the previous step. Then, in this class, we calculate a deepest scatter matrix (since uniqueness cannot be suspected at all) in which the largest and smallest eigenvalues attainable for point mass contaminations coincide with those of the bounds found in the concentration inequalities. On the one hand we show that in case of having a sequence of matrices such that the maximum eigenvalue tends to infinity its depth could only converge to a value less than or equal to $\min\left(\varepsilon,1-\varepsilon\right).$ On the other hand, we prove that the deepest estimator in the case of using point mass contaminations is less than or equal to $\left(1-\varepsilon\right)/2$. Therefore we can conclude that if we have a level of contamination $\varepsilon$ such that the deepest estimator has smallest and largest eigenvalues bounded above and below for any contamination, the level of contamination must be less than or equal to $1/3$ and we get that the breakdown $\varepsilon^{\ast}=1/3,$ which coincides with that of Tukey’s median. The reminiscence of Tukey’s median is even emphasized since we can use a similar reasoning to that of Chen and Tyler (2002) to get a bound for the asymptotic maximum bias curve.
Let us state the notation and results briefly described in the previous paragraph. We know that $\hat{\Gamma}\left(P_{0}\right)$ is Fisher consistent up to a constant, that is $\hat{\Gamma}\left(P_{0}\right)=\sqrt{\beta}\Sigma.$ Without loss of generality to study the breakdown point and maximum bias, we can assume that $\Sigma=I.$ If $F$ is a centrosymmetric distribution in $\mathbb{R}^{p}$ and $\mathbf{X}\sim F\,$with then $\mathbf{w}^{t}\mathbf{X}\sim\mathbf{z}^{t}\mathbf{X}$ for all $\mathbf{w,z\in\mathcal{S}}^{p-1}.$ Let us take a symmetric positive matrix $\Gamma,$ with eigenvalues $l_{1}\geq l_{2}\geq\dots$ $\geq l_{p}>0$ and eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{p}$ respectively, $\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}.$ Observe that
$P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}\mathbf{X}\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)$
$P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $=$ $P_{0}\left(\mathbf{u}^{t}\mathbf{X}\leq-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+P_{0}\left(\mathbf{u}^{t}\mathbf{X}\geq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$
Let $g$ be the function,
$g$ $:$ $\emph{S}^{p-1}\rightarrow\left[0,1\right],$
$g\left(\mathbf{u}\right)$ $=$ $P_{0}\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}\mathbf{X}\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right).$
Then, we have that the depth under a centrosymmetric model of any positive definite matrix is given by
Lemma 4.1
$D\left(\Gamma,P_{0}\right)=\min_{\mathbf{u\in}\emph{S}^{p-1}}\left(g\left(\mathbf{u}\right),1-g\left(\mathbf{u}\right)\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right).$
Moreover, we can easily derive the deepest estimator, which confirms the Fisher-consistency of the procedure.
Corollary 4.2
The deepest estimator is a multiple of the identity matrix and its depth is $1/2.$
The following result shows that if a sequence of contaminations distributions yields a sequence of depth estimators exploding or imploding, the level of contamination must be greater than $1/3.$
Corollary 4.3
Let $\left\{P_{\varepsilon,n}\right\}_{n=1}^{\infty}$ be such that $P_{\varepsilon,n}=\left(1-\varepsilon\right)P_{0}+\varepsilon P_{n}.$ If the associated depth estimators $\left\{\Gamma_{n}\right\}_{n=1}^{\infty}$, $\Gamma_{n}=\sum_{j=1}^{p}l_{j}^{(n)}\mathbf{v}_{j}^{(n)}\left[\mathbf{v}_{j}^{(n)}\right]$ have either their largest eigenvalue $l_{1}^{(n)}\rightarrow\infty$ or their smallest eigenvalue $l_{p}^{(n)}\rightarrow 0$ then $\varepsilon>1/3.$
Given a unit vector $\mathbf{e}$ and $r>0,$ take the point mass contaminations $\delta_{r}=\delta_{r\mathbf{e}}$. If we have $P_{\varepsilon,r}=\left(1-\varepsilon\right)P_{0}+\varepsilon\delta_{r}$ and $\delta$ stands for the Kronecker delta (it is $1$ if the inequality holds, 0 otherwise), we consider the function $h^{\mathbf{e}}\left(\mathbf{v}\right)=\mathbf{v}^{t}\Gamma\mathbf{v/}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}$ and $D\left(\Gamma,P_{\varepsilon,r}\right)=\min_{\mathbf{u\in}\mathcal{S}^{p-1}}$ $m_{r}\left(\mathbf{u}\right),$ with
$m_{r}\left(\mathbf{u}\right)=\min\left\{\left(1-\varepsilon\right)g\left(\mathbf{u}\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{u}\right)\geq r^{2}\right),\left(1-\varepsilon\right)\left(1-g\left(\mathbf{u}\right)\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{u}\right)\leq r^{2}\right)\right\}.$
Notation 4.4
The function $h^{\mathbf{e}}\left(\mathbf{v}\right)$ plays a crucial role in the derivation of the maximum bias. Take the sets
$B_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)<r^{2}\right\}$
$F_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)=r^{2}\right\}$
$A_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)>r^{2}\right\},$
and the related quantities,
$gb_{r}^{\mathbf{e}}=\inf_{B_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }Gb_{r}^{\mathbf{e}}=\sup_{B_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }Ga_{r}^{\mathbf{e}}=\sup_{A_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right),\text{ \ \ }ga_{r}^{\mathbf{e}}=\inf_{A_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right).$
If either $B_{r}^{\mathbf{e}}=\emptyset$ or $A_{r}^{\mathbf{e}}=\emptyset$ put $gb_{r}^{\mathbf{e}}=\infty$ and $Gb_{r}^{\mathbf{e}}=$-$\infty$ or $ga_{r}^{\mathbf{e}}=\infty$ and $Ga_{r}^{\mathbf{e}}=-\infty.$
The following lemma gives us the depth of any symmetric positive matrix $\Gamma$ under point mass contaminations.
Lemma 4.5
It holds that
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\\ \left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\end{array}\right\}.$ (13)
From this lemma we can derive the depth for a matrix whose eigenvector associated with the largest eigenvalue coincides with the direction of the point mass contamination.
Corollary 4.6
Take $\Gamma$ to be a matrix whose eigenvector $\mathbf{v}_{1}$ coincides with $\mathbf{e.}$ If $r>l_{1}^{1/2}$ consider $\mathbf{v}_{m,r}=\arg\min_{B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$ and $\mathbf{v}_{M,r}=\arg\max_{A_{r}^{e}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v.}$ Then
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\\ \left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\end{array}\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{m,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{M,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right.$.
The following theorem restricted to matrices whose eigenvector associated with the largest eigenvalue shares the same direction as the contamination gives us which should be the deepest estimator in this class.
Lemma 4.7
Given $\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$, $l_{1}\geq...\geq l_{p},$ $\mathbf{v}_{i}^{t}\mathbf{v}_{j}=\delta_{ij}$ and $\mathbf{v}_{1}=\mathbf{e.}$ Then the deepest matrices in this class are given by
$\Gamma=\left\{\begin{array}[]{cc}\sqrt{\beta}I,&\text{if }r\leq l_{1}^{-1/2},\\ \sum l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t},&\text{if }r>l_{1}^{-1/2}.\end{array}\right$.
with $\beta=\left[\Phi^{-1}\left(\frac{3-4\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},$ $l_{1}=\left[\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}=\left[\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}=l_{p-1}=\cdots=l_{2}<\Phi^{-1}\left(\frac{3}{4}\right)$ and $D\left(\Gamma,P_{\varepsilon,r}\right)=\left(1-\varepsilon\right)/2.$
Remark 4.8
$\Phi\left(l_{1}^{1/2}\right)$ and $\Phi\left(l_{p}^{1/2}\right)$ are the quantities displayed by the bounds in Chen, Gao and Ren (2018a), $B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)-1\right]$ and $B_{I}\left(\varepsilon\right)=1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right).$
In case of having point mass contaminations going to infinity and taking matrices whose largest eigenvalue goes to infinity as well with certain rate, their depth should converge to the level of contamination as we can establish in the following result.
Lemma 4.9
Let us consider the case of having a family of matrices $\Gamma_{r}$ with the largest eigenvalue $l_{1}^{(r)}\rightarrow$ $\infty$, $r^{2}l_{1}^{{}^{(r)}-1/2}\rightarrow 1/2,$ and the other eigenvalues going to 0. Then, $\lim_{r\rightarrow\infty}D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left(\varepsilon,1-\varepsilon\right).$
The following lemma will give us a bound for the depth of the deepest estimator under point mass contaminations.
Lemma 4.10
If $\hat{\Gamma}$ stands for a deepest estimator, then $D\left(\hat{\Gamma},P_{\varepsilon,r}\right)\leq\left(1-\varepsilon\right)/2$.
Finally we can show the asymptotic breakdown point for the deepest estimator for multivariate scatter.
Theorem 4.11
The asymptotic breakdown point of the deepest estimator is $1/3.$
Next, we follow closely Theorems 4.1 and 4.2 of Chen and Tyler (2002). Let us take the contaminated distribution $P_{\varepsilon,Q}=\left(1-\varepsilon\right)P_{0}+\varepsilon Q$ and $A\subset\left\{\Gamma\succeq\mathbf{0}\right\}$, define
$\left\|A\right\|=\sup_{\Gamma\in A}\left\{\frac{\left\|\Gamma\right\|_{op}}{\left\|\beta I\right\|_{op}},\frac{\left\|\beta I\right\|_{op}}{\left\|\Gamma^{-1}\right\|_{op}}\right\}$
$L\left(\eta,P_{0}\right)$ $=$ $\left\{\Gamma\succeq 0:D\left(\Gamma,P_{0}\right)\geq D_{M}\left(P_{0}\right)-\eta\right\}$
$\Lambda\left(\varepsilon,P_{0}\right)$ $=$ $\inf_{Q}D_{M}\left(P_{\varepsilon,Q}\right)$
$\delta\left(\varepsilon,P_{0}\right)$ $=$ $\frac{\Lambda\left(\varepsilon,P_{0}\right)-\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)}{1-\varepsilon}$
$M\left(P\right)$ $=$ $\left\{\Gamma\succeq 0:D\left(\Gamma,P\right)=D_{M}\left(P\right)\right\}=\bigcap_{0<\varepsilon<D_{M}\left(P\right)}L\left(\varepsilon,P\right)$
Lemma 4.12
Let $\varepsilon<1/3$ and $P_{\varepsilon,Q}=\left(1-\varepsilon\right)P_{0}+\varepsilon Q.$ It holds that
  1. 1.
    $\Lambda\left(\varepsilon,P_{0}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$ and $D_{M}\left(P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)+\varepsilon.$
  2. 2.
    Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ If $\Gamma\notin L\left(\alpha.P\right)$ then $D\left(\Gamma,P_{\varepsilon,Q}\right)<\Lambda\left(\varepsilon,P_{0}\right)$ and $\Gamma$ cannot be a deepest estimator.
  3. 3.
    $B\left(\hat{\Gamma},\varepsilon,P_{0}\right)\leq\left\|L\left(\frac{\varepsilon}{1-\varepsilon},P_{0}\right)\right\|=\max\left\{\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\sqrt{\beta}}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}$.
Thus, we have the following result.
Theorem 4.13
$B\left(\hat{\Gamma},\varepsilon,P_{0}\right)=\max\left\{\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\sqrt{\beta}}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}.$
All proofs are deferred to the Appendix.
5 Depth as a residual smallness concept
So far we have shed light on the relationship between concentration inequalities and the concept of asymptotic maximum bias for depth estimators in the multivariate and regression setting. We next want to strengthen some connections between some equivalent concepts of depth. Carrizosa (1996) and Adrover, Maronna and Yohai (2002) independently came up with a residual smallness concept which comprises the notion of depth given in the univariate and multivariate model as well as the univariate regression model. More precisely, we can define the depth of $\boldsymbol{\theta}\in\mathbf{\mathbb{R}}^{p}$ as
$\mathcal{D}_{T}^{E}\left(\boldsymbol{\theta},P\right){\ =}\inf_{\left\|\boldsymbol{\lambda}\right\|=1,\boldsymbol{\gamma}\mathbf{\in\mathbb{R}}^{p}}P\left(\left|\boldsymbol{\lambda}^{t}(\mathbf{x-}\boldsymbol{\theta}\mathbf{)}\right|\leq\left|\boldsymbol{\lambda}^{t}(\mathbf{x-}\boldsymbol{\gamma}\mathbf{)}\right|\right)$ (14)
It is proved that $\mathcal{D}_{T}^{E}\left(\boldsymbol{\theta},P\right)=\mathcal{D}_{T}\left(\boldsymbol{\theta},P\right).$ The idea behind that definition is that the depth of a fit $\boldsymbol{\theta}$ is determined by the bad performance displayed by the residuals $\left|\boldsymbol{\lambda}^{t}(\mathbf{x-}\boldsymbol{\theta}\mathbf{)}\right|$ compared to the best competitor $\mathbf{\gamma,}$ whose residuals $\left|\boldsymbol{\lambda}^{t}(\mathbf{x-}\boldsymbol{\theta}\mathbf{)}\right|$ have the minimum probability of being worse than those of $\left|\boldsymbol{\lambda}^{t}(\mathbf{x-}\boldsymbol{\theta}\mathbf{)}\right|$, therefore the $\boldsymbol{\theta}$ with the best worst performance is singled out.
In the regression setting, (14) takes the form
$\mathcal{D}_{R}^{E}\left(\boldsymbol{\theta},P\right){=}\inf_{\boldsymbol{\gamma}\mathbf{\in\mathbb{R}}^{p}}P\left(\left|y-\boldsymbol{\theta}^{t}\mathbf{x}\right|\leq\left|y-\boldsymbol{\gamma}^{t}\mathbf{x}\right|\right)$ (15)
and the deepest regression estimator is taken to be
$\mathbf{\hat{\theta}}_{R}\left(P\right)=\arg\sup_{\mathbf{\theta}}\mathcal{D}_{R}^{E}\left(\boldsymbol{\theta},P\right).$
The equivalence between (4) and (15) in the regression setting can be found in Adrover, Maronna and Yohai (2002). In the multivariate regression model we can also generalize the concept of depth given in (15) by defining,
$\mathcal{D}_{MR}^{E}\left(B,P\right){\ =}\inf_{U\mathbf{\in\mathbb{R}}^{p\times m}}P\left(\left\|Y-B^{t}\mathbf{X}\right\|\leq\left\|Y-U^{t}\mathbf{X}\right\|\right).$
Then, it can be easily seen that this approach coincides with that of (3).
Lemma 5.1
$\mathcal{D}_{MR}^{E}\left(B,P\right)=\mathcal{D}_{MR}\left(B,P\right).$
5.1 Depth in the location-scale model
The residual smallness concept in (14) can be easily adapted for joint estimation of location and scale as follows. In the univariate case the model (1) switches to the usual location-scale model $Y=\mu_{0}+\sigma_{0}U$, $\sigma_{0}>0$. Then, the depth of $\left(\mu,\sigma\right)\in\mathbb{R}\times(0,\infty)$ is taken to be,
$D_{LS}^{1}\left(\mu,\sigma,P\right)$ $=$ $\min\left\{\begin{array}[]{c}\inf_{\lambda\in\mathbb{R}}P\left(\left[\left|Y-\mu\right|\leq\left|Y-\lambda\right|\right]\right),\\ \inf_{\gamma>0}P\left(\left[\left|\left|\frac{Y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{Y-\mu}{\gamma}\right|-1\right|\right]\right)\end{array}\right\}$ (18)
$\left(\hat{\mu}_{1},\hat{\sigma}_{1}\right)$ $=$ $\arg\max_{\mu,\sigma}D_{LS}^{1}\left(\mu,\sigma,P\right).$
With this definition of depth, we obtain very well known functionals for location and scale, as it is stated in the following lemma.
Lemma 5.2
If $Y\sim P,$ then $\hat{\mu}_{1}=med_{P}\left(Y\right)$ and $\hat{\sigma}_{1}=med_{P}\left(\left|Y-med_{P}\left(Y\right)\right|\right)$ and $D_{LS}^{1}\left(\hat{\mu}_{1},\hat{\sigma}_{1},P\right)\geq 0.5.$
[…]
A usual way to model the presence of spurious data in a sample is the $\varepsilon$-contamination neighborhood, which is defined to be,
[…]
and the asymptotic breakdown point of $\left(\varepsilon_{L}^{\ast},\varepsilon_{S}^{\ast}\right)$ of a functional $\left(\hat{\mu},\hat{\sigma}\right):\mathcal{P}_{\varepsilon}\left(F_{0}\right)\rightarrow\Theta$ is defined to be,
[…]
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}\max\left\{\hat{\sigma}\left(\cdot\right),\hat{\sigma}^{-1}\left(\cdot\right)\right\}=\infty\right\},\text{ }$
[…]
It is well known that $\hat{\mu}_{1}$ as well as $\hat{\sigma}_{1}$ are estimators whose breakdown point is $0.5,$ the best breakdown point attainable for location and scale equivariant estimators. We want to introduce now a slight modification which yields a substantial modification of the estimators. Let us suppose that we bind together the estimation of both parameters in one expression, by considering the following second proposal for location-scale depth,
$D_{LS}^{2}\left(\mu,\sigma,P\right)$
$=$ $\inf_{\gamma>0,\lambda\in\mathbb{R}}P\left(\left[\left|Y-\mu\right|\leq\left|Y-\lambda\right|\right]\cap\left[\left|\left|\frac{Y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{Y-\mu}{\gamma}\right|-1\right|\right]\right),$
and the deepest estimator as
$\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\arg\max_{\mu\in\mathbb{R},\sigma>0}D_{LS}^{2}\left(\mu,\sigma,P\right),$
which we will call the modified location-scale deepest estimator. It is easy to see that
$D_{LS}^{2}\left(\mu,\sigma,P\right)=\min\left\{\begin{array}[]{cc}P\left(\mu-\sigma\leq y\leq\mu\right),&P\left(\mu\leq y\leq\mu+\sigma\right),\\ P\left(\left[y\leq\mu-\sigma\right]\right),&P\left(\left[y\geq\mu+\sigma\right]\right).\end{array}\right\}.$
If $P=F_{0}\left(\sigma_{0}^{-1}\left(\cdot-\mu_{0}\right)\right),$ then the modified estimator verifies that $\left(\hat{\mu}_{2}(P),\hat{\sigma}_{2}(P)\right)=\left(\mu_{0},\sigma_{0}\right)$ (that is, it is Fisher-consistent). If $P$ is a symmetric distribution, $\hat{\mu}^{2}=med\left(P\right),$ $\hat{\sigma}^{2}=med\left(P\right)-q_{1}\left(P\right),$ $q_{1}\left(P\right)$ is the first quartile of $P.$ However, we can prove that $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ has breakdown point much lower than that of the $D_{LS}^{1}.$ We will assume that $F_{0}=\Phi$, the standard normal distribution. Let $\varepsilon_{0}$ be the solution to the equation
$\delta=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right).$
with $h\left(x,y\right)=\Phi\left(x\right)-\Phi\left(\frac{y+x}{2}\right)$ and $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$ Then we obtain that the breakdown point for the modified estimator goes down.
Theorem 5.3
$\varepsilon^{\ast}\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)=\varepsilon_{0}$ with $1/5<\varepsilon_{0}<1/4.$
The second proposal for depth estimators in the location-scale model yields estimators with much lower breakdown point compared to that of the first proposal. This phenomenon can be considered in a broader scenario showing a behavior also described in some other simultaneous procedures in the robust literature. While the estimation of $\mu$ and $\sigma$ is performed separately in (18), the estimation of both parameters is jointly performed in (5.1). For instance, a similar behavior is observed in the simultaneous M-estimation for location and scale (Huber and Ronchetti (2009), p.141) and in the context of simultaneous M-estimation for regression and scale in Maronna and Yohai (1991) which seems to prevent one from doing simultaneous estimation in order to improve its robustness.
6 Numerical study
In the previous section we derived the maximum bias curve for the deepest scatter matrix as well as its breakdown point. Only a few maxbias curves have been derived for multivariate dispersion measures, despite their being a very informative overall measure of robustness. Therefore, we cannot rely on comparisons among many theoretical maxbias curves to understand the behaviour of dispersion estimators.
On the one hand, maxbias curves are population measures; that is, they describe the behaviour of estimators as functionals over contamination neighborhoods. On the other hand, finite-sample effects should also be taken into account. In this section, we report the results of a Monte Carlo simulation study to investigate these effects, explicitly including the estimation of the unknown location parameter, whose impact is typically avoided in the maxbias derivation for scatter matrices because of the intractability to perform a theoretical analysis.
In Maronna and Yohai (2017) and Maronna et al. (2018), the behavior of scatter matrix estimators is studied under the assumption of known location, using the Kullback-Leibler divergence as a performance measure. Chen, Gao and Ren (2018a) addresses the performance of estimators under unknown location by using the operator norm. We consider the empirical version of the bias given in formulae (3.1) and (3.1) as performance measures. Following Maronna and Yohai (1995), we also use the condition number of an estimator as a measure of asymptotic bias.
In the following subsections, we describe the estimators and contamination scenarios, explain how the estimated bias measures are empirically computed from replicated samples of the contaminated model, and then report and discuss the main results.
6.1 The estimators
Let us consider $\mathbf{X}=(X_{1},\ldots,X_{p})^{T}$ a random vector and $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}$ a random sample from $\mathbf{X}$. Let $\mathbb{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}\subset\mathbb{R}^{p}$ denote a dataset. Consider $\mathbf{x},\boldsymbol{\mu}\in\mathbb{R}^{p}$ and $\Sigma\in\mathbb{R}^{p\times p}$, and let $d(\mathbf{x},\boldsymbol{\mu},\Sigma)=(\mathbf{x}-\boldsymbol{\mu})^{\prime}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})$ be the squared Mahalanobis distance. Let $d_{i}=d(\mathbf{x}_{i},\boldsymbol{\mu},\Sigma)$ and $\mathbf{d}(\boldsymbol{\mu},\Sigma)=(d_{1},\ldots,d_{n}),$ $i=1,\dots,n$.
We include several estimators in our simulation study. To compute them, we use functions from R packages available on the Comprehensive R Archive Network (CRAN), with default argument values in all cases. A brief description of them is as follows.
  • 1.
    The sample covariance matrix (SCOV), defined by $S_{n}=\frac{1}{n}\sum_{i=1}^{n}(\boldsymbol{x}_{i}-\bar{\mathbf{x}}_{n})(\mathbf{x}_{i}-\bar{\mathbf{x}}_{n})^{\prime}$, where $\bar{\mathbf{x}}_{n}$ is the sample mean, included to provide a benchmark for comparison with the robust estimators that follow.
  • 2.
[…]
    $\left(\hat{\boldsymbol{\mu}},\hat{\Sigma}\right)=\arg\underset{\left(\boldsymbol{t},\Sigma\right)\in C_{h,c}}{\min}\det\left(\Sigma\right).$
    $c$ is chosen to yield a consistent estimator $\hat{\Sigma}$ at the central model, and $h=[(n+p+1)/2]$ ensures the maximal breakdown point, as described in Van Aelst and Rousseeuw (2009). MVE is computed in R using the function fastmve from the package RobStatTM (Yohai et al., 2024).
  • 3.
    The minimum covariance determinant estimator (MCD), was proposed by Rousseeuw (1985)). Given the $n$ observations, this method finds $h$ observations $\mathbf{x}_{i_{1}},\ldots,\mathbf{x}_{i_{h}}$ with sample covariance matrix $S_{h}$ having the lowest determinant, with $\lfloor(n+p+1)/2\rfloor\leq h\leq n.$ $S_{h}$ and the sample mean $\bar{\mathbf{x}}_{n}$ of the $h$ observations are the MCD estimators for multivariate scatter and location. It is computed in R using the CovMcd function from the package rrcov developed by Todorov and Filzmoser (2009).
  • 4.
    S-estimators (SE) were introduced by Davies (1987). Given the constant $\delta\in(0,1),$ an M-scale $S=S(\mathbf{d}(\boldsymbol{\mu},\Sigma))$ is defined through the equation
    $\frac{1}{n}\sum_{i=1}^{n}\rho\!\left(\frac{d_{i}}{S}\right)=\delta,$
[…]
    $(\widehat{\boldsymbol{\mu}},\widetilde{\Sigma})=\arg\min_{\boldsymbol{\mu}\in\mathbb{R}^{p},\Sigma\in\mathcal{E},|\Sigma|=1}S\bigl(\mathbf{d}(\boldsymbol{\mu},\Sigma)\bigr),$
    where $\delta$ controls the breakdown point, then it is taken as $\delta=1/2$. Hence, the S-estimator for multivariate scatter $\Sigma$ is defined as $\widehat{\Sigma}=S\,\widetilde{\Sigma}$. SE is computed using the function CovSest from the package rrcov, choosing method = "bisquare", with the bisquare function.
    $\rho(t)=\begin{cases}1-(1-t)^{3}&\text{if }t\leq 1,\\ 1&\text{if }t>1.\end{cases}$
  • 5.
    S-estimators with non-monotonic weight functions (Rocke ) were introduced by Rocke (1996). Note that S-estimators satisfy the system
    $\frac{1}{n}\sum_{i=1}^{n}W\left(\frac{d_{i}}{S}\right)(\mathbf{x}_{i}-\boldsymbol{\mu})(\mathbf{x}_{i}-\boldsymbol{\mu})^{\prime}$ $=$ $\Sigma,$
    $\frac{1}{n}\sum_{i=1}^{n}W\left(\frac{d_{i}}{S}\right)(\mathbf{x}_{i}-\boldsymbol{\mu})$ $=$ $0,$
    where $W=\rho^{\prime}$ is usually called a weight function. Rocke is computed using the function CovSest from the package rrcov with option method ="rocke", that by default sets $\alpha=0.1$. This routine considers a modification of Rocke’s “biflat” weight function given by
    $W(t)=\left[1-\left(\frac{t-1}{\gamma}\right)^{2}\right]\mathrm{1}_{(1-\gamma,1+\gamma)}(t),$
    with tuning constant $\gamma=\min\!\left(1,\frac{\chi_{p,1-\alpha}^{2}}{p}-1\right)$ where $\alpha$ is chosen to control efficiency.
  • 6.
    MM-estimators (MM) were proposed by Lopuhaa (1992). Let $\left(\widehat{\boldsymbol{\mu}}_{0},\widehat{\Sigma}_{0}\right)$ be an initial high breakdown point estimator and the squared Mahalanobis distances $d_{i}^{0}=d\left(\mathbf{x}_{i},\widehat{\boldsymbol{\mu}}_{0},\widehat{\Sigma}_{0}\right)$. Consider an M-scale $S^{0}$ solving
    $\frac{1}{n}\sum_{i=1}^{n}\rho\left(\frac{d_{i}^{0}}{S}\right)=\delta.$
    In the end, consider the location and shape estimator defined as
    $(\widehat{\boldsymbol{\mu}},\widetilde{\Sigma})=\arg\min_{u\in\mathbb{R}^{p},\Sigma\in\mathcal{E},|\Sigma|=1}\sum_{i=1}^{n}\rho\left(\frac{d_{i}}{cS^{0}}\right),$
[…]
    $\rho_{\mathrm{SHR}}(d)=\frac{1}{6.494}\begin{cases}d&\text{if }d\leq 4,\\ s(d)&\text{if }4<d\leq 9,\\ 6.494&\text{if }d>9,\end{cases}$
    where
    $s(d)=3.534-1.944\,d+0.864\,d^{2}-0.104\,d^{3}+0.004\,d^{4}.$
    For more details, see Maronna and Yohai (2017).
  • 7.
    The Stahel-Donoho location scatter estimator (SD) was proposed by Stahel (1981) and Donoho (1982), weighting the observations according to a notion of outlyingness. Given $\mathbf{u}\in\mathcal{S}^{p-1}$, denote by $\mathbf{u}^{\prime}\mathbb{X}=\left\{\mathbf{u}^{\prime}\mathbf{x}_{1},\ldots,\mathbf{u}^{\prime}\mathbf{x}_{n}\right\}$. The outlyingness with respect to $\mathbb{X}$ of $\mathbf{x}\in\mathbb{R}^{p}$ along $\mathbf{u}$ is defined by
    $t(\mathbf{x},\mathbf{u})=\frac{\mathbf{x}^{\prime}\mathbf{u}-L\left(\mathbf{u}^{\prime}X\right)}{S\left(\mathbf{u}^{\prime}X\right)},$
    where $L$ and $S$ are robust univariate location and scale estimators, for example the median and the MAD respectively. Now, the outlyingness of $\mathbf{x}$ is defined by
    $t(\mathbf{x})=\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}|t(\mathbf{x},\mathbf{u})|.$
[…]
    $\widehat{\boldsymbol{\mu}}$ $=$ $\frac{1}{\sum_{i=1}^{n}w_{i1}}\sum_{i=1}^{n}w_{i1}\mathbf{x}_{i},$
    $\widehat{\Sigma}$ $=$ $\frac{1}{\sum_{i=1}^{n}w_{i2}}\sum_{i=1}^{n}w_{i2}\left(\mathbf{x}_{i}-\widehat{\boldsymbol{\mu}}\right)\left(\mathbf{x}_{i}-\widehat{\boldsymbol{\mu}}\right)^{t}$
    It is computed using the function CovSde from the rrcov package. This function uses the piecewise polynomial weight $W_{\mathrm{opt}}$ described in Maronna and Yohai (2017).
  • 8.
    The Deepest Estimator (MDepth) proposed by Chen, Gao and Ren (2018a). We use the function matrix_depth_by_descent from the R package DepthDescent. For more details see Chen, Gao and Ren (2018b).
We do not compute Tyler’s M-estimator of scatter, since we are only using routines available at the CRAN Project and the R package ICSNP takes the sample mean as multivariate location estimator by default, which makes the comparison with other robust proposals unreliable.
6.2 Contamination scenarios
Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon$-contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$, $\boldsymbol{\theta}=\mathbf{0}$ and $\Sigma=I_{p}$, the $p\times p$ identity matrix. We assume that the contaminating distribution is $G_{k}=\delta_{x_{k}}$, a point mass at $\mathbf{Z}_{k}=(k,\ldots,k)^{t}$, with $k\in\mathbb{N}\cup\{0\}$. Consider $\boldsymbol{X}^{0}=(X_{1}^{0},\ldots,X_{p}^{0})^{\prime}\sim P_{0}$ and $B\sim Ber\left(\varepsilon\right).$ If $\boldsymbol{X}^{0}$ and $B$ are independent then
$\boldsymbol{X}=(1-B)\boldsymbol{X}^{0}+B\boldsymbol{Z}_{k}\sim P_{k},$ (20)
with $P_{k}=(1-\varepsilon)P_{\mathbf{0}}\left(\cdot\right)+\varepsilon G_{k}(\cdot)\in\mathcal{P}_{\varepsilon}\left(P_{0}\right).$ For a contamination rate $\varepsilon$ and a fixed constant $k$, using (20), we generate a sample $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ from the random vector $\mathbf{X}\sim P_{k}$.
We will consider the following scenarios:
  • Contamination proportions $\varepsilon=0.1$, $0.2$.
  • $k$ $\in$ $\left\{0,1,5,10,15,20,25\right\}$.
  • Dimensions $p=2,5$, 1010, 1515.
  • Sample size: $n=10p$ (see Maronna and Yohai (2017), Table 3), $n=40p$, $n=500p$.
For each combination of $p$, $\varepsilon$, and $k$, we generate $R=50$ independent data sets, denoted as $\mathbb{X}_{r}=\{\mathbf{x}_{1}^{(r)},\ldots,\mathbf{x}_{n}^{(r)}\}$, for $r=1,\ldots,R$.
6.3 Empirical bias
The effect of the distortion caused by having $P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ was measured through formulas (3.1) and (3.1). Following Maronna and Yohai (1995), we also use the condition number (CN) of an estimator $\hat{\Gamma}$ as an additional measure of bias, abbreviated as $B_{\text{CN}}$, and defined as the supremum of $\text{CN associated with }\hat{\Gamma}$ over the contamination neighbourhood.
To obtain empirical versions of $B$ and $B_{\text{CN}}$, we proceed as follows:
  • Step 1. For every $k=0,1,\ldots,25$, generate $R=50$ independent data sets, denoted by
    $\mathbb{X}_{k}^{(r)}=\{\mathbf{x}_{1}^{(r)},\ldots,\mathbf{x}_{n}^{(r)}\},\quad r=1,\ldots,R,$
    where each $\mathbb{X}_{k}^{(r)}$ is a sample from the distribution $P_{k}$. Let $\hat{\Gamma}_{k}^{(r)}$ denote the corresponding estimated scatter matrix. For every $\hat{\Gamma}_{k}^{(r)}$, compute $\hat{b}_{E,k}^{(r)}$ and $\hat{b}_{I,k}^{(r)}$, $r=1,\ldots,R$, by using the empirical version of $P_{k}$ corresponding to $\mathbb{X}_{k}^{(r)}$, denoted by $P_{n,k}^{(r)}$. Then, define the bias of the scatter matrix $\hat{\Gamma}_{k}^{(r)}$, based on the $r$-th sample from $P_{k}$, as
    $\hat{b}_{k}^{(r)}=\max\left\{\hat{b}_{I,k}^{(r)},\hat{b}_{E,k}^{(r)}\right\}=\max\left\{\lambda_{(1)}^{(r)},\lambda_{(p)}^{(r)-1}\right\},$
    where $\lambda_{(1)}^{(r)}$ and $\lambda_{(p)}^{(r)}$ denote the largest and the smallest eigenvalues of $\hat{\Gamma}_{k}^{(r)}$. For $1\leq r\leq R,$ let $_{k}^{(r)}=CN$$\left(\hat{\Gamma}_{k}^{(r)}\right)=\lambda_{(1)}^{(r)}/\lambda_{(p)}^{(r)}$ be the condition number of the estimated scatter matrix $\hat{\Gamma}_{k}^{(r)}.$ If $l$ is a location estimator, then $\hat{b}_{k}$ and CNk stand for
    $\hat{b}_{k}=l\left(\{\hat{b}_{k}^{(r)}\}\right),$ (21)
    $CN_{k}=l\{CN_{k}^{(r)}\}.$ (22)
    Step 2. The measures
    $\hat{B}=\max_{0\leq k\leq 25}\{\hat{b}_{i,k}\}\quad\text{and}\quad\hat{B}_{\text{CN}}=\max_{0\leq k\leq 25}\{CN_{k}\},$ (23)
    are empirical approximations to the maximum biases $B$ and $B_{\text{CN}}$, respectively.
In order to place the measures $\hat{b}_{I,k}^{(r)},\hat{b}_{E,k}^{(r)}$ and $CN_{k}^{(r)}$on a more comparable scale, we consider their logarithms. When sampling from a contaminated distribution $P_{k},$ the empirical distributions of $\log\left(\hat{b}_{I,k}^{(r)}\right),\log\left(\hat{b}_{E,k}^{(r)}\right)$ and $\log\left(CN_{k}^{(r)}\right)$ exhibit asymmetry and the presence of outliers, as shown by the boxplots in Figures 1 to 4. Therefore, it is more appropriate to use the median as the location measure $l$ in (21) and (22) rather than the mean. For the sake of simplicity, we refer to $\hat{B}$ and $\hat{B}_{\mathrm{CN}}$ as the “maximum medians of $B$ and CNCN”, respectively in the tables and figures. Further details are provided in the Section 6.5.
6.4 Efficiency
When the data come from the symmetric central model, the empirical distributions of $\log\left(CN_{k}^{(r)}\right)$ and $\log\left(b_{k}^{(r)}\right)$ tend to be more symmetric, and the mean is therefore more representative. Thus, given a scatter estimator, we can take $\hat{B}$ and $\hat{B}_{\mathrm{CN}}$ as defined in (23), but using the mean rather than the median in (21) and (22), which will be called the (empirical) mean absolute error, abbreviated as MAE. Furthermore, if $\hat{\Gamma}$ denotes a scatter estimator and $S_{n}$ denotes the sample covariance matrix, the ratio of the MAE for $S_{n}$ to the corresponding value for $\hat{\Gamma}$
[…]
may be considered a measure of efficiency at the normal central model $P_{0}$; see Maronna and Yohai (1995).
6.5 Results and conclusions
In the Tables 1 to 4 the efficiencies of scatter estimators are computed for different combinations of $n$ and $p$. For most estimators, efficiency remains stable or even slightly increases with growing $p$, suggesting favorable finite-sample behavior when the ratio $p/n$ is moderate. In particular, the MM estimator maintains high efficiency across all considered dimensions and sample sizes.
For the S-E, as can be seen in Tables 14, for small dimensions the efficiency is low, but it increases considerably as $p$ grows. This is a phenomenon described in the literature; see, for instance, Section 6.4.4. of Maronna et al. (2018). As discussed therein, when $p$ is large enough, almost all observations receive similar weights, except for observations far from the bulk of the data, yielding an estimator which closely approximates the sample covariance matrix. But increasing efficiency usually entails a loss of robustness when the dimension grows.
n MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.254 0.688 0.541 0.359 0.917 0.589 0.647
$p=5$ 0.339 0.736 0.826 0.565 0.941 0.601 0.798
$p=10$ 0.381 0.700 0.910 0.524 0.958 0.585 0.774
Table 1: Eff in log-scale based on CN (means) for $n=50$ over dimensions.
MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.446 0.690 0.580 0.433 0.806 0.617 0.282
$p=5$ 0.328 0.687 0.803 0.518 0.919 0.564 0.399
$p=10$ 0.355 0.678 0.876 0.444 0.952 0.582 0.471
Table 2: Eff in log-scale based on $B$ (means) for $n=50$ over dimensions.
n MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.224 0.630 0.515 0.378 0.944 0.567 0.630
$p=5$ 0.303 0.825 0.855 0.635 0.940 0.723 0.770
$p=10$ 0.385 0.856 0.951 0.706 0.951 0.786 0.846
$p=15$ 0.416 0.858 0.970 0.698 0.933 0.746 0.843
Table 3: Eff in log-scale based on CN (means) for $n=200$ over dimensions.
MVE MCD SE ROCKE $M$ SD MDEPTH
$p=2$ 0.368 0.667 0.583 0.452 0.875 0.636 0.151
$p=5$ 0.308 0.800 0.848 0.622 0.903 0.718 0.255
$p=10$ 0.339 0.811 0.935 0.652 0.935 0.754 0.331
$p=15$ 0.364 0.821 0.965 0.640 0.917 0.714 0.379
Table 4: Eff in log-scale based on $B$ (means) for $n=200$ over dimensions.
Figures 1 to 4 show boxplots of the biases $\{\hat{b}_{k}^{(r)}\}$, on a log scale, across the $R$ replicates for each scatter estimator, under selected scenarios (combinations of $p$, $n$, and $k$). The boxplots highlight the need for using medians rather than means as representative measures of the behavior of the bias measures. Note that most empirical distributions exhibit heavy tails, skewness, and outliers, especially when the contamination level is $\varepsilon=0.2$.
Figures 5 to 8 display the behavior of $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (see (21) and (22)), for each scatter estimator, as functions of $k$, for selected dimensions $p$ and sample sizes $n$, under both contamination levels. Tables 5 to 12 show the maximum medians of $CN$ and $B$ for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2,\ldots,15$. From the tables we can conclude that MM tends to give the smallest maximum medians of $B$ for small and moderate sample sizes and for lower dimensions, showing the best overall performance. ROCKE often outperforms MM for larger sample sizes and in higher dimensions ($p\geq 10$). The figures suggest that, in low-dimensional settings ($p=2$), MM and MCD perform best. For larger sample sizes, MM remains the best-performing estimator, followed by SE and ROCKE, while MCD also shows competitive performance. In moderate and higher dimensions ($p=5,10$), MM consistently exhibits strong performance across contamination levels. Generally speaking, the bias curves decrease as $k$ grows, reflecting the robustness of these estimators to extreme outliers.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
20 0.10 4.72 2.51 1.13 1.58 1.93 0.69 1.24 0.97
20 0.20 5.36 3.03 1.95 2.35 2.38 0.81 1.69 1.32
80 0.10 4.81 1.49 0.60 0.79 0.94 0.40 0.72 0.62
80 0.20 5.54 1.71 1.26 1.29 1.40 0.46 1.14 1.05
1000 0.10 4.86 0.98 0.63 0.61 0.80 0.24 0.52 0.48
1000 0.20 5.55 1.35 1.33 1.17 1.30 0.45 1.02 0.96
Table 5: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
20 0.10 4.77 1.29 1.37 1.48 1.65 0.78 1.30 1.65
20 0.20 5.35 1.51 1.67 1.90 2.20 0.82 1.60 1.82
80 0.10 4.74 0.75 0.48 0.51 0.65 0.34 0.48 1.18
80 0.20 5.32 0.86 1.14 0.99 1.14 0.52 0.93 1.62
1000 0.10 4.76 0.49 0.48 0.41 0.51 0.21 0.40 1.14
1000 0.20 5.31 0.68 1.18 0.80 0.90 0.52 0.77 1.52
Table 6: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator across sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=2$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
50 0.10 6.30 3.26 1.73 1.40 2.17 1.14 2.09 1.60
50 0.20 7.07 4.15 2.97 2.34 3.00 1.20 3.51 2.22
200 0.10 6.00 1.98 1.04 0.99 1.18 0.72 1.27 1.04
200 0.20 6.65 2.34 1.87 1.51 1.58 1.11 2.09 1.77
2500 0.10 5.81 1.57 0.66 0.76 0.75 0.52 0.82 0.92
2500 0.20 6.50 1.86 1.31 1.26 1.09 0.86 1.53 1.67
Table 7: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under contamination levels $\varepsilon$ and different sample sizes $n$ for dimension $p=5$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
50 0.10 5.75 1.97 1.24 0.89 1.55 0.69 1.34 1.73
50 0.20 6.24 2.41 2.35 1.42 2.26 0.78 2.09 2.32
200 0.10 5.69 1.08 0.67 0.55 0.66 0.41 0.70 1.37
200 0.20 6.22 1.27 1.34 0.87 0.96 0.63 1.23 1.82
2500 0.10 5.65 0.84 0.36 0.49 0.46 0.29 0.53 1.15
2500 0.20 6.22 1.02 0.80 0.83 0.60 0.47 1.04 1.54
Table 8: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under contamination levels $\varepsilon$ and different sample sizes $n$ for dimension $p=5$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
100 0.10 7.00 3.32 2.06 1.64 1.92 1.40 2.52 2.52
100 0.20 7.80 4.56 3.35 2.24 2.48 1.58 4.41 3.97
400 0.10 6.70 2.23 1.29 1.30 1.14 1.09 1.51 2.03
400 0.20 7.41 2.75 2.33 1.86 1.39 1.56 2.64 3.59
5000 0.10 6.51 2.11 0.94 1.06 0.75 0.85 1.11 1.79
5000 0.20 7.22 2.48 1.79 1.64 0.97 1.34 2.06 3.34
Table 9: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=10$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
100 0.10 6.34 1.87 1.40 0.90 1.33 0.77 1.35 2.04
100 0.20 6.88 2.41 2.49 1.31 1.79 0.81 2.67 2.39
400 0.10 6.30 1.26 0.69 0.86 0.61 0.66 0.97 1.53
400 0.20 6.89 1.48 1.50 1.31 0.77 1.01 1.85 2.32
5000 0.10 6.33 1.13 0.65 0.85 0.53 0.65 0.87 1.17
5000 0.20 6.91 1.40 1.01 1.34 0.62 1.04 1.69 2.00
Table 10: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=10$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
150 0.10 7.48 3.33 2.41 1.92 1.95 1.64 3.03 3.60
150 0.20 8.22 4.57 3.87 2.67 2.43 1.93 5.08 5.66
600 0.10 7.15 3.33 2.41 1.92 1.95 1.37 3.03 3.60
600 0.20 7.88 3.24 2.76 2.16 1.29 1.87 3.11 5.19
7500 0.10 6.92 2.46 1.17 1.30 0.69 1.10 1.06 2.25
7500 0.20 7.62 3.00 2.15 1.91 0.84 1.65 1.73 4.77
Table 11: Empirical maximum bias $\hat{\text{B}}_{\text{CN}}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension $p=15$.
$n$ $\varepsilon$ SCOV MVE MCD SE ROCKE $M$ SD MDEPTH
150 0.10 6.68 1.96 1.37 1.10 1.32 0.87 1.79 2.37
150 0.20 7.28 2.39 2.43 1.69 1.77 1.05 3.47 2.62
600 0.10 6.73 1.47 0.92 1.15 0.56 0.93 1.31 2.23
600 0.20 7.31 1.77 1.56 1.65 0.71 1.35 2.36 1.90
7500 0.10 6.73 1.33 0.88 1.12 0.50 0.92 0.89 1.23
7500 0.20 7.31 1.70 1.35 1.66 0.55 1.39 1.46 2.02
Table 12: Empirical maximum bias $\hat{B}$ (log scale) for each scatter estimator under different sample sizes $n$ and contamination levels $\varepsilon$ for dimension for dimension $p=15$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 1: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=2$, $n=20$ and $k=1$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 2: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=2$, $n=20$ and $k=20$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 3: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=15$, $n=150$ and $k=0$
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 4: Boxplots of the biases $\{\hat{b}_{k}^{(r)}\}_{1\leq r\leq R}$ (log-scale) through the $R=50$ replicates for each scatter estimator with $p=15$, $n=150$ and $k=25$
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 5: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=2$ and $n=20$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 6: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=2$ and $n=1000$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 7: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=10$ and $n=100$.
$\varepsilon=0.1$
$\varepsilon=0.2$
Figure 8: $\hat{b}_{k}=\text{median}_{1\leq r\leq R}\{\hat{b}_{k}^{(r)}\}$ (log-scale) versus $k$ for each scatter estimator, under contamination levels $\varepsilon$. Dimension $p=10$ and $n=5000$.
7 Appendix
Proof of Lemma 3.1. Chen, Gao and Ren (2018a), in the proof of Theorem 2.1, showed that, for $\mathbf{\theta}=\mathbf{0}$
[…]
$=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$
The upper bound becomes useless for $\varepsilon\geq 1/3$ since it is greater than 1. Therefore, we need to consider $\varepsilon<1/3-c.$ Put $b\left(p,n\right)=40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{p+1}{n}}+\frac{7}{2}\sqrt{\frac{\log\left(1/\delta\right)}{n}}$ and $a\left(\varepsilon\right)=\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}=\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}.$ Thus, (7) is equivalent to
$\left\|\hat{\theta}\right\| \displaystyle\leq$ $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)-\Phi^{-1}\left(\frac{1}{2}\right)+\Phi^{-1}\left(\frac{1}{2}\right)$
$=$ $\frac{1}{\phi\left(\Phi^{-1}\left(\gamma_{C}\right)\right)}\left(\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)\leq C\left(\sqrt{\frac{p}{n}}\vee\varepsilon+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),\text{ }$
$\text{with }\gamma_{C}$ $\in$ $\left(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right).$
Therefore, consider $\Phi^{-1}\left(a\left(\varepsilon\right)\right)$ rather than $\Phi^{-1}\left(\frac{1}{2}\right)$ to center $\Phi^{-1}\left(\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}+b\left(p,n\right)\right)$ and we get
$\left\|\hat{\theta}\right\| \displaystyle\leq$ $\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)-\Phi^{-1}\left(a\left(\varepsilon\right)\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
$=$ $\frac{1}{\phi\left(\Phi^{-1}\left(\eta_{B}\right)\right)}b\left(p,n\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right),\text{ \ }\eta_{B}\in\left(a\left(\varepsilon\right),a\left(\varepsilon\right)+b\left(p,n\right)\right)$
$\leq$ $\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+b\left(p,n\right)\right)\right)}b\left(p,n\right)+\Phi^{-1}\left(a\left(\varepsilon\right)\right)$
If $\varepsilon<1/3-c$, where $c$ is a positive constant, $a\left(\varepsilon\right)$ is increasing on [0,1/3c)[0,1/3-c), let $a(m_{c})$ be the maximum value. Take $d$ such that $1-a\left(m_{c}\right)>d>0$ and $\left(p,n\right)$ $\in$ $A_{c,d}$ == $\left\{\left(p,n\right):b\left(p,n\right)<1-a\left(m_{c}\right)-d\right\}.$ Consequently, there exists a constant
$C_{c,d}=\sup_{\left(p,n\right)\in A_{c,d}}\left[\phi\left(\Phi^{-1}\left(a\left(m_{c}\right)+b\left(p,n\right)\right)\right)\right]^{-1}$
such that
$\left\|\hat{\theta}\right\|\leq\tilde{C}_{c,d}\left(\sqrt{\frac{p}{n}}\vee B_{L}\left(\mathbf{\hat{\theta}}_{T},\varepsilon,\Phi\right)+\sqrt{\frac{\log\left(1/\delta\right)}{n}}\right),$
[…]
Proof of Lemma 3.2. In the process of obtaining (11), Chen, Gao and Ren (2018a) (p. 1955) showed that, with probability greater than $1-2\delta,$
$\sup_{\mathbf{u\in}\emph{S}^{p-1}}\left|\Phi\left(\sqrt{\beta}\right)-\Phi\left(\frac{\mathbf{u}^{t}\hat{\Gamma}\mathbf{u}}{\mathbf{u}^{t}\Sigma\mathbf{u}}\right)\right|\leq\frac{\varepsilon}{2(1-\varepsilon)}+\sqrt{\frac{2400e\pi}{1-e^{-1}}}\sqrt{\frac{3+2p}{n}}+\frac{7}{4}\sqrt{\frac{\log\left(1/\delta\right)}{n}}.$ (26)
holds. Put $a_{n,\delta}=40\sqrt{\frac{6e\pi}{1-e^{-1}}}\sqrt{\frac{3+2p}{n}}+\frac{7}{2}\sqrt{\frac{\log(1/\delta)}{n}}$ and $\sqrt{\beta}\sqrt{\frac{\mathbf{u}^{T}\left(\hat{\Gamma}/\beta\right)u}{\mathbf{u}^{T}\Sigma\mathbf{u}}}=\sqrt{\beta}x_{\mathbf{u}}$, with $a_{n,\delta}\geq 0$. Then,
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|$ (27)
$=$ $\max\left(\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}),\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)\right)$ (28)
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$, then
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right).$
If $1\leq\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}},$ then
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}x_{\mathbf{u}}\right)\right|=\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta}).$
If $\inf_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\leq 1\leq\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}$ we have to analyze both cases. If the maximum on the right-hand side of (27) occurs in $\Phi\left(\sqrt{\beta}\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)-\Phi(\sqrt{\beta})$ then $\sup_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\geq 1,$ and we have,
$\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right)\leq\Phi(\sqrt{\beta})+\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}.$
By calling $a\left(\varepsilon\right)=\frac{3}{4}+\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}$ an increasing function in $\varepsilon$, we have
$\sup_{u\in S^{p-1}}x_{u}$ $\leq$ $\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)+1$
$+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]$
Then there exists $\eta\in\left[a\left(\varepsilon\right),a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right]$ such that
$\sup_{u\in S^{p-1}}x_{u}-1\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(\eta\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right],\text{ }$
$0$ $\leq$ $\sup_{u\in S^{p-1}}x_{u}-1$
$\leq$ $\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(a\left(\varepsilon\right)+\frac{1}{2}a_{n,\delta}\right)\right)}a_{n,\delta}+\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right]=g_{S}\left(\varepsilon,n,\delta,p\right).$
Then,
$B_{E}\left(\varepsilon\right)=\left[\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(a\left(\varepsilon\right)\right)-1\right].\text{ }$
If the maximum in (27) occurs at $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\right),$ then $\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq 1$ and we call $b\left(\varepsilon\right)=\frac{3}{4}-\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)},$ which is decreasing in $\varepsilon.$ Thus, we get that
$0$ $\leq$ $\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)\leq\frac{\varepsilon}{2\left(1-\varepsilon\right)}+\frac{1}{2}a_{n,\delta}$
$\Phi\left(\sqrt{\beta}\inf_{u\in S^{p-1}}x_{u}\right)$ $\geq$ $\Phi(\sqrt{\beta})-\frac{\varepsilon}{2\left(1-\varepsilon\right)}-\frac{1}{2}a_{n,\delta}$
$\inf_{u\in S^{p-1}}x_{u}$ $\geq$ $\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1+1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)$
$+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)-\frac{1}{2}a_{n,\delta}\right)$
Then, there exists $\eta\in\left(b\left(\varepsilon\right)-\frac{1}{2}a_{n,\delta},b\left(\varepsilon\right)\right)$ such that
$\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}-1$ $\geq$ $-\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(\eta\right)\right)}a_{n,\delta}+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1,\text{ }$
$\geq$ $-\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(b\left(\varepsilon\right)\right)\right)}a_{n,\delta}+\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)-1.$
Then,
$0\leq 1-\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}}\leq\frac{1}{2\sqrt{\beta}}\frac{1}{\phi\left(\Phi^{-1}\left(b\left(\varepsilon\right)\right)\right)}a_{n,\delta}+1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(b\left(\varepsilon\right)\right)=g_{I}\left(\varepsilon,n,\delta,p\right).$
Call
$B_{I}\left(\varepsilon\right)=1-\frac{1}{\sqrt{\beta}}\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right).$
It is easy to show that $B_{I}\left(\varepsilon\right)\leq B_{E}\left(\varepsilon\right).$ Observe that there is an $\eta\in\left[\beta\inf_{\mathbf{u}\in S^{p-1}}x_{\mathbf{u}},\beta\right]\cup\left[\beta,\beta\sup_{\mathbf{u\in}S^{p-1}}x_{\mathbf{u}}\right]$ such that
$\sup_{\mathbf{u}\in S^{p-1}}\left|\Phi(\sqrt{\beta})-\Phi\left(\sqrt{\beta x_{\mathbf{u}}}\right)\right|$ $=$ $\sup_{\mathbf{u}\in S^{p-1}}\left|\left(\beta-\beta x_{\mathbf{u}}\right)\frac{\varphi\left(\eta\right)}{\sqrt{\eta}}\right|$
$\geq$ $\frac{\varphi\left(1+g_{S}\left(\varepsilon,n,p,\delta\right)\right)}{\sqrt{1+g_{S}\left(\varepsilon,n,p,\delta\right)}}\sup_{\mathbf{u}\in S^{p-1}}\left|\left(\beta-\beta x_{\mathbf{u}}\right)\right|.$
Then, we can conclude that
$\inf_{\Sigma\in\mathcal{F}\left(M\right),P\in\mathcal{P}_{\varepsilon}\left(P_{0}\right)}P\left(\left\|\hat{\Sigma}-\Sigma\right\|_{op}^{2}\leq C^{\ast}\left(\max\left\{\frac{p}{n},B_{E}^{2}(\varepsilon)\right\}+\frac{\log\left(1/\delta\right)}{n}\right)\right)\geq\alpha.$
[…]
Proof of Lemma 3.4. Since we have regression through the origin model (a zero intercept model) and the covariables have an ellipsoidal distribution and we are dealing with affine equivariant estimators we can assume that $B=0,$ $\Sigma=I$ and $\sigma=1$, Then, $P_{B},$ the joint distribution of $\left(\mathbf{X},\mathbf{Y}\right)\in\mathbb{R}^{p}\times\mathbb{R}^{m},$ is given by
$\left(\begin{array}[]{c}\mathbf{Y}\\ \mathbf{X}\end{array}\right)\sim N_{p+m}\left(\left(\begin{array}[]{c}0_{m\times 1}\\ 0_{p\times 1}\end{array}\right),\left(\begin{array}[]{cc}I_{m\times m}&0_{m\times p}\\ 0_{p\times m}&I_{p\times p}\end{array}\right)\right)$
Then, if $A,U\in\mathbb{R}^{p\times m},$ it holds that
$\left(\begin{tabular}[]{l}$\mathbf{Y}-A^{t}\mathbf{X}$\\ $U^{t}\mathbf{X}$\end{tabular}\right)\sim N_{2m}\left(\left(\begin{array}[]{c}0_{m\times 1}\\ 0_{m\times 1}\end{array}\right),\left(\begin{array}[]{cc}I_{m\times m}+B^{t}B&-B^{t}U\\ -U^{t}B&U^{t}U\end{array}\right)\right).$
Given $D\left(A,P_{B}\right)=\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0),$ from the proof of Theorem 4.1, p.1167 of Gao (2020), we have the following inequality
$D\left(\hat{A},P_{B}\right)\geq\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}-C_{1}\left(\sqrt{\frac{mp+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right),$ (29)
with $C_{1}$ an absolute constant (independent of $n,p,m,\varepsilon)$). Then,
$P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0)$ $=$
$=$ $E_{X}\left[P\left(\left.X^{t}UA^{t}X\leq X^{t}UY\right|X\right)\right]\text{ \ }(\left.X^{t}UY\right|X\sim N\left(0,\left\|X^{t}U\right\|^{2}\right)$
$=$ $E\left[1-\Phi\left(\frac{X^{t}UA^{t}X}{\left\|X^{t}U\right\|}\right)\right]=1-E\Phi\left(\left(\frac{U^{t}X}{\left\|U^{t}X\right\|},A^{t}X\right)\right).$
Therefore,
$\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)$ $\geq$ $0)=1-\sup_{U\in\mathbb{R}^{p\times m}}E\Phi\left(\left(\frac{U^{t}X}{\left\|U^{t}X\right\|},A^{t}X\right)\right)$
$=$ $1-E\Phi\left(\left\|A^{t}X\right\|\right).$
Take the spectral decomposition of the non-negative matrix $A^{t}\in\mathbb{R}^{p\times p}, AA$$A^{t}=PDP^{t}$ $P\in\mathbb{R}^{p\times p}$ orthonormal and $D=diag\left(\lambda_{1},...,\lambda_{p}\right)$ with $0\leq\lambda_{0}\leq\lambda_{1}\leq\dots\leq\lambda_{p}$, then $Z=PX\sim N_{p}\left(0,I\right).$ Therefore,
$E\Phi\left(\sqrt{X^{t}AA^{t}X}\right)=E\Phi\left(\sqrt{Z^{t}DZ}\right)=E\Phi\left(t\sqrt{Z^{t}\frac{D}{tr\left(D\right)}Z}\right)=E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right),$
with $t=\sqrt{tr\left(D\right)},$ $\left\{\theta_{i}\right\}_{i=1}^{p}\subset\left[0,1\right]$ and $\sum_{i=1}^{p}\theta_{i}=1$. The square root is a concave function, and we can obtain that
$\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\geq\sum_{i=1}^{p}\theta_{i}\left|Z_{i}\right|.$
$\Phi$ is also a concave function on $0,\infty)$, then take the functions $h:[0,\infty)\times\left[0,1\right]^{p}\rightarrow\left[0.5,1\right]$ and $g:[0,\infty)\rightarrow\left[0.5,1\right]$ as $h\left(t,\theta_{1},\dots,\theta_{p}\right)=\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ and $g\left(t\right)=h\left(t,1,0,...,0\right),$
$\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ $\geq$ $\sum_{i=1}^{p}\theta_{i}\Phi\left(t\left|Z_{i}\right|\right)$
$E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)$ $\geq$ $\sum_{i=1}^{p}\theta_{i}E\Phi\left(t\left|Z_{i}\right|\right)=E\Phi\left(t\left|Z_{1}\right|\right)=g\left(t\right),$
since $E\Phi\left(t\left|Z_{1}\right|\right)=\dots=E\Phi\left(t\left|Z_{p}\right|\right).$ Then
$\inf_{U\in\mathbb{R}^{p\times m}}P_{B}\left((U^{t}X,Y-A^{t}X\right)\geq 0)=1-E\Phi\left(t\sqrt{\sum_{i=1}^{p}\theta_{i}Z_{i}^{2}}\right)\leq 1-E\Phi\left(t\sqrt{Z_{1}^{2}}\right)=1-g\left(t\right).$
Then, from (29) we can say,
$D\left(\hat{A},P_{B}\right)$ $=$ $1-g\left(\hat{t}\right)\geq\frac{1-3\varepsilon}{2\left(1-\varepsilon\right)}-C_{1}\left(\sqrt{\frac{mp+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)$
[…]
Since $g$ is increasing, we get that
$\hat{t} \displaystyle\leq$ $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right)$
$=$ $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right)-g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)+g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$
$=$ $C_{1}\left(g^{-1}\right)^{\prime}\left(\xi\right)\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)+g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right),\text{ }$
$\xi$ $\in$ $\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)},\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}+C_{1}\left(\sqrt{\frac{pm+p}{n}}+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right)\right).$
Thus,
$\hat{t}\leq C\left[\left(\frac{pm}{n}\vee g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)\right)+\sqrt{\frac{\log\left(1/\delta\right)}{2n}}\right]$
[…]
Remark 7.1
Observe that $g^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$ does not depend either on the number of dependent variables $m$ or the number of independent variables $p.$
Proof of Lemma 3.5. Since we are dealing with affine equivariant estimators we can assume that $\boldsymbol{\beta}=0,$ $\Sigma=I$ and $\sigma=1$. With this background we get that
$\left(\begin{array}[]{c}y\\ \mathbf{X}\end{array}\right)\sim N_{p+1}\left(\left(\begin{array}[]{c}0\\ \mathbf{0}_{p\times 1}\end{array}\right),\left(\begin{array}[]{cc}1&\mathbf{0}_{p\times 1}^{t}\\ \mathbf{0}_{p\times 1}&I_{p\times p}\end{array}\right)\right)$
Then, it holds that
$\left(\mathbf{\lambda}^{t}\mathbf{X},y-\mathbf{\alpha}^{t}\mathbf{X}\right)^{t}\sim N_{2}\left(\left(\begin{array}[]{c}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}1&-\mathbf{\alpha}^{t}\mathbf{\lambda}\\ -\mathbf{\alpha}^{t}\mathbf{\lambda}&1+\left\|\mathbf{\alpha}\right\|^{2}\end{array}\right)\right)$
for a unitary vector $\mathbf{\lambda}.$ Since
$P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)\geq 0)=1-E\Phi\left(\frac{\mathbf{\lambda}^{t}\mathbf{XX}^{t}\mathbf{\alpha}}{\sigma\left|\mathbf{\lambda}^{t}\mathbf{X}\right|}\right).$
Since $h\left(-\rho\right)=1-h\left(\rho\right)$ and $\dot{h}:\left[-1,1\right]\rightarrow\left[0,1\right]$ is increasing, we have that, for $b=\left\|\alpha\right\|$
$P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)$ $\geq$ $0)=1-P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)<0)$
$=$ $1-\left[1-h\left(-\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)\right]$
$=$ $h\left(-\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)=1-h\left(\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)$
This entails that,
$\inf_{\left\|\mathbf{\lambda}\right\|=1}P_{B}\left(\left(\mathbf{\lambda}^{t}\mathbf{X}\right)(y-\mathbf{X}^{t}\mathbf{\alpha})\right)$ $\geq$ $0)=1-\sup_{\left\|\lambda\right\|=1}E\Phi\left(\frac{\mathbf{\lambda}^{t}\mathbf{XX}^{t}\mathbf{\alpha}}{\left|\mathbf{\lambda}^{t}\mathbf{X}\right|}\right)$
$=$ $1-E\Phi\left(\left\|\mathbf{\alpha}\right\|\frac{\left|\mathbf{\alpha}^{t}\mathbf{X}\right|}{\left\|\mathbf{\alpha}\right\|}\right)=1-E\Phi\left(\left\|\mathbf{\alpha}\right\|\left|Z\right|\right)=1-g\left(\left\|\mathbf{\alpha}\right\|\right)$
$=$ $1-\sup_{\left\|\lambda\right\|=1}h\left(\frac{\mathbf{\alpha}^{t}\mathbf{\lambda}}{\sqrt{1+b^{2}}}\right)=1-h\left(\frac{b}{\sqrt{1+b^{2}}}\right)$
Then,
$h\left(\frac{b}{\sqrt{1+b^{2}}}\right)=g\left(b\right)=\frac{1+\varepsilon}{2\left(1-\varepsilon\right)},$
[…]
7.1 Maximum bias of the deepest scatter estimator
Proof of Lemma 4.1. Given $g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right),$ the Lagrangian is given by
$h\left(\mathbf{u,}\lambda\right)=g\left(\mathbf{u}\right)+\lambda\left(\mathbf{u}^{t}\mathbf{u-}1\right).$
By differentiating the function $h$ we get
$\mathbf{0}\mathbf{=}\frac{\partial h\left(\mathbf{u,}\lambda\right)}{\partial\mathbf{u}}=\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\left[\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}\right]$ $-\varphi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\left[-\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}\right]+2\lambda\mathbf{u}$
$0$ $\mathbf{=}$ $\frac{\partial h\left(\mathbf{u,}\lambda\right)}{\partial\lambda}=\mathbf{u}^{t}\mathbf{u-}1.$
Therefore, we obtain that
$\mathbf{0}\mathbf{=}\frac{\partial h\left(\mathbf{u,\lambda}\right)}{\partial\mathbf{u}}=\left[2\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right]\frac{\Gamma\mathbf{u}}{\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}}+2\lambda\mathbf{u}$
Call $b\left(\mathbf{u}\right)=-\left[2\varphi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right]/\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}<0,$ which yields $b\left(\mathbf{u}\right)\Gamma\mathbf{u}\mathbf{=}2\lambda\mathbf{u}$ with $\lambda=\frac{1}{2}b\left(\mathbf{u}\right)\mathbf{u}^{t}\Gamma\mathbf{u}$ and consequently, $\left(\Gamma\mathbf{-}\left[\mathbf{u}^{t}\Gamma\mathbf{u}\right]I\right)\mathbf{u}\mathbf{=}\mathbf{0.}$ Thus, this entails that the critical points of the Lagrangian are the eigenvectors of $\Gamma.$ Therefore, the function,
$g\left(\mathbf{u}\right)=\Phi\left(\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)-\Phi\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)$
has the following Lagrangian critical points we have
$g\left(\mathbf{v}_{j}\right)$ $=$ $\Phi\left(l_{j}^{1/2}\right)-\Phi\left(-l_{j}^{1/2}\right)=2\Phi\left(l_{j}^{1/2}\right)-1,\text{ }j=1,...,p$
$g\left(\mathbf{v}_{p}\right)$ $\leq$ $g\left(\mathbf{v}\right)\leq g\left(\mathbf{v}_{1}\right)\text{ for all }\mathbf{v\in}\emph{S}^{p-1}$
$1-g\left(\mathbf{v}_{j}\right)$ $=$ $1-\left[\Phi\left(l_{j}^{1/2}\right)-\Phi\left(-l_{j}^{1/2}\right)\right]=2\left[1-\Phi\left(l_{j}^{1/2}\right)\right],\text{ }j=1,...,p$
$1-g\left(\mathbf{v}_{p}\right)$ $\geq$ $g\left(\mathbf{v}\right)\geq 1-g\left(\mathbf{v}_{1}\right)\text{ for all }\mathbf{v\in}\emph{S}^{p-1}$
[…]
Proof of Corollary 4.2 Since $g(\mathbf{v}_{p})\leq g(\mathbf{v}_{1})$ and $g$ is an increasing function of $\sqrt{\mathbf{v}^{\top}\Gamma\mathbf{v}}$, $\min\bigl(g(\mathbf{v}_{p}),\,1-g(\mathbf{v}_{1})\bigr)$ is maximized when $g(\mathbf{v}_{p})=1-g(\mathbf{v}_{1})=0.5$, which implies $\ell_{p}^{1/2}=\ell_{1}^{1/2}=\Phi^{-1}(3/4)$. Hence, $\hat{\Gamma}=\left[\Phi^{-1}(3/4)\right]^{2}I_{p},$ and the corresponding maximum depth equals to $1/2$.
Proof of Corollary 4.3 For every vector $\mathbf{w}\in\mathbb{R}^{p}$, define
$Z_{n}(\mathbf{w})=\frac{\mathbf{w}^{t}\Gamma_{n}^{-1/2}\mathbf{X}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\in\mathbb{R}.$
Let
$c_{n}(\mathbf{w})=\frac{1}{\displaystyle\sum_{j=1}^{p}l_{j}^{(n)-1}\left(\mathbf{w}^{t}\mathbf{e}_{j}\right)^{2}}.$
and consider the probabilities
$a_{\varepsilon,n}(\mathbf{w})$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}(\mathbf{w})^{2}\leq c_{n}(\mathbf{w})\right)+\varepsilon\,P_{n}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\leq c_{n}(\mathbf{w})\right),$
$b_{\varepsilon,n}(\mathbf{w})$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\geq c_{n}(\mathbf{w})\right)+\varepsilon\,P_{n}\!\left(Z_{n}(\mathbf{w}_{r})^{2}\geq c_{n}(\mathbf{w})\right).$
[…]
Case i) $l_{1}^{(n)}\rightarrow\infty$.
By definition of the depth, $D(\Gamma_{n},P_{n})=\min\{a_{\varepsilon,n}(\mathbf{w}_{n}),b_{\varepsilon,n}(\mathbf{w}_{n})\}\leq\min\{a_{\varepsilon,n}(\mathbf{w}),b_{\varepsilon,n}(\mathbf{w})\}$ for every vector $\mathbf{w}$. Since
$D(\Gamma_{n},P_{n})=\min\{a_{\varepsilon,n}(\mathbf{w}_{n}),b_{\varepsilon,n}(\mathbf{w}_{n})\}\leq\min\{a_{\varepsilon,n}(\mathbf{v}_{1}^{(n)}),b_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\},$
it holds that $a_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\geq D(\Gamma_{n},P_{n})\quad b_{\varepsilon,n}(\mathbf{v}_{1}^{(n)})\geq D(\Gamma_{n},P_{n})$. By considering that $\{\mathbf{v}_{j}^{(n)}\}_{j=1}^{p}$ is an orthonormal basis, we have that
$\frac{1}{\sum_{j=1}^{p}l_{j}^{(n)-1}\bigl(\mathbf{v}_{1}^{(n)T}\mathbf{e}_{j}\bigr)^{2}}=l_{1}^{(n)}.$
and
$b_{\varepsilon,n}\!\left(\mathbf{v}_{1}^{(n)}\right)$ $=$ $(1-\varepsilon)\,P_{0}\!\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)+\varepsilon\,P_{n}\!\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)$
$\geq$ $D(\Gamma_{n},P_{n}).$
Since $P_{0}\left(Z_{n}\!\left(\mathbf{v}_{1}^{(n)}\right)^{2}\geq l_{1}^{(n)}\right)\rightarrow 0$ when $l_{1}^{(n)}\rightarrow\infty$, we may conclude that $\varepsilon\geq$ $\lim_{n\rightarrow\infty}D\left(\Gamma_{n},P_{n}\right)\geq$ $\lim_{n\rightarrow\infty}D\left(I,P_{n}\right)\geq$ $\left(1-\varepsilon\right)D\left(I,P_{0}\right)=\left(1-\varepsilon\right)\frac{1}{2}$which says that $\varepsilon\geq 1/3.$
Case ii) $l_{p}^{(n)}\rightarrow 0$. With a similar analysis to the case i) we can conclude that
$b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)\text{ and }a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right).$
Since
$a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)=(1-\varepsilon)\,P_{0}\!\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)+\varepsilon\,P_{n}\!\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)\geq D\!\left(\Gamma_{n},P_{n}\right).$
and considering that $P_{0}\left(Z_{n}\!\left(\mathbf{v}_{p}^{(n)}\right)^{2}\leq l_{p}^{(n)}\right)\rightarrow 0$ when $l_{p}^{(n)}\rightarrow 0$, we get
$\varepsilon\geq\lim_{n\rightarrow\infty}D\left(\Gamma_{n},P_{n}\right)\geq\lim_{n\rightarrow\infty}D\left(I,P_{n}\right)\geq\left(1-\varepsilon\right)D\left(I,P_{0}\right)=\left(1-\varepsilon\right)\frac{1}{2}$
which says that $\varepsilon\geq 1/3.\quad\square$
[…]
Lemma 7.2
The critical points of $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)$ (respectively $1-\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right))$ subject to $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\geq 0$ (idem $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\leq 0)$ are either $\mathbf{v}_{p},$ $\mathbf{v}_{1}$ or occurs at $F_{r}^{\mathbf{e}}.$
Proof. Take the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\lambda\left(\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right),$ Karush, Kuhn, Tucker condition says that if the minimum local occurs in $\mathbf{\tilde{v},}$ then, either $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}$ == 0 or we have to consider the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right).$ In the last case, it was shown that the critical points are the eigenvectors of $\Gamma.$ Suppose that $\mathbf{v}_{j_{0}\text{,}}$ $j_{0}<p,$is the eigenvector corresponding to the minimum $l_{j_{0}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$, by the continuity of the functions we could find another $\mathbf{\bar{v}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>\mathbf{\bar{v}}_{\text{ }}^{t}\Gamma\mathbf{\bar{v}}>r^{2}\left(\mathbf{\bar{v}}^{t}\mathbf{e}\right)^{2}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$ which is a contradiction. Therefore the critical points occur at $\mathbf{\tilde{v}=v}_{p}$ $\left(\mathbf{v}_{1}\right)$ or $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0.$
Proof of Corollary 4.6 Set $q_{j}=\frac{l_{j}^{1/2}}{\mathbf{v}_{j}^{t}\mathbf{e}},$ then $q_{1}=l_{1}^{1/2}$ and $q_{j}=\infty,$ j=2,p.j=2...,p. If $r<l_{1}^{1/2}$ then
$B_{r}^{\mathbf{e}}=\emptyset,Gb_{r}^{\mathbf{e}}=-\infty\text{ and }gb_{r}^{\mathbf{e}}=\infty.$
$\left\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{p}\right\}\subseteq A_{r}^{\mathbf{e}}\text{ then }Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right)\text{ and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
Thus,
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\right\}.$
If $r>l_{1}^{1/2}$ then
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}$ $\subseteq$ $\left(B_{r}^{\mathbf{e}}\right)^{c},\mathbf{v}_{1}\in B_{r}^{\mathbf{e}},\text{ }Gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right),$
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}$ $\subseteq$ $A_{r}^{\mathbf{e}},\text{ }\mathbf{v}_{1}\notin A_{r}^{\mathbf{e}},ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
We need to determine $gb_{r}^{\mathbf{e}}$ and $Ga_{r}^{\mathbf{e}}.$ Then the function $g\left(\mathbf{v}\right)$ has a minimum in $B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}$ (a compact set) and the minimum has to occur in $F_{r}^{\mathbf{e}}.$ Likewise, the function $g\left(\mathbf{v}\right)$ has a maximum in $A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}$ which must occur at $F_{r}^{\mathbf{e}}.$ In both cases, Lemma 7.2 supports that statement. We use spherical coordinates,
$\mathbf{v}^{t}\mathbf{v}_{1}$ $=$ $\cos\eta_{0},\text{ }$
$\mathbf{v}^{t}\mathbf{v}_{2}$ $=$ $\sin\eta_{0}\cos\eta_{1},$
$\mathbf{v}^{t}\mathbf{v}_{3}$ $=$ $\sin\eta_{0}\sin\eta_{1}\cos\eta_{2},\text{ }0\leq\eta_{0},..\eta_{p-3}\leq\pi$
$\vdots$
$\mathbf{v}^{t}\mathbf{v}_{p-1}$ $=$ $\sin\eta_{0}\sin\eta_{1}\sin\eta_{2}\dots\cos\eta_{p-2},0\leq\eta_{p-2}\leq 2\pi$
$\mathbf{v}^{t}\mathbf{v}_{p}$ $=$ $\sin\eta_{0}\sin\eta_{1}\sin\eta_{2}\dots\sin\eta_{p-2},$
($\eta_{j}$ is the angle between $\mathbf{v}_{j+1}$ and $\mathbf{v,}$ $j=0,...,p-1\mathbf{).}$ Then
$\mathbf{v}^{t}\Gamma\mathbf{v=}r^{2}\left(\mathbf{v}^{t}\mathbf{v}_{1}\right)^{2}$
is equivalent to
$l_{1}\cos^{2}\eta_{0}+\sin^{2}\eta_{0}\left[l_{2}\cos^{2}\eta_{1}+l_{3}\sin^{2}\eta_{1}\cos^{2}\eta_{2}+\dots+l_{p}\sin^{2}\eta_{1}\sin^{2}\eta_{2}\dots\sin^{2}\eta_{p-2}\right]=r^{2}\cos^{2}\eta_{0}$
$tg^{2}\eta_{0}=\frac{r^{2}-l_{1}}{(l_{2}\cos^{2}\eta_{1}+\sin^{2}\eta_{1}(\dots(l_{p-2}\cos^{2}\eta_{p-3}+\sin^{2}\eta_{p-3}\left(l_{p-1}\cos^{2}\eta_{p-2}+l_{p}\sin^{2}\eta_{p-2}\right))...)}$
Observe that the denominator are nested convex combinations. Thus, we can ensure that
$\frac{r^{2}-l_{1}}{l_{2}}$ $\leq$ $tg^{2}\eta_{0}\leq\frac{r^{2}-l_{1}}{l_{p}}$
$\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$ $\leq$ $\left|tg\left(\eta_{0}\right)\right|\leq\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}$
$\text{either }\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$ $\leq$ $tg\left(\eta_{0}\right)\leq\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\text{ or }-\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\leq tg\left(\eta_{0}\right)\leq-\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$
$\text{either }0$ $\leq$ $\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\leq\eta_{0}\leq\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\leq\pi/2\text{ }$
$\text{or }\pi/2$ $\leq$ $\arctan\left(-\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\leq\eta_{0}\leq\arctan\left(-\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\leq\pi$
Then, by similarity of triangles we can say that, for either $l=l_{2}$ or $l_{p},$
$\frac{\sqrt{1+\frac{r^{2}-l_{1}}{l}}}{1}=\frac{1}{\cos\left(\theta\right)},\text{ \ \ \ }\frac{\sqrt{1+\frac{r^{2}-l_{1}}{l}}}{1}=\frac{\sqrt{\frac{r^{2}-l_{1}}{l}}}{\sin\left(\theta\right)}$
Therefore, we obtain that
$\max_{A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$ $\mathbf{=}$ $\max_{F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v=}r^{2}\cos^{2}\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\right]=r^{2}\frac{l_{2}}{r^{2}+l_{2}-l_{1}}\text{ }$
$\mathbf{v}_{M,r}$ $=$ $\arg\max_{F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$
$\mathbf{=}$ $\left(\cos\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right),\sin\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right),0,\dots,0)\right)$
$=$ $\left(\sqrt{\frac{l_{2}}{r^{2}+l_{2}-l_{1}}},\sqrt{\frac{r^{2}-l_{1}}{r^{2}+l_{2}-l_{1}}},0,\dots,0\right)$
$\min_{B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$ $\mathbf{=}$ $r^{2}\cos^{2}\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right]=r^{2}\frac{l_{p}}{r^{2}+l_{p}-l_{1}}$
$\mathbf{v}_{m,r}$ $=$ $\arg\min_{B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$
$\mathbf{=}$ $\left(\cos\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right),0,\dots,0,\sin\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right)$
$=$ $\left(\sqrt{\frac{l_{p}}{r^{2}+l_{p}-l_{1}}},0,\dots,0,\sqrt{\frac{r^{2}-l_{1}}{r^{2}+l_{p}-l_{1}}}\right)$
which means that
$gb_{r}^{\mathbf{e}}$ $=$ $g\left(\mathbf{v}_{m,r}\right)=2\Phi\left(r\cos\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right]\right)-1,$
$Ga_{r}^{\mathbf{e}}$ $=$ $g\left(\mathbf{v}_{M,r}\right)=2\Phi\left(r\cos\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\right]\right)-1.$
Since $\mathbf{v}_{M,r}$ is a maximum over $A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}},$ $g\left(\mathbf{v}_{2}\right)\leq g\left(\mathbf{v}_{M,r}\right)\leq g\left(\mathbf{v}_{1}\right).$ Then, we have that
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\\ \left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\end{array}\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{m,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{M,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right.$. (30)
[…]
Proof of Lemma 4.7 If $r<l_{1}^{1/2}$ then the deepest matrix should verify
$\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon$ $=$ $\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)$
$g\left(\mathbf{v}_{p}\right)+g\left(\mathbf{v}_{1}\right)$ $=$ $\frac{1-2\varepsilon}{1-\varepsilon}.$
To increase the depth, $g\left(\mathbf{v}_{p}\right)$ should increase and $g\left(\mathbf{v}_{1}\right)$ should decrease, and then they would reach
$2g\left(\mathbf{v}_{p}\right)=2g\left(\mathbf{v}_{1}\right)=\frac{1-2\varepsilon}{1-\varepsilon}.$
This implies that
$g\left(\mathbf{v}_{p}\right)$ $=$ $g\left(\mathbf{v}_{1}\right)=\frac{1-2\varepsilon}{2(1-\varepsilon)}$
$l_{p}^{1/2}$ $=$ $l_{1}^{1/2}=\Phi^{-1}\left(\frac{3-4\varepsilon}{4(1-\varepsilon)}\right)$
and the deepest matrix is a multiple of the identity matrix.
Let us take $l_{1}^{1/2}<r.$ Thus we get
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{m,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{M,r}\right)\right)\right)\end{array}\right\}.$
with $g(\mathbf{v}_{p})\leq...\leq g(\mathbf{v}_{2})\leq g\left(\mathbf{v}_{M,r}\right)\leq g\left(\mathbf{v}_{1}\right).$ The best estimator should be selected so that $g\left(\mathbf{v}_{m,r}\right)=g\left(\mathbf{v}_{M,r}\right)$ (by making $l_{2}=\dots=l_{p},$ $g\left(\mathbf{v}_{m,r}\right)$ increases) and $\mathbf{v}_{m,r}=\mathbf{v}_{M,r}.$ Call $\mathbf{v}_{0,r}=\mathbf{v}_{m,r}$ and $\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right)=(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right),$ which entails that $g\left(\mathbf{v}_{0,r}\right)=1/2$ or $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=\mathbf{\Phi}^{-1}\left(3/4\right).$  Then we take matrices $\Gamma$ such that $\Phi^{-1}\left(3/4\right)=0.6745\leq l_{1}^{1/2}$ . All the four quantities involved in $D\left(\Gamma,P_{\varepsilon,r}\right)$ must be equal in order to get the best estimator In this case we get that
$\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon$ $=$ $\frac{\left(1-\varepsilon\right)}{2}=\left(1-\varepsilon\right)g_{p}+\varepsilon$
$\left(1-g_{1}\right)+\frac{\varepsilon}{1-\varepsilon}$ $=$ 12 ; $\frac{1}{2}\text{ \ };\text{ \ }g_{p}+\frac{\varepsilon}{1-\varepsilon}$$\frac${1}{2}
[…]
$g_{1}$ $=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\text{ \ \ ; \ \ }g_{p}=\frac{1-2\varepsilon}{2\left(1-\varepsilon\right)},$
which implies that
$g_{1}$ $=$ $2\Phi\left(l_{1}^{1/2}\right)-1=\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}$
$\Phi\left(l_{1}^{1/2}\right)$ $=$ $\frac{3}{4}+\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-3\varepsilon+2\varepsilon}{4\left(1-\varepsilon\right)}=\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}.$
$g_{p}$ $=$ $2\Phi\left(l_{p}^{1/2}\right)-1=\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}$
$\Phi\left(l_{p}^{1/2}\right)$ $=$ $\frac{3}{4}-\frac{\varepsilon}{2\left(1-\varepsilon\right)}=\frac{3-3\varepsilon-2\varepsilon}{4\left(1-\varepsilon\right)}=\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}.$
[…]
Proof of Lemma 4.9 We have to analyze,
$\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\mathbf{XX}^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\leq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\leq 1\right),\\ \left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\mathbf{XX}^{t}\frac{\Gamma_{r}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}_{r}^{t}\Gamma_{r}^{-1/2}\right\|}\geq\frac{1}{\sum_{j=2}^{p}l_{j}^{{}^{(r)}-1}\left(\mathbf{w}_{r}^{t}\mathbf{e}_{j}\right)^{2}}\right)+\\ \varepsilon\delta\left(r^{2}l_{1}^{{}^{(r)}-1/2}\mathbf{w}_{r}^{t}\mathbf{ee}^{t}\mathbf{w}_{r}\geq 1\right)\end{array}\right\},$
where $\mathbf{w}_{r}$ is a vector which yields the minimum.
[…]
(ii) If $\mathbf{w}_{r}$ $\in\left[L\left(\mathbf{e}_{2},\dots,\mathbf{e}_{p}\right)\right]^{c},$ we get for $r$ sufficiently large, $\min\left\{\varepsilon,\left(1-\varepsilon\right)\right\}.$ Then, for $r$ large enough, since $\varepsilon\leq 1/2,$
$D\left(\Gamma_{r},P_{\varepsilon,r}\right)=\varepsilon,$ (31)
[…]
Proof of Lemma 4.10. Take $\Gamma=\sum_{j=1}^{p}l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$. The depth of a matrix is given by, if $r>l_{1}^{1/2},$
$\min\left\{\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\right\}\text{.}$
Call $g_{m}=\min_{\mathbf{v\in}F_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right)$ and $g_{M}=\arg\max_{\mathbf{v\in}F_{r}^{\mathbf{e}}}g\left(\mathbf{v}\right)$. The different configurations that a matrix $\Gamma$ might have are given in the following table depending on where $\mathbf{v}_{1}$ and $\mathbf{v}_{p}$ belong to,
$A_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$
$B_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$
$F_{r}^{\mathbf{e}}$ $\mathbf{v}_{1},\mathbf{v}_{p}$ $\mathbf{v}_{p}$ $\mathbf{v}_{1}$ $\mathbf{v}_{1}$ $\mathbf{v}_{p}$
(i) $\mathbf{v}_{1},\mathbf{v}_{p}\in A_{r}^{\mathbf{e}}$: the depth of $\Gamma$ should be
$\min\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{m},\left(1-\varepsilon)\left(1-g_{1}\right)\right)$
(ii) $\mathbf{v}_{1},\mathbf{v}_{p}\in B_{r}^{\mathbf{e}}:$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{m}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(iii) $\mathbf{v}_{1},\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}:$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(iv) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in B_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{m}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(v) $\mathbf{v}_{p}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{1}\in B_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{m},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(vi) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(vii) $\mathbf{v}_{1}\in B_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}},$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{M}\right)\right)\right\}$
(viii) $\mathbf{v}_{1}\in A_{r}^{\mathbf{e}}$, $\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{M}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
(ix) $\mathbf{v}_{1},\mathbf{v}_{p}\in F_{r}^{\mathbf{e}}$
$\min\left\{\left(1-\varepsilon\right)\left(1-g_{1}\right)+\varepsilon,\left(1-\varepsilon\right)g_{p}+\varepsilon,\left(1-\varepsilon\right)g_{p},\left(1-\varepsilon)\left(1-g_{1}\right)\right)\right\}$
[…]
Proof of Theorem 4.11. (i) Lemma above says that if the estimators moves to the boundary of the domain of the eigenvalues, then $\varepsilon\geq 1/3$ and $\varepsilon^{\ast}\geq 1/3.$
[…]
Proof of Lemma 4.12. Since $P_{0}=\Phi$ then $D_{M}\left(\Phi\right)=1/2$ . Observe that $\delta\geq 0$ since given $Q$ any distribution on $\mathbb{R}^{p},$
[…]
$P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$ $\geq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\geq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)$
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $\geq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)$
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$D_{M}\left(P_{\varepsilon,Q}\right)$ $\geq$ $D\left(\beta I,P_{\varepsilon,Q}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$
$\Lambda\left(\varepsilon,P_{0}\right)$ $=$ $\inf_{Q}D_{M}\left(P_{\varepsilon,Q}\right)\geq\left(1-\varepsilon\right)D_{M}(P_{0})$
Since
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$ $\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
$\min\left\{\begin{array}[]{c}P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}$ $\leq$ $P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)$
$\leq$ $\left(1-\varepsilon\right)P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)+\varepsilon$
Then,
$\min\left\{P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{\varepsilon,Q}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}$
$\leq$ $\left(1-\varepsilon\right)\min\left\{P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P_{0}\left(\left|\mathbf{u}^{t}\mathbf{X}\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}+\varepsilon$
and
$D\left(\Gamma,P_{\varepsilon,Q}\right)\leq\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon.$
Consequently, (i) follows.
(ii) Set $\alpha=\frac{e}{1-\varepsilon}-\delta\left(\varepsilon,P_{0}\right).$ Then $0\leq\alpha<1/2.$ If $\Gamma\notin L\left(\alpha.P_{0}\right)$ we have that $D\left(\Gamma,P_{0}\right)<D_{M}\left(P_{0}\right)-\alpha.$ Therefore,
$D\left(\Gamma,P_{\varepsilon,Q}\right)$ $\leq$ $\left(1-\varepsilon\right)D\left(\Gamma,P_{0}\right)+\varepsilon<\left(1-\varepsilon\right)\left[D_{M}\left(P_{0}\right)-\frac{\varepsilon}{1-\varepsilon}+\delta\left(\varepsilon,P_{0}\right)\right]+\varepsilon$
$\leq$ $\left(1-\varepsilon\right)\left[D_{M}\left(P_{0}\right)-\frac{\varepsilon}{1-\varepsilon}+\frac{\Lambda\left(\varepsilon,P_{0}\right)-\left(1-\varepsilon\right)D_{M}\left(P_{0}\right)}{1-\varepsilon}\right]+\varepsilon=\Lambda\left(\varepsilon,P_{0}\right)$
and (ii) follows.
(iii) $L\left(\alpha,P_{0}\right)=L\left(\varepsilon/\left(1-\varepsilon\right)-\delta,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ Then, $\Gamma\in\left(L\left(\alpha,P_{0}\right)\right)^{c}$ implies, by (ii) that $\Gamma\in\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ or $\left(L\left(\alpha,P_{0}\right)\right)^{c}\subset\left(M\left(P_{\varepsilon,Q}\right)\right)^{c}$ which says that $M\left(P_{\varepsilon,Q}\right)\subset L\left(\alpha,P_{0}\right)\subset L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ for all distribution $Q.$ Since $D\left(\Gamma,P_{0}\right)=\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)$ and $D_{M}\left(P_{0}\right)=1/2,$ we have that
$L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)$ $=$ $\left\{\Gamma\succeq 0:\min\left(g\left(\mathbf{v}_{p}\right),1-g\left(\mathbf{v}_{1}\right)\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$=$ $\left\{\Gamma\succeq 0:1-g\left(\mathbf{v}_{1}\right)\geq g\left(\mathbf{v}_{p}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$\cup\left\{\Gamma\succeq 0:g\left(\mathbf{v}_{p}\right)\geq 1-g\left(\mathbf{v}_{1}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}$
$=$ $\left\{\Gamma\succeq 0:\frac{1}{2}+\varepsilon/\left(1-\varepsilon\right)\geq g\left(\mathbf{v}_{1}\right)\geq g\left(\mathbf{v}_{p}\right)\geq\frac{1}{2}-\varepsilon/\left(1-\varepsilon\right)\right\}.$
Then, $\left\|\Gamma\right\|_{op}\leq\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),$ $\left\|\Gamma^{-1}\right\|_{op}\leq 1/\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)$ if $\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P\right)$ and
$\left\|L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)\right\|$ $=$ $\sup_{\Gamma\in L\left(\varepsilon/\left(1-\varepsilon\right),P_{0}\right)}\max\left\{\frac{\left\|\Gamma\right\|_{op}}{\beta},\frac{\beta}{\left\|\Gamma^{-1}\right\|_{op}}\right\}$
$=$ $\max\left\{\frac{1}{\beta}\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right),\frac{\beta}{\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)}\right\}$
The bias functions turn out to be
$b_{S}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\beta^{-1}\lambda_{(1)}\left(\hat{\Gamma}\left(P\right)\right)=\beta^{-1}\left\|\hat{\Gamma}\left(P\right)\right\|_{op}$
$b_{I}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{1/2}\hat{\Gamma}^{-1}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{1/2}\right)=\beta\lambda_{(1)}\left(\hat{\Gamma}^{-1}\left(P\right)\right)=\beta\left\|\hat{\Gamma}^{-1}\left(P\right)\right\|_{op}.$
$B\left(\hat{\Gamma},\varepsilon,P_{0}\right)=\max\left\{\beta^{-1}\sup_{P\in\mathcal{P}_{\varepsilon}}\left\|\hat{\Gamma}\left(P\right)\right\|_{op},\text{ }\beta\sup_{P\in\mathcal{P}_{\varepsilon}}\left\|\hat{\Gamma}^{-1}\left(P\right)\right\|_{op}\right\}$
[…]
7.2 Depth as residual smallness concept and depth in the location-scale model
Proof of Lemma 5.1. Let $\mathcal{D}_{MR}^{E}\left(B,P\right){\ =}\inf_{U\mathbf{\in\mathbb{R}}^{p\times m}}P\left(\left\|Y-B^{t}\mathbf{X}\right\|\leq\left\|Y-U^{t}\mathbf{X}\right\|\right).$ Thus,
$P\left(\left\|Y-B^{t}\mathbf{X}\right\|\leq\left\|Y-U^{t}\mathbf{X}\right\|\right)$ $=$ $P\left(\left\|Y-B^{t}\mathbf{X}\right\|^{2}\leq\left\|Y-B^{t}\mathbf{X+}\left(B-U\right)^{t}\mathbf{X}\right\|^{2}\right)$
$=$ $P\left(0\leq-2\left\langle Y-B^{t}\mathbf{X,}\left(B-U\right)^{t}\mathbf{X}\right\rangle+\left\|\left(B-U\right)^{t}\mathbf{X}\right\|^{2}\right).$
Take U=BtV/2,U=B-tV/2, $V$ any matrix in $\mathbf{\mathbb{R}}^{p\times m}$, $t\geq 0.$ Then, we get that
$P\left(0\leq-2\left\langle Y-B^{t}\mathbf{X,}\left(U-B\right)^{t}\mathbf{X}\right\rangle+\left\|\left(U-B\right)^{t}\mathbf{X}\right\|^{2}\right)$
$=$ $P\left(0\leq t\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle+\frac{t^{2}}{4}\left\|V^{t}\mathbf{X}\right\|^{2}\right)$
$=$ $P\left(0\leq t\left(\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle+\frac{t}{4}\left\|V^{t}\mathbf{X}\right\|^{2}\right)\right)$
$=$ $P\left(0\leq\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle+\frac{t}{4}\left\|V^{t}\mathbf{X}\right\|^{2}\right).$
Since $\left[0\leq\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle+\frac{t}{4}\left\|V^{t}\mathbf{X}\right\|^{2}\right]\supseteq\left[0\leq\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle\right]$ for all $t\geq 0,$ we have that
$\inf_{U\mathbf{\in\mathbb{R}}^{p\times m}}P\left(\left\|Y-B^{t}\mathbf{X}\right\|\leq\left\|Y-U^{t}\mathbf{X}\right\|\right)=\underset{V\mathbf{\in\mathbb{R}}^{p\times m}}{\inf}\left[0\leq\left\langle Y-B^{t}\mathbf{X,}V^{t}\mathbf{X}\right\rangle\right]$
[…]
Proof of Lemma 5.2. It is easily proved that $\min_{\lambda}P\left[\left|y-\mu\right|\leq\left|y-\lambda\right|\right]=\min\left\{P(y\leq\mu),P(y\geq\mu)\right\}.$ Let us compute $\inf_{\gamma}P\left(\left[\left|\left|\frac{y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{y-\mu}{\gamma}\right|-1\right|\right]\right).$ Call $z=\left|y-\mu\right|.$ Then,
$\left\{\left|\left|\frac{y-\mu}{\sigma}\right|-1\right|\leq\left|\left|\frac{y-\mu}{\gamma}\right|-1\right|\right\}$ $=$ $\left\{\left|\left|\frac{y-\mu}{\sigma}\right|-1\right|^{2}\leq\left|\left|\frac{y-\mu}{\gamma}\right|-1\right|^{2}\right\}$
$=$ $\text{ }\left\{\frac{z^{2}}{\sigma^{2}}-2\frac{z}{\sigma}\leq\frac{z^{2}}{\gamma^{2}}-2\frac{z}{\gamma}\right\}=\left\{\frac{z}{\sigma^{2}}-\frac{2}{\sigma}\leq\frac{z}{\gamma^{2}}-\frac{2}{\gamma}\right\}$
$=$ $\left\{z\left(\frac{1}{\sigma^{2}}-\frac{1}{\gamma^{2}}\right)\leq 2\left(\frac{1}{\sigma}-\frac{1}{\gamma}\right)\right\}$
[…]
Therefore, we get that,
$\min_{\gamma}P\left(\left|\frac{z}{\sigma}-1\right|^{2}\leq\left|\frac{z}{\gamma}-1\right|^{2}\right)$
$=$ $\min_{\gamma}\left\{P\left(z\leq\frac{2}{\frac{1}{\sigma}+\frac{1}{\gamma}}\right)1_{\left(\sigma,\infty\right)}\left(\gamma\right)+P\left(z\geq\frac{2}{\frac{1}{\sigma}+\frac{1}{\gamma}}\right)1_{\left(0,\sigma\right)}\left(\gamma\right)\right\}$
[…]
$=$ $\min\left(P\left(z\leq\sigma\right),P\left(z\geq\sigma\right)\right)$
$=$ $\min\left(P\left(\mu-\sigma\leq y\leq\mu+\sigma\right),P\left(\left[y\geq\mu+\sigma\right]\cup\left[y\leq\mu-\sigma\right]\right)\right)$
and the depth turns out to be
$D_{LS}^{1}\left(\mu,\sigma,P\right)=\min\left\{\begin{array}[]{c}\min\left\{P(y\leq\mu),P(y\geq\mu)\right\},\\ \min\left(P\left(\mu-\sigma\leq y\leq\mu+\sigma\right),P\left(\left[y\geq\mu+\sigma\right]\cup\left[y\leq\mu-\sigma\right]\right)\right)\end{array}\right\}.$
[…]
Let $g\left(\delta\right)=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ with $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$ Let $\varepsilon_{0}$ be the fixed point of $g$, i.e.
$\varepsilon_{0}=g(\varepsilon_{0}).$ (36)
Let $P_{n}=(1-\varepsilon)\Phi\left(\cdot\right)+\varepsilon H_{n}\left(\cdot\right)$ be a sequence of contaminated distributions in the $\varepsilon$-contamination neighborhood $\mathcal{P}_{\varepsilon}\left(\Phi\right)$, $D_{n}$ the depth induced by these contaminated distributions and $\left\{\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right\}_{n}$ the deepest estimators associated to those distributions. Take the sets determined by $\left(\mu,\sigma\right)\in\mathbb{R}\times(0,\infty),$ $A^{(1)}\left(\mu,\sigma\right)=(-\infty,\mu-\sigma],$ $A^{(2)}\left(\mu,\sigma\right)=[\mu-\sigma,\mu],$ $A^{(3)}\left(\mu,\sigma\right)=[\mu,\mu+\sigma]$ and $A^{(4)}\left(\mu,\sigma\right)=[\mu+\sigma,\infty)$. Let $P_{\varepsilon}=(1-\varepsilon)\Phi$. Note that
$\forall n\geq 1:\,P_{n}(A^{(j)}(\mu,\sigma))\geq P_{\varepsilon}(A^{(j)}),\;j=1,2,3,4.$
Lemma 7.3
The following statements hold.
  • (i)
    Given $y_{0}<0$ fixed, the maximum of the function $h_{0}(x)=h\left(x,y_{0}\right),\;x>0$ occurs at $M\left(y_{0}\right)=\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$.
  • (ii)
    For $\epsilon\in(0,1/3)$ the following equivalence holds: $\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$ if and only if $\varepsilon<\varepsilon_{0}$.
Proof. (i) Fixed $y_{0},$ we want to maximize $h_{0}\left(x\right)=h\left(x,y_{0}\right)=\Phi\left(x\right)-\Phi\left(\frac{y_{0}+x}{2}\right).$ If $\varphi$ stands for the derivative of $\Phi,$ the derivative of $h_{0}\$turns out to be $h_{0}^{\prime}\left(x\right)=\varphi\left(x\right)-\frac{1}{2}\varphi\left(\frac{y_{0}+x}{2}\right)\geq 0$ if and only if $2exp\left(-\frac{1}{2}x^{2}\right)\geq exp\left(-\frac{1}{8}\left(\left(x+y_{0}\right)^{2}\right)\right)$ if and only if $\ln 2-\frac{3x^{2}}{8}+\frac{xy_{0}}{4}+\frac{y_{0}^{2}}{8}\geq 0$ if and only if $\frac{1}{3}\left(y_{0}-2\sqrt{y_{0}^{2}+6\ln 2}\right)$ $\leq$ $x$ $\leq$ $\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$. Then $M\left(y_{0}\right)=\frac{1}{3}\left(y_{0}+2\sqrt{y_{0}^{2}+6\ln 2}\right)$ yields the maximum value given by
$h_{0}\left(M\left(y_{0}\right)\right)=\Phi\left(\frac{y_{0}}{3}+\frac{2}{3}\sqrt{y_{0}^{2}+6\ln 2}\right)-\frac{1}{2}\Phi\left(\frac{2}{3}y_{0}+\frac{1}{3}\sqrt{y_{0}^{2}+6\ln 2}\right).$
(ii) If $\varepsilon<\varepsilon_{0}$, using (36), we have that
$(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)>(1-\varepsilon_{0})h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=g(\varepsilon_{0})=\varepsilon_{0}>\varepsilon.$
Assume now that $\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$ holds. Using again (36), we have that
$\varepsilon<(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}(1-\varepsilon_{0}))h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)=\frac{1-\varepsilon}{1-\varepsilon_{0}}\varepsilon_{0},$
[…]
Lemma 7.4
The function $g\left(\delta\right)=\left(1-\delta\right)h\left(\frac{1}{3}\left(y_{\delta}+2\sqrt{y_{\delta}^{2}+6\ln 2}\right),y_{\delta}\right)$ is strictly decreasing for $\delta\in\left(0,1/3\right).$ Therefore, $\delta<g\left(\delta\right)$ if and only if $\delta<\varepsilon_{0}.$
Proof. Note that $g(\delta)=(1-\delta)f(y_{\delta})=f\left(\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)\right)$with
$f\left(y\right)$ $=$ $h\left(\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2},y\right)$
$=$ $\Phi\left(\frac{1}{3}y+\frac{2}{3}\sqrt{y^{2}+6\ln 2}\right)-\Phi\left(\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}\right).$
Let us see that the function $f$ is decreasing in $y\in(-\infty,0).$ Note that
$\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2}=\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}+\frac{1}{3}\left(\sqrt{y^{2}+6\ln 2}-y\right)$
and $\frac{y}{3}+\frac{2}{3}\sqrt{y^{2}+6\ln 2}>0$ for all $y.$ The derivative of this function turns out to be
$f^{\prime}\left(y\right)$ $=$ $\varphi\left(\frac{1}{3}y+\frac{2}{3}\sqrt{y^{2}+6\ln 2}\right)\left(\frac{2}{3}+\frac{1}{3}\frac{y}{\sqrt{y^{2}+6\ln 2}}+\frac{1}{3}\left(\frac{y}{\sqrt{y^{2}+6\ln 2}}-1\right)\right)$
$-\varphi\left(\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}\right)\left(\frac{2}{3}+\frac{1}{3}\frac{y}{\sqrt{y^{2}+6\ln 2}}\right).$
Set $a=\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2},$ $b=\frac{1}{3}\left(\sqrt{y^{2}+6\ln 2}-y\right)>0$ and
$f^{\prime}\left(y\right)=\frac{1}{\sqrt{y^{2}+6\ln 2}}\left(\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)\right).$
Observe that $\frac{2}{3}y+\frac{1}{3}\sqrt{y^{2}+6\ln 2}>0$ if and only if $y>-\sqrt{\ln\left(4\right)}.$ Therefore, if $y\leq-\sqrt{\ln\left(4\right)},$ then $a\leq 0,$ $\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)<0$ and
$f^{\prime}\left(y\right)=\frac{1}{\sqrt{y^{2}+6\ln 2}}\left[\varphi\left(a+b\right)a-\varphi\left(a\right)\left(a+b\right)\right]<0$
[…]
Lemma 7.5
The following statements are valid.
  • (i)
    If $\varepsilon<1/5$ then there exists $\infty<x_{L}<x_{U}<\infty$ such that
    $\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\geq\varepsilon\text{.}$
  • (ii)
    If there exists $0<\alpha\leq 1$ verifying that $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))\geq\alpha$ and $\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}<\varepsilon<\varepsilon_{0},$ then there exists $\infty<x_{L}<x_{U}<\infty$ such that
    $\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon\alpha\text{.}$
    Note: $\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}=0$.1068.1068.
  • (iii)
    If $P_{\varepsilon}\left(A^{(j)}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)\underset{n\rightarrow\infty}{\rightarrow}0$ for some $j\in\left\{1,2,3,4\right\}$ then $\underset{n\rightarrow\infty}{\overline{\lim}}D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)$ $\leq$ $\varepsilon$ and $\varepsilon>1/5.$
  • (iv)
    Given $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ such that $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\mu_{n},\sigma_{n}\right)>0$ then $\underset{n\rightarrow\infty}{\underline{\lim}}P_{n}\left(A^{(j)}\left(\mu_{n},\sigma_{n}\right)\right)$ >> 0 for any $j\in\left\{1,2,3,4\right\}.$
Proof. (i) It suffices to take $x_{U}=-x_{L}$ such that $\Phi\left(x_{L}\right)=1/4.$ Indeed, considering $\mu=(x_{L}+x_{U})/2$ and $\sigma=(x_{L}-x_{U})/2$, the probabilities that define the notion of depth satisfy: $P_{n}(A^{(1)}\left(\mu,\sigma)\right)\geq P_{\varepsilon}(A^{(1)}(\mu,\sigma))=(1-\epsilon)\Phi(x_{L})=(1-\epsilon)\frac{1}{4}$, $P_{n}(A^{(2)}(\mu,\sigma)))\geq P_{\varepsilon}(A^{(2)}(\mu,\sigma))=(1-\epsilon)[\Phi(0)-\Phi(x_{L})]=(1-\epsilon)\frac{1}{4}$, and similarly for the other terms. Therefore $D\left(0,-x_{L}\right)\geq\left(1-\varepsilon\right)1/4>\varepsilon$ and the statement follows.
(ii) Given $y_{0}=y_{\varepsilon_{0}},$ and using Lemma 7.3 we can write $(1-\epsilon)h_{0}(x)$ == $P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+x}{2},\frac{x-y_{\varepsilon_{0}}}{2}\right)\right)$.The maximum of $h_{0}$ occurs at
$M\left(y_{\varepsilon_{0}}\right)=\frac{1}{3}\left(y_{\varepsilon_{0}}+2\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right)$
and the maximum value is given by
$h_{0}\left(M\left(y_{\varepsilon_{0}}\right)\right)=\Phi\left(\frac{y_{\varepsilon_{0}}}{3}+\frac{2}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right)-\frac{1}{2}\Phi\left(\frac{2}{3}y_{\varepsilon_{0}}+\frac{1}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}\right).$
If $\varepsilon<\varepsilon_{0}$, by (ii) of Lemma 7.3
$P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)=(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)>\varepsilon$
Moreover
$P_{\varepsilon}\left(A^{(1)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)=\varepsilon_{0}>\varepsilon,$
and then,
$P_{n}\left(A^{(j)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)>\varepsilon,j=1,3.$
Observe that $\frac{2}{3}y_{\delta}+\frac{1}{3}\sqrt{y_{\delta}^{2}+6\ln 2}>0$ if and only if $y_{\delta}\in\left[-\ln\left(4\right),\ln\left(4\right)\right]$ if and only if $\delta\geq\frac{\Phi\left(-\sqrt{\ln 4}\right)}{1+\Phi\left(-\sqrt{\ln 4}\right)}.$ Therefore $\frac{2}{3}y_{\varepsilon_{0}}+\frac{1}{3}\sqrt{y_{\varepsilon_{0}}^{2}+6\ln 2}>0,$and we can say that
$P_{\varepsilon}\left(A^{(2)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$ $>$
$P_{\varepsilon}\left(A^{(3)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$ $>$ $\varepsilon,$
and then $P_{n}\left(A^{(2)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)>\epsilon$.
The remaining set $A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)$ has probability greater than $\varepsilon\alpha.$ Set
$\eta=\frac{(1-\varepsilon)}{2}\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right),$
then we get, for $n$ large enough that
$(1-\varepsilon)\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$
$+\varepsilon H_{n}\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)$
$\geq$ $(1-\varepsilon)\Phi\left(A^{(4)}\left(\frac{y_{\varepsilon_{0}}+M\left(y_{\varepsilon_{0}}\right)}{2},\frac{M\left(y_{\varepsilon_{0}}\right)-y_{\varepsilon_{0}}}{2}\right)\right)+\varepsilon\alpha-\eta=\varepsilon\alpha+\eta,$
and (ii) follows.
(iii) Take $x_{L}$ obtained in item (i) and $j_{0}$ such that $\lim_{n}P_{\varepsilon}\left(A^{(j_{0})}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)=0.$ Then,
$\frac{(1-\varepsilon)}{4}\leq D\left(0,-x_{L}\right)\leq D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\leq P_{\varepsilon}\left(A^{(j_{0})}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right)+\varepsilon$
and the result follows.
[…]
Lemma 7.6
If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ then there exists $\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon$
if and only if $\varepsilon<\varepsilon_{0}.$
Proof. If $\varepsilon<\varepsilon_{0}$ then Lemma 7.5 (i) and (ii) says the conclusion. If $\underset{x>0}{inf}$ $\underset{n\rightarrow\infty}{\underline{lim}}H_{n}\left([x,\infty\right))=1$ and there exists $\infty<x_{L}<x_{U}<\infty$ such that $\varepsilon$ << $\underset{n\rightarrow\infty}{\underline{\lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)$ then
$P_{\varepsilon}\left(A^{(j)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)>\varepsilon\text{ forall }j\in\left\{1,2,3\right\}.$
If $P_{\varepsilon}\left(A^{(1)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)=(1-\epsilon)\Phi(x_{L})>\varepsilon$ then $x_{L}>\Phi^{-1}(\epsilon/(1-\epsilon))=y_{\epsilon}$. Moreover, by (i) of Lemma 1
$\varepsilon<P_{\varepsilon}\left(A^{(3)}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)\right)$ $=$ $(1-\epsilon)\left(\Phi(x_{U})-(\Phi(\frac{x_{L}+x_{U}}{2})\right)$
$=$ $(1-\epsilon)h(x_{L},x_{U})$
$\leq$ $(1-\varepsilon)h\left(M\left(y_{\varepsilon_{0}}\right),y_{\varepsilon_{0}}\right)$
$=$ $g(\epsilon).$
[…]
Lemma 7.7
Given the contaminated distributions $P_{n}=(1-\varepsilon)\Phi\left(\cdot\right)+\varepsilon\delta_{a_{n}}\left(\cdot\right)$ with point mass contaminations, $\delta_{a_{n}}$ verifying that $\lim_{n\rightarrow\infty}a_{n}=\infty,$ then there exists $\left(M_{0},s_{0}\right)\in\left(0,\infty\right)\times\left(0,\infty\right)$ such that the deepest estimators $\left\{\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\right\}_{n\in\mathbb{N}}\in\left(-M_{0},M_{0}\right)\times\left(s_{0},M_{0}\right)$ if and only $\varepsilon<\varepsilon_{0}.$
Proof: Take $\left\{\left(\mu_{n},\sigma_{n}\right)\right\}_{n\in\mathbb{N}}$ any sequence such that
$\lim_{n\rightarrow\infty}\mu_{n}+\sigma_{n}=\infty\text{ and}\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\mu_{n},\sigma_{n}\right)>0.$ (37)
Then,
(i) If $\lim_{n\rightarrow\infty}\mu_{n}-\sigma_{n}=$-$\infty$ then $\lim_{n\rightarrow\infty}P_{n}\left(A^{(1)}(\mu_{n},\sigma_{n})\right)=0$ and Lemma 7.5 (iv) yields a contradiction.
(ii) If $\mu_{n}+\sigma_{n}>a_{n}$ for $a_{n}$ large enough, then $\lim_{a\rightarrow\infty}P_{n}(A^{(1)}(\mu_{n},\sigma_{n}))=0$ and Lemma 7.5 (iv) yields a contradiction. Consequently,
[…]
Set $b=\lim_{n\rightarrow\infty}\mu_{n}-\sigma_{n}.$ This allows to say that $\lim_{n\rightarrow\infty}\mu_{n}=\infty$ since $\lim_{n\rightarrow\infty}\mu_{n}+\sigma_{n}-\left(\mu_{n}-\sigma_{n}\right)=\infty.$
(iii) If $\mu_{n}+\sigma_{n}<a_{n}$ for $n$ sufficiently large then $\lim_{a\rightarrow\infty}P_{n}\left(A^{(3)}(\mu_{n},\sigma_{n})\right)=0,$ and Lemma 7.5 (iv) yields a contradiction. Then $\mu_{n}+\sigma_{n}=a_{n}$ for $n$ sufficiently large.
(iv) Summing up, we can assume without loss of generality that
$\lim_{n\rightarrow\infty}P_{n}\left(A^{(1)}(\mu_{n},\sigma_{n})\right)$ $=$ $\left(1-\varepsilon\right)\Phi\left(b\right),\text{ \ \ }$
$\lim_{a\rightarrow\infty}P_{n}\left(A^{(2)}(\mu_{n},\sigma_{n})\right)$ $=$ $\left(1-\varepsilon\right)\left(1-\Phi\left(b\right)\right),\text{ \ \ }$
$\lim_{n\rightarrow\infty}P_{n}\left(A^{(3)}(\mu_{n},\sigma_{n})\right)$ $=$ $\varepsilon,\text{ \ \ }\lim_{a\rightarrow\infty}P_{n}\left(A^{(4)}(\mu_{n},\sigma_{n})\right)=\varepsilon.$
By Lemma 7.6 there exists $\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon$
if and only if $\varepsilon<\varepsilon_{0}.$ Therefore, for $n$ large enough,
$D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>D_{n}\left(\mu_{n},\sigma_{n}\right)$ (38)
[…]
Proof of Theorem 5.3. Given $\varepsilon>0,$
(i) Assume that the deepest estimator remains bounded in the $\varepsilon$-contamination neighborhood, that is there exists $M>0$ and $r>0$ such that
$\left(\mu\left(P\right),\sigma\left(P\right)\right)\in\left[-M,M\right]\times\left[r,M\right]$
for all $P\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$. An special case is given by the point mass contaminations. Thus Lemma 7.7 let us infer that $\varepsilon<\varepsilon_{0}$ and the breakdown point verifies that $\varepsilon^{\ast}\leq\varepsilon_{0}.$
(ii) Assume that there is a sequence of distributions $P_{n}\in\mathcal{F}_{\varepsilon}\left(\Phi\right)$ such that the associated deepest estimators $\left\{\left(\hat{\mu}_{n}\hat{\sigma}_{n}\right)\right\}$ verifies that $\lim_{n\rightarrow\infty}\left(\hat{\mu}_{n}+\hat{\sigma}_{n}\right)=\infty.$ Set $A_{n}^{(j)}=A^{(j)}(\hat{\mu}_{n},\hat{\sigma}_{n}).$ Then $\lim_{n\rightarrow\infty}P_{\varepsilon}\left(A_{n}^{(j)}\right)=0$ for $j=4$ and Lemma 7.5 (iii) says that $\varepsilon>1/5.$ Without loss of generality we can assume that there is a -$\infty\leq b$ such that
$\lim_{n\rightarrow\infty}P_{n}\left(A_{n}^{(1)})\right)$ $=$ $\left(1-\varepsilon\right)\Phi\left(b\right)+\varepsilon\alpha_{1},\text{ }$
$\lim_{a\rightarrow\infty}P_{n}\left(A_{n}^{(2)}\right)$ $=$ $\left(1-\varepsilon\right)\left(1-\Phi\left(b\right)\right)+\varepsilon\alpha_{2},\text{ \ \ }$
$\lim_{n\rightarrow\infty}P_{n}\left(A_{n}^{(3)}\right)$ $=$ $\varepsilon\alpha_{3},\text{ \ \ }\lim_{a\rightarrow\infty}P_{n}\left(A_{n}^{(4)}\right)=\varepsilon\alpha_{4}.$
Then,
$\lim_{n\rightarrow\infty}D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)\leq\varepsilon\alpha,\text{ with }\alpha=\min\left(\alpha_{3},\alpha_{4}\right).$ (39)
Suppose by contradiction that $\varepsilon<\varepsilon_{0}.$ By Lemma 7.5 (ii), there exists $\infty<x_{L}<x_{U}<\infty$ such that
$\underset{n\rightarrow\infty}{\underline{lim}}D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>\varepsilon\alpha.$ (40)
Then, by (39) and (40), for $n$ large enough we have
$D_{n}\left(\frac{x_{L}+x_{U}}{2},\frac{x_{U}-x_{L}}{2}\right)>D_{n}\left(\hat{\mu}_{n},\hat{\sigma}_{n}\right)$
[…]